### v1

# Protect Before Generate: Error Correcting Codes within Discrete Deep Generative Models

Abstract
Despite significant advancements in deep probabilistic models, learning low-dimensional discrete latent representations remains a challenging task. In this paper, we introduce a novel method that enhances variational inference in discrete latent variable models by leveraging Error-Correcting Codes to introduce redundancy in the latent representations. This redundancy is then exploited by the variational posterior to yield more accurate estimates, thereby narrowing the variational gap. Inspired by ECCs commonly used in digital communications and data storage, we demonstrate proof-of-concept using a Discrete Variational Autoencoder (DVAE) with binary latent variables and block repetition codes. We further extend this idea to a hierarchical structure based on polar codes, where certain latent bits are more robustly protected. Our method improves generation quality, data reconstruction, and uncertainty calibration compared to the uncoded DVAE, even when trained with tighter bounds such as the Importance Weighted Autoencoder (IWAE) objective. In particular, we demonstrate superior performance on MNIST, FMNIST, CIFAR10, and Tiny ImageNet datasets. The general approach of integrating ECCs into variational inference is compatible with existing techniques to boost variational inference, such as importance sampling or Hamiltonian Monte Carlo. We also outline the key properties ECCs must have to effectively enhance discrete variational inference.
Keywords: deep generative models, variational inference, discrete representations, error correcting codes
1 Introduction
Discrete latent space models seek to represent data using a finite set of features. Recent progress in generative models has increasingly favored these representations, as they are well-suited for datasets characterized by naturally discrete hidden states. However, effective learning of low-dimensional discrete latent representations is technically challenging. Vector Quantized-Variational Autoencoders [31, 21] stand out as solutions for this problem but rely on a non-probabilistic autoencoder, which does not provide uncertainty quantification in the latent space. To fit a fully probabilistic Variational Autoencoder (VAE) model [12], a common approach considers either Concrete [18] or Gumble-Softmax [11] approximations to sample from a discrete latent distribution in a reparameterizable manner [20, 17]. However, this approach leads to instabilities since the gradient variance is sensitive to the temperature that controls these approximations. The DVAE in [22, 30, 29], augments the binary latent representations with a set of continuous random variables, pairing each bit with a continuous counterpart where reparameterization can be done in a more stable manner after marginalizing the latent bits. The key distinction between the DVAE [22] and the DVAE++ [30] lies in their smoothing transformations: while [22] introduces spike-and-exponential transformations, [30] uses overlapping exponential distributions. These overlapping transformations are generalized in [29], enabling tighter variational bounds. We demonstrate the effectiveness of our method over a simplified version of the DVAE++ [30].
This work presents a novel method to improve variational inference and representation learning in generative models with discrete latent variables. In particular, for a latent variable model, we argue that one should use ECCs to introduce redundancy into the latent sample before the reconstruction decoder network processes it to generate the data. The variational approximation to the true posterior distribution can then exploit the added redundancy to provide more accurate estimates, reducing, in turn, the variational gap to the data likelihood.
Our approach is based on well-known digital communications and data storage techniques where information is protected with ECCs before transmission/storage to reduce the overall error rate during recovery. For different datasets, our results demonstrate that, compared to the uncoded DVAE, the DVAE with ECCs (Coded-DVAE) achieves superior generation quality, better data reconstruction, and critically calibrated uncertainty in the latent space. In Fig. 1, we highlight some representative results for both MNIST and FMNIST data sets. We note that the use of ECCs is a general design approach that is perfectly compatible with state-of-the-art techniques for improved variational inference, such as importance sampling [3, 27] or Hamiltonian Monte Carlo [32, 4]. In summary, our main contributions are:
  • We provide proof-of-concept results demonstrating that training deep generative models can be improved by ECC techniques, an idea that, to the best of our knowledge, is completely novel in the literature.
  • We formulate a coded version of DVAE using block repetition codes. We show that encoding/decoding of the block repetition code can be efficiently done with linear complexity.
  • We show that Coded-DVAE improves reconstruction, generation, and uncertainty calibration in the latent space when compared to the uncoded case using the same latent dimension, even when the uncoded DVAE is trained with tighter bounds such as the IWAE objective [3].
  • We discuss the generalization of this method to other coding schemes and introduce a hierarchical structure, inspired by polar codes [2], that effectively separates high-level information from finer details.
  • Through an extensive ablation study, we show that the enhancement in performance is not attributed to the increased dimensionality introduced by the redundancy from the ECC.
Figure 1: Comparison between uncoded and coded DVAE models with 8 latent bits, where the uncoded models are identified by the number of latent bits, and the coded models by their code rate. Fig. (a) presents examples of uncurated generation for FMNIST. Fig. (b) illustrates the evolution of the reconstruction and regularization terms of the loss. Fig. (c) depicts the graphical model of the uncoded DVAE, while Fig. (d) shows the graphical model of the coded DVAE.
2 Our baseline: the uncoded DVAE
This section introduces a simplified version of the uncoded DVAE [22, 30, 29], serving as the foundational model upon which the subsequent aspects of our work are constructed. Let $\boldsymbol{X}=\{\boldsymbol{x}_{0},\ldots,\boldsymbol{x}_{N}\}$ denote a collection of unlabelled data, where $\boldsymbol{x}_{i}$ represents a K-dimensional feature vector. While [22], [30] and [29] use Boltzmann machine priors, we consider a generative probabilistic model characterized by a simple low-dimensional binary latent variable $\boldsymbol{m}\in\{0,1\}^{M}$ comprising independent and identically distributed (i.i.d.) Bernoulli components $p(\boldsymbol{m})=\prod_{j=1}^{M}p(m_{j})=\prod_{j=1}^{M}\text{Ber}(\nu)$. Since backpropagation through discrete variables is generally not possible, a smoothing transformation of these binary variables is introduced. While the smoothing transformations proposed in [22] are limited to spike-and-X type of transformations, [30] show better results by using truncated exponential distributions:
$p(\boldsymbol{z}|\boldsymbol{m})=\prod_{j=1}^{M}p(z_{j}|m_{j}),\quad p(z_{j}|m_{j})$ $=\frac{e^{-\beta(z_{j}-m_{j})}}{Z_{\beta}},$ (1)
for $m_{j}\in\{0,1\}$, $z_{j}\in[0,1]$, and $Z_{\beta}=(1-e^{-\beta})/\beta$. The parameter $\beta$ serves as an inverse temperature term, similar to the one in the Gumbel-Softmax relaxation [11]. Given the simplicity of the defined binary prior, the complexity of the model is primarily determined by the likelihood function $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})=p(f_{\boldsymbol{\theta}}(\boldsymbol{z}))$, where the likelihood is a Neural Network (NN) (referred to as the decoder) with parameter set $\boldsymbol{\theta}$.
Variational family and inference
Following [22], we assume an amortized variational family of the following form:
$q_{\boldsymbol{\eta}}(\boldsymbol{m},\boldsymbol{z}|\boldsymbol{x})$ $=q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})p(\boldsymbol{z}|\boldsymbol{m}),\quad q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})=\prod_{j=1}^{M}\text{Ber}(g_{j,\boldsymbol{\eta}}(\boldsymbol{x})),$ (2)
where $g_{\boldsymbol{\eta}}(\boldsymbol{x})$ represents a parameterized function; here, a NN (referred to as the encoder) with parameter set $\boldsymbol{\eta}$. Inference is achieved by maximizing the Evidence Lower Bound (ELBO), which can be expressed as
[…]
where the first term corresponds to the reconstruction of the observed data and the second term is the Kullback-Leibler (KL) Divergence between the variational family and the binary prior distribution, which acts as a regularization term. This can be computed in closed form as $\mathcal{D}_{KL}\big{(}q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})||p(\boldsymbol{m})\big{)}=\sum_{j=1}^{M}\Big{[}q_{j}\log\frac{q_{j}}{\nu}+(1-q_{j})\log\frac{1-q_{j}}{1-\nu}\Big{]}$, where $q_{j}=q_{\boldsymbol{\eta}}(m_{j}=1|\boldsymbol{x})$. The reconstruction term needs to be approximated via Monte Carlo. Since $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})$ does not depend on the binary latent variable $\boldsymbol{m}$, we can marginalize the posterior distribution as
$q_{\boldsymbol{\eta}}(\boldsymbol{z}|\boldsymbol{x})$ $=\prod_{j=1}^{M}q_{\boldsymbol{\eta}}(z_{j}|\boldsymbol{x}),\quad q_{\eta}(z_{j}|\boldsymbol{x})=\sum_{k=0}^{1}q_{\eta}(m_{j}=k|\boldsymbol{x})p(z_{j}|m_{j}=k).$ (4)
As shown in [30], the corresponding inverse Cumulative Density Function (CDF) is given by
$F^{-1}_{q_{\boldsymbol{\eta}}(z_{j}|\boldsymbol{x})}(\rho)=-\frac{1}{\beta}\log\left(\frac{-b+\sqrt{b^{2}-4c}}{2}\right),$ (5)
where $b=\big{(}\rho+e^{-\beta}(q_{j}-\rho)\big{)}/(1-q_{j})-1$ and $c=-[q_{j}e^{-\beta}]/(1-q_{j})$. The (5) is a differentiable function that converts a sample $\rho$ from an independent uniform distribution $\mathcal{U}(0,1)$ into a sample from $q_{\eta}(\boldsymbol{z}|\boldsymbol{x})$. Thus, we can apply the reparameterization trick to sample from the latent variable $\boldsymbol{z}$ and optimize the ELBO with respect to the model’s parameters. Algorithm 1 outlines the training procedure for the uncoded DVAE.
Algorithm 1 Training the model with uncoded inference.
1:  Input: training data $\boldsymbol{x}_{i}$.
2:  repeat
3:     $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{m}|\boldsymbol{x}_{i})\leftarrow$ forward encoder $g_{\boldsymbol{\eta}}(\boldsymbol{x}_{i})$
4:     $\boldsymbol{z}\leftarrow$ sample from (5)
5:     $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})\leftarrow$ forward decoder $f_{\boldsymbol{\theta}}(\boldsymbol{z})$
6:     Compute ELBO according to (3)
7:     $\boldsymbol{\theta},\boldsymbol{\eta}\leftarrow Update(ELBO)$
8:  until convergence
3 Improving inference by adding redundancy to latent vectors
In the DVAE framework [22, 30, 29], the authors use Boltzmann machines as priors instead of the independent prior $p(\boldsymbol{m})$ presented in Section 2. While these complex priors increase the model’s flexibility and can produce competitive results, our objective is to enhance inference through model design by maintaining the simpler independent prior. This approach would improve interpretability and encourage the model to learn independent components in the latent space, which is essential for capturing potentially disentangled representations.
VAEs [12] are often viewed as lossy compression models, where the goal is to minimize reconstruction error while imposing regularization through a prior distribution. However, our approach is better understood from a generative standpoint. We first sample a latent vector $\boldsymbol{m}$, generate an observation $\boldsymbol{x}$, and focus on minimizing the error rate when recovering $\boldsymbol{m}$ from $\boldsymbol{x}$. Achieving this requires the variational approximation to be sufficiently accurate. In fields where reliable data transmission or storage is important, introducing ECCs is a well-established approach to reduce the error rate when estimating a discrete source $\boldsymbol{m}$ transmitted through a noisy channel with output $\boldsymbol{x}$. Estimating $\boldsymbol{m}$ from $\boldsymbol{x}$ implies approximating the true and unknown posterior distribution $p(\boldsymbol{m}|\boldsymbol{x})$ with a proposed $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$. The gap between $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ and $p(\boldsymbol{m}|\boldsymbol{x})$ is precisely the variational gap. We propose employing ECCs to safeguard $\boldsymbol{m}$ with controlled and known redundancy that can be leveraged by the variational posterior $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ by design. This way, it is possible to reduce the mistakes committed when comparing $\boldsymbol{m}$ with samples drawn from $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$, obtaining a tighter approximation to the true posterior $p(\boldsymbol{m}|\boldsymbol{x})$, therefore reducing the gap to optimal inference.
ECCs play a crucial role in information theory and digital communications by enabling reliable data transmission over unreliable channels [19]. They introduce redundancy into the transmitted data, allowing the receiver to detect errors and, in many cases, correct them without retransmission. In his seminal work, Shannon [25] demonstrated that arbitrarily reliable communication is possible through error correction. Our approach builds on the idea that the generative model in Fig. 1(c) can be conceptualized as a communication system, where the bits sampled from $p(\boldsymbol{m})$ undergo continuous modulation into $\boldsymbol{z}$ and are then transmitted through a nonlinear communication channel (in this setting, the decoder NN) characterized by the input/output response $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})=p(f_{\boldsymbol{\theta}}(\boldsymbol{z}))$. In this scenario, the complexity of the channel is essential since it is necessary to account for the intricate nature of the data at its output (e.g., complex images). Following this idea, the process of inference via $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ can be thought of as deciphering the latent variable $\boldsymbol{m}$ given the observed data $\boldsymbol{x}$, where the encoder NN plays the role of the channel equalizer, trying to reverse the channel’s effects without knowing the bit correlations from the ECC.
4 Coded DVAE
This section extends the previously described DVAE, introducing an ECC over $\boldsymbol{m}$. We refer to this model as coded DVAE. In ECCs, we augment the dimensionality of the binary latent space from $M$ to $D$ in a controlled and deterministic manner, where $R=M/D$ is the coding rate. An ECC is typically designed so that the $2^{M}$ possible codewords are separated as much as possible in the space of binary vectors of $D$ bits. This facilitates algorithms in detecting and/or correcting errors by searching for the nearest code word. A random choice of the codewords brings what is known as a random block code [25]. While they are known to be very robust and amenable to theoretical analysis, their lack of structure makes them computationally intractable since we have to rely on codeword enumeration during the encoding/decoding process. In Appendix J, we include the formulation of a random code’s encoding/decoding process within the DVAE model.
Instead, we adopt a much simpler linear coding scheme, namely repetition codes. In a repetition code, each bit of the original message $\boldsymbol{m}$ is repeated multiple times to create the encoded message $\boldsymbol{c}$. Intuitively, the more times an information bit is repeated, the better it is protected. Our experiments consider uniform $(M,D)$ repetition codes where all bits are repeated $L$ times, resulting in codewords of dimension $D=ML$ and a coding rate of $R=1/L$. Note that repetition codes represent a special case of linear ECCs since each codeword can be deterministically computed by multiplying a binary vector $\boldsymbol{m}$ by an $M\times D$ generator matrix $\mathbf{G}$, such that $\boldsymbol{c}=\boldsymbol{m}^{T}\mathbf{G}$, where $u$-th row, with $u=1,\ldots,M$, has entries equal to one at columns $L(u-1)+1,L(u-1)+2,\dots,Lu$, and zero elsewhere. For example, for $M=3$ and $L=2$, the generator matrix of the $(3,6)$ repetition code is
$\mathbf{G}=\begin{bmatrix}1&1&0&0&0&0\\ 0&0&1&1&0&0\\ 0&0&0&0&1&1\end{bmatrix}.$ (6)
The generative process of the coded DVAE follows similarly to the uncoded case, and it is represented in Fig. 1(d). We assume the same prior distribution $p(\boldsymbol{m})$, but in this case the samples $\boldsymbol{m}$ are deterministically encoded using $\mathbf{G}$. Now, the smoothing $\boldsymbol{z}$ transformations are defined over $\boldsymbol{c}$
$p(\boldsymbol{z}|\boldsymbol{c})=\prod_{j=1}^{D}p(z_{j}|c_{j}),\quad p(z_{j}|c_{j})$ $=\frac{e^{-\beta(z_{j}-c_{j})}}{Z_{\beta}},$ (7)
for $z_{j}\in[0,1]$, $c_{j}\in\{0,1\}$ and $Z_{\beta}=(1-e^{-\beta})/\beta$. The likelihood $p(\boldsymbol{x}|\boldsymbol{z})$ is again of the form $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})=p(f_{\boldsymbol{\theta}}(\boldsymbol{z}))$. Note that, compared to the uncoded case, we have a larger input dimensionality to the decoder NN $f_{\boldsymbol{\theta}}(\boldsymbol{z})$. When comparing uncoded vs. coded DVAEs, the structure of the decoder NN $f_{\boldsymbol{\theta}}(\boldsymbol{z})$ (detailed in Appendix A) is equal in both cases except for the first Multilayer Perceptron (MLP) layer that attacks the input $\boldsymbol{z}$. Therefore, if a rate $R=1/L$ repetition code is used, the number of additional parameters of the $f_{\boldsymbol{\theta}}(\boldsymbol{z})$ NN is given by $(L-1)\times h$, where $h$ is the dimension of the first hidden space of $f_{\boldsymbol{\theta}}(\boldsymbol{z})$.
Variational family and inference
The repetition code introduces correlations between the bits in $\boldsymbol{c}$ that we will exploit to obtain an improved variational bound. We again assume a variational family factorizing as
$q_{\boldsymbol{\eta}}(\boldsymbol{m},\boldsymbol{z}|\boldsymbol{x})=q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})p(\boldsymbol{z}|\boldsymbol{c})$ (8)
where $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})=\prod_{u=1}^{M}q_{\boldsymbol{\eta}}(m_{u}|\boldsymbol{x})$ is computed in two steps. First, we construct an encoder NN $\mathbf{g}_{\boldsymbol{\eta}}(\boldsymbol{x})$ similar to that of (2), that retrieves the probabilities of the bits in $\boldsymbol{c}$ from $\boldsymbol{x}$ without exploiting the correlations introduced by the repetition code:
$q^{u}_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x})$ $=\prod_{j=1}^{D}\text{Ber}(g_{j,\boldsymbol{\eta}}(\boldsymbol{x})),$ (9)
where the $u$ superscript serves as a reminder that this posterior does not exploit the redundancy introduced by the ECC.
Now, we utilize the known redundancy introduced by the ECC to constrain the solution of $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}|\boldsymbol{x})$, given that each bit from $\boldsymbol{m}$ has been repeated $L$ times to create $\boldsymbol{c}$. To do so, we follow a soft decoding approach, where the marginal posteriors of the information bits are derived from the marginal posteriors of the encoded bits, exploiting the repetition code’s known structure. In the case of repetition codes, we compute the all-are-zero and the all-are-ones products of probabilities of the bits in $\boldsymbol{c}$ that are copies of the same message bit and renormalize as
$q(m_{u}=1|\boldsymbol{x})$ $=\frac{1}{Z}\prod_{j=L(u-1)+1}^{Lu}g_{j,\boldsymbol{\eta}}(\boldsymbol{x})\doteq\frac{g^{+}_{u,\boldsymbol{\eta}}(\boldsymbol{x})}{Z},$ (10)
$q(m_{u}=0|\boldsymbol{x})$ $=\frac{1}{Z}\prod_{j=L(u-1)+1}^{Lu}(1-g_{j,\boldsymbol{\eta}}(\boldsymbol{x}))\doteq\frac{g^{-}_{u,\boldsymbol{\eta}}(\boldsymbol{x})}{Z},$ (11)
for $u=1,\ldots,M$ and $Z=\left(g^{+}_{u,\boldsymbol{\eta}}(\boldsymbol{x})+g^{-}_{u,\boldsymbol{\eta}}(\boldsymbol{x})\right)^{-1}$. This approach can be seen as a soft majority voting strategy, enabling the recovery of the original information vector even if some bits in the inferred encoded word are corrupted. All operations in (10) preserve the gradients concerning the parameters in the encoder $\mathbf{g}_{\boldsymbol{\eta}}(\boldsymbol{x})$. We implement them in the log domain for stability.
When compared to the uncoded case, as in the likelihood term, we consider the same NN structure for the encoder $\mathbf{g}_{\boldsymbol{\eta}}(\boldsymbol{x})$ where both cases only differ in the last MLP layer. The additional overhead in the coded cases requires $(L-1)\times h^{\prime}$ parameters in the last layer, where $h^{\prime}$ is the dimension at the output of the last layer. In Appendix H.2, we conduct an ablation study on the number of trainable parameters to demonstrate that the improvement in performance does not stem from this increase in the number of parameters, but rather from the incorporation of the ECC in the latent space.
Algorithm 2 Training the coded DVAE with repetition codes.
1:  Input: training data $\boldsymbol{x}_{i}$, matrix $\mathbf{G}$.
2:  repeat
3:     $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}|\boldsymbol{x}_{i})\leftarrow$ forward encoder $g_{\boldsymbol{\eta}}(\boldsymbol{x}_{i})$
4:     $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x}_{i})\leftarrow$ soft decoding by aggregating $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}|\boldsymbol{x}_{i})$ according to (10)
5:     $q_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x}_{i})\leftarrow$ repeat posterior bit probabilities $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x}_{i})$ according to $\mathbf{G}$
6:     $\boldsymbol{z}\leftarrow$ sample from (5)
7:     $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})\leftarrow$ forward decoder $f_{\boldsymbol{\theta}}(\boldsymbol{z})$
8:     Compute ELBO according to (3)
9:     $\boldsymbol{\theta},\boldsymbol{\eta}\leftarrow Update(ELBO)$
10:  until convergence
Soft encoding for efficient reparameterization
Given the variational family in (8), the ELBO matches the expression in (3). However, the reparameterization trick in (5) requires independent bits, which is not the case in $\boldsymbol{c}$. To efficiently circumvent this issue during training, we employ a soft encoding approach. With soft encoding, a marginal probability is computed for each bit in the codeword $\boldsymbol{c}$, taking into account the structure of the ECC and the marginal probabilities of the information bits. For a repetition code, this involves simply replicating the posterior probabilities $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})=\prod_{u=1}^{M}q_{\boldsymbol{\eta}}(m_{u}|\boldsymbol{x})$ for each copy of the same information bit. Hence, we treat the bits in $\boldsymbol{c}$ as independent but distributed according to $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$. The Algorithm 2 shows the training pseudo-code.
When marginalizing $\boldsymbol{c}$ using the soft encoding marginals, we disregard potential correlations between the coded bits. For instance, with repetition codes, sampling from the marginals could produce inconsistent bits, leading to an invalid codeword. However, since we do not sample the coded bits during training but instead propagate their marginal probabilities, we consider this approximation to have minimal negative impact. In fact, it can be seen as a form of probabilistic dropout, which enhances robustness during training. It is important to note that when sampling from the generative model in test time, we use hard bits encoded into valid codewords, yielding visually appealing samples, indicating that our training approach is reliable.
Related work
While the use of deep neural networks and generative models in digital communications problems have been profusely reported in recent years (see [36], [5], [9], [33], and [26] for representative examples), the use of ECC techniques as a design tool in machine learning is scarce. The most prominent example is [7], where the authors address multiclass learning problems via ECCs. In [1], the authors proposed using linear codes for applications that may require removing the trace of a sample from the system, e.g., a user requests their data to be deleted or corrupted data is discovered. They address a regression problem by introducing a coded learning protocol that employs linear encoders to divide the training data before the learning phase. More recently, in [35] the authors introduced ECCs to improve code-to-code translation using transformers.
4.1 Reconstruction
We first evaluate the model’s performance of reconstructing data by examining its uncoded and coded versions across different configurations, varying the number of information bits and code rates. The introduction of the repetition code led to improved reconstruction and smaller KL values, indicating that the posterior latent features are disentangled and less correlated. In Appendices B, C, D and E, we show the behavior of the ELBO loss function for all the models and datasets.
Image reconstruction quality.   In the table included in Fig. 2, we first quantify the quality of the reconstructions in FMNIST by measuring the Peak Signal-To-Noise Ratio (PSNR) in the test set. The results for the rest of the datasets are provided in sections B, C, D and E of the Appendix. In all the cases, the coded models yield higher PSNR values than their uncoded counterparts, indicating a superior performance in reconstruction. This improvement is also evident by visual inspection of Fig. 2, where the coded models exhibit a greater ability to capture details in the images for the same latent dimension. We observe a general improvement in PSNR as we increase the number of information bits, i.e., as we augment the latent dimensionality of the model. This increase in the number of available latent vectors provides greater flexibility, enabling the models to capture the underlying structure of the data more effectively. We also observe a general improvement in PSNR as we decrease the code rate, i.e., as we add more redundancy. Note that adding redundancy does not increase the model’s flexibility, since the information bits determine the number of latent vectors. However, coded models yield more accurate and detailed reconstructions.
Model PSNR Acc Conf. Acc Entropy
uncoded 5 14.477 0.536 0.536 0.237
coded 5/50 16.241 0.647 0.700 1.899
coded 5/80 16.624 0.688 0.748 2.180
coded 5/100 16.702 0.700 0.757 2.256
uncoded 8 15.598 0.594 0.595 0.467
coded 8/80 17.318 0.750 0.816 2.905
coded 8/160 17.713 0.783 0.831 3.637
coded 8/240 17.861 0.799 $\mathbf{0.893}$ 4.000
uncoded 10 16.000 0.644 0.648 0.659
coded 10/100 17.694 0.790 0.850 3.879
coded 10/200 18.009 0.814 0.871 4.609
coded 10/300 $\mathbf{18.111}$ $\mathbf{0.817}$ 0.870 5.076
Figure 2: Reconstruction performance over the test set in FMNIST. The figure at the left shows an example of reconstructed test images obtained with different model configurations. Observe that more details are visualized as we increase the number of bits in the latent space and decrease the coding rate. The table at the right includes reconstruction metrics. Acc is the semantic accuracy and Conf. Acc the confident semantic accuracy. Entropy is the average entropy of $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ in the test set.
Semantic accuracy.   As the PSNR operates at the pixel level, it does not account for the semantic errors committed by the model. For example, if the model incorrectly reconstructs a nine instead of a four in the MNIST dataset, the PSNR may still yield a large value due to the similarity between the two images. Nonetheless, this would represent a severe failure in correct reconstruction of the intended class. Therefore, we additionally evaluate the reconstruction accuracy, ensuring that the model successfully reconstructs images within the same class as the original ones. For this purpose, we trained an image classifier for each dataset and compared the reconstructed images’ predicted labels against the originals’ ground truth labels. Additionally, we provide a confident reconstruction accuracy. While the reconstruction accuracy is computed across the entire dataset partitions, for the confident accuracy, we only consider those images projected into a latent vector with a probability exceeding 0.4.111Namely, we do not count errors when the Maximum a Posteriori (MAP) value of $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ is below 0.4.. Results for FMNIST are detailed in the table included in Fig.2, and corresponding results for MNIST are available in Table 2 within Appendix C. In light of the results, we can conclude that introducing an ECC in the model allows for latent spaces that better capture the semantics of the images while employing the same number of latent vectors, significantly outperforming the uncoded models in terms of accuracy in all the cases.
Figure 3: Example of erroneous reconstructions in FMNIST using the 4 most-probable a posteriori latent vectors. The first column in each image shows the original input to the model, while the second column displays the reconstructions. The $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ probability is indicated in each row.
Posterior uncertainty calibration.   Finally, also in the table included in Fig.2, we report the average entropy of the variational posterior $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ over the test set. The low entropy observed in the uncoded models suggests a low uncertainty when the model projects data points into the latent space, which could be advantageous if the model consistently assigned high probability to the correct latent vectors. However, the semantic accuracy results demonstrate this is not true in the uncoded model. In other words, the uncoded variational family projects images into the wrong class with high confidence. This indicates the uncertainty of the uncoded case is severely miscalibrated.
Coded models, on the other hand, improve semantic accuracy and present a larger entropy. This suggests that i) the coded DVAE is aware that multiple latent vectors might be related to the image class and ii) that the model posterior shows large uncertainties (high entropy) for certain images for which the model has not properly identified the class. We illustrate this in Fig. 3, where we show some images that were selected so that the MAP latent word from $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ induces class reconstruction errors. We display the reconstruction of the 4 most probable latent vectors and their corresponding probabilities. Observe that the uncoded model is confident no matter the reconstruction outcome while, in the coded posterior, the uncertainty is much larger. These results are also observed for MNIST, indicating that the posterior distribution in coded models exhibits a better uncertainty calibration. Note also that the increase in the number of latent bits (from 8 to 10) does not result in an excessive increase in the entropy despite the exponential growth of the number of vectors.
4.2 Generation
Model BER WER LL train LL test
uncoded 5 0.051 0.195 -266.157 -267.703
coded 5/50 0.011 0.046 -239.379 -241.882
coded 5/80 0.008 0.039 -227.550 -232.992
coded 5/100 0.010 0.049 -238.206 -241.404
uncoded 8 0.089 0.384 -247.964 -249.880
coded 8/80 0.021 0.144 -227.550 -232.992
coded 8/160 0.027 0.189 -228.585 -235.819
coded 8/240 0.037 0.231 -231.679 -238.459
uncoded 10 0.142 0.622 -242.842 -244.997
coded 10/100 0.040 0.321 -222.011 -230.772
coded 10/200 0.044 0.341 -223.748 -234.849
coded 10/300 0.045 0.349 -226.504 -238.647
Figure 4: Evaluation of generation in FMNIST. The figure at the left shows an example of randomly generated, uncurated FMNIST images. The table at the right shows the quantitative results on the evaluation of the BER, WER, and log-likelihood (LL).
In this section, we evaluate the model for the image generation task. In Fig. 4, we show examples of randomly generated images using different model configurations in FMNIST. Results for the rest of the datasets are available in Appendices B, C, D, and E. These results are consistent with the ones obtained in reconstruction since we can observe that the coded models can generate more detailed and diverse images. Both uncoded and coded models generate more intricate and varied images with increased information bits. However, if the number of latent vectors becomes too large for the dataset’s complexity, not all words in the codebook are specialized during model training. This leads to generation artifacts, images where different classes of objects are overlapped. A visual inspection of Fig. 4 suggests these artifacts are more frequent in the uncoded case. Note that, since we are dealing with discrete latent variables, we could simply detect and prune uninformative vectors.
Accuracy metrics in generation.   The improved inference given by the repetition code can also be tested by generating images using the generative model and counting errors using the MAP solution of the variational distribution $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$. The table included in Fig. 4 reports the BER and WER for FMNIST. As expected, at the same number of latent bits, the coded models significantly reduce both the BER and WER w.r.t. the uncoded case. Note also that the error rates grow with the number of latent bits, which is expected due to the increased complexity of the inference process. We may commit more errors by taking the MAP, but errors typically fall in consistent reconstructions (latent words that also reconstruct the same type of image), as the results presented in Section 4.1 indicated.
Log-likelihood.   We additionally estimated the log-likelihood (LL). Results for FMNIST with different model configurations, estimated through importance sampling with 300 samples per observation, are presented in the table included in Fig. 4, please refer to Appendix I for further details. We observe that coded models consistently outperform their uncoded counterparts for both train and test sets, aligning with the results previously presented. Thus, we can argue that the introduction of repetition codes in the definition of the model allows for an improved inference and tighter posterior approximation. We draw similar conclusions for other datasets.
We observe a general improvement in LL values as we increase the number of information bits, i.e., as we augment the latent dimensionality of the model and its flexibility. However, reducing the code rate does not lead to an improvement in log-likelihood. We argue that this might indicate overfitting of the decoder, as the LL deteriorates while reconstruction metrics improve. We must note that we use feed-forward networks at the decoder’s input. However, this may not be appropriate for the correlations we present in our coded words. We might overcome this overfitting tendency by using an architecture that properly leverages these coded bits correlations.
Figure 5: Reconstruction results for CIFAR10 (left) and Tiny ImageNet (right).
Figure 6: Generation results for CIFAR10 (left) and Tiny ImageNet (right).
4.3 Additional results with CIFAR10 and Tiny ImageNet
Since MNIST-like datasets are rather simple, it is difficult to assess the true gain in performance resulting from the introduction of ECCs proposed in our model. This section presents additional results using CIFAR10 and Tiny ImageNet, which contain colored images with more intricate shapes, patterns, and greater diversity than the previous datasets. We trained uncoded and coded models using different configurations to gain intuition regarding the effect of introducing the ECC. For a reference, in the case of the DVAE++ [30], the authors needed 128 binary latent variables to achieve state-of-the-art performance in generation and reconstruction for this dataset. They employed a more intricate model than the one introduced in this study, featuring Boltzmann Machine priors. In Fig. 5 we show examples of reconstruction using different configurations of the model and in Fig. 6 we show examples of randomly generated images. Additional results are provided in Appendices D and E. The results are consistent with those presented in previous sections; but in this case, the difference in performance is even more pronounced. We observe that the uncoded DVAE cannot decouple spatial information from the images and project it in the latent space. Nevertheless, the coded DVAE shows particular promise for learning low-dimensional discrete latent representations in complex datasets. Note that we used a rather simple architecture as we want to focus on the gain obtained only by introducing ECCs in the latent space.
5 Beyond repetition codes
We have presented compelling proof-of-concept results that incorporating ECCs, like repetition codes, into DVAEs can improve performance. We believe this opens a new path for designing latent probabilistic models with discrete latent variables. Although a detailed analysis of the joint design of ECC and encoder-decoder networks is beyond the scope of this work, we will outline key properties that any ECCs must satisfy to be integrated within this framework.
  • Scalable hard encoding $\big{(}\boldsymbol{m}\rightarrow\boldsymbol{c}\big{)}$. Our model requires hard encoding for generation once the model is trained. This process should have linear complexity in $M$.
  • Scalable soft encoding $\big{(}p(\boldsymbol{m})\rightarrow p(\boldsymbol{c})\big{)}$. Soft encoding is required during training for reparameterization. This process should also have linear complexity in $M$.
  • Scalable soft decoding $\big{(}p(\boldsymbol{c})\rightarrow p(\boldsymbol{m})\big{)}$. Our model employs soft-in soft-out (SISO) decoding during inference. This process should again be linearly complex in $M$.
  • Differentiability. Both encoding and decoding processes must be differentiable w.r.t the inputs to enable gradient computation and backpropagation.
Figure 7: Graphical model of the hierarchical coded DVAE.
Since Shannon’s seminal work [25], researchers have developed effective ECC schemes that meet these properties, including state-of-the-art ECCs such as Low Density Parity Check (LDPC) codes [8], or polar codes [2]. Efforts have also focused on developing efficient SISO decoders, such as the sum-product algorithm [15].
Inspired by polar codes [2], we present a hierarchical coded DVAE with two layers of latent bits. In this model, the latent bits $\boldsymbol{m}_{1}$ are encoded using a repetition code in the first layer, producing $\boldsymbol{c}_{1}$ and $\boldsymbol{z}_{1}$. Simultaneously, the vector $\boldsymbol{m}_{2}$ is linearly combined with $\boldsymbol{m}_{1}$ using modulo 2 operations $\big{(}\boldsymbol{m}_{1}\oplus\boldsymbol{m}_{2}\big{)}$ and then encoded using a repetition code, yielding $\boldsymbol{c}_{2}$ and $\boldsymbol{z}_{2}$. Both soft vectors are concatenated and fed to the decoder NN to generate $\boldsymbol{x}$. The model provides stronger protection for $\boldsymbol{m}_{1}$, as it appears in both branches of the generative model. Inference follows a similar approach to the coded DVAE, incorporating the linear combination of $\boldsymbol{m}_{1}$ and $\boldsymbol{m}_{2}$ used in the second branch. This hierarchical structure allows the model to effectively separate high-level information from finer details, as we show in the results presented in Appendix G.
6 Conclusion
This paper presents the first proof-of-concept demonstration that safeguarding latent information with ECCs within deep generative models holds promise for enhancing overall performance. By integrating redundancy into the latent space, the variational family can effectively refine the inference network’s output according to the structure of the ECC. Our findings underscore the efficacy of simple and efficient ECCs, like repetition codes, showcasing remarkable improvements over a lightweight version of the DVAE introduced in [30].
Furthermore, our work reveals numerous avenues for future research. Firstly, investigating decoder architectures capable of efficiently utilizing the correlations and structure introduced by the ECCs, in contrast to the feed-forward networks employed in this study. We also contemplate exploring more complex and robust coding schemes, conducting theoretical analyses aligned with Shannon’s channel capacity and mutual information concepts to determine the fundamental parameters of the ECC needed to achieve reliable variational inference, exploring different modulations, and integrating these concepts into state-of-the-art models based on discrete representations.
7 Acknowledgements
The work was partially supported by the Ministerio de Ciencia, Innovación y Universidades, Agencia Estatal de Investigación, under grant PID2021-123182OB-I00, MCIN/AEI/10.13039/501100011033, co-financed by European Regional Development Fund (ERDF), “A way of making Europe”. Pablo M. Olmos also acknowledges the support by BBVA Foundation (Beca Leonardo a Investigadores y Creadores Culturales 2024) and by the Comunidad de Madrid under grants IND2022/TIC-23550 and ELLIS Unit Madrid. David Mitchell acknowledges support from the National Science Foundation under Grant No. CCF-2145917 and travel support from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 951847. Grace Villacrés acknowledges support from Comunidad de Madrid within the 2023-2026 agreement with Universidad Rey Juan Carlos for the granting of direct subsidies for the promotion and encouragement of research and technology transfer, Line of Action A: Emerging Doctors, under Project OrdeNGN.
Appendix A Architecture
In this section, we detail the architecture used to obtain the experimental results with FMNIST and MNIST (28x28 gray-scale images). Note that across experiments we only modify the output layer of the encoder and the input layer of the decoder to adapt to the different configurations of the model. This modification leads to a minimal alteration in the total number of parameters. In Section H.2, we conduct an ablation study on the number of trainable parameters to show that the enhancement in performance is not attributed to the increased dimensionality introduced by redundancy. For the additional CIFAR10 experiments, we change the input of the encoder and the output of the decoder to process the 32x32 color images. For the Tiny ImageNet experiments, we do the same to process 64x64 color images. The rest of the architecture remains unchanged.
A.1 Encoder
The encoder NN consists of 3 convolutional layers followed by two fully connected layers. We employed Leaky ReLU as the intermediate activation function and a Sigmoid as the output activation, as the encoder outputs bit probabilities. The full architecture is detailed in Fig. 8.
Figure 8: Block diagram of the encoder architecture for FMNIST and MNIST.
A.2 Decoder
The decoder architecture is inspired by the one proposed in [24]. It is composed of two fully connected layers, followed by transposed convolutional layers with residual connections and Squeeze-and-Excitation (SE) layers [10]. We employed Leaky ReLU as the intermediate activation function and a Sigmoid as output activation, given that we consider datasets with gray-scale images. The complete architecture is detailed in Fig. 9.
Figure 9: Block diagram of the decoder architecture for FMNIST and MNIST.
Appendix B FMNIST results
In this section, we present supplementary results obtained with the FMNIST dataset.
B.1 Training
We present the evolution of the ELBO and its terms throughout the training process. The models were trained for 200 epochs using an Adam optimizer with a learning rate of $10^{-4}$, and a batch size of 128. Fig. 10 displays the results for configurations with 5 information bits, Fig. 11 for 8 information bits, and Fig. 12 for 10 information bits. The colors in all plots represent the various code rates.
Across all cases, coded models achieve superior bounds. The main differences in the ELBO come from the different performances in reconstruction. As we have observed in the different experiments, coded models are capable of generating more detailed images and accurate reconstructions. Introducing the repetition code also leads to smaller KL values, indicating that the posterior latent features are disentangled and less correlated.
We observe that, as we decrease the code rate, we obtain better bounds in general. Adding redundancy does not increase the model’s flexibility, since the information bits determine the number of latent vectors. However, the introduction of ECCs in the model allows for latent spaces that better capture the structure of the images while employing the same number of latent vectors.
Figure 10: Evolution of the ELBO during training with 5 information bits on FMNIST.
Figure 11: Evolution of the ELBO during training with 8 information bits on FMNIST.
Figure 12: Evolution of the ELBO during training with 10 information bits on FMNIST.
B.2 Reconstruction and generation
In this section, we augment the results presented in the main text, including outcomes obtained with the training dataset in Table 1. We include again the results obtained in the test to facilitate comparison. The results remain consistent across the two data partitions, and the analysis conducted for the test set also applies to training data.
In all the cases, the coded models yield higher PSNR values than their uncoded counterparts, indicating a superior performance in reconstruction. We observe a general improvement in PSNR as we increase the number of information bits (i.e., as we augment the latent dimensionality of the model) and decrease the code rate (i.e., as we introduce more redundancy).
As we discussed in the main text, the PSNR does not account for the semantic errors committed by the model. Therefore, we additionally report the semantic accuracy and the confident semantic accuracy. While the reconstruction accuracy is computed across the entire dataset partitions, for the confident accuracy, we only consider those images projected into a latent vector with a probability exceeding $0.4$. We observe that coded models better capture the semantics of the images while employing the same number of latent vectors, significantly outperforming the uncoded models in terms of accuracy in all the cases.
In Fig 13 we include additional examples of randomly generated images using different model configurations. We observe that coded models can generate more detailed and diverse images than their uncoded counterparts.
Figure 13: Example of randomly generated, uncurated images using different model configurations.
Table 1: Evaluation of reconstruction performance in FMNIST.
Model
PSNR
Acc
Conf. Acc.
PSNR
Acc
Conf. Acc.
uncoded 5 14.490 0.541 0.541 14.477 0.536 0.536
coded 5/50 16.375 0.656 0.702 16.241 0.647 0.700
coded 5/80 16.824 0.694 0.751 16.624 0.688 0.748
coded 5/100 17.001 0.708 0.760 16.702 0.700 0.757
uncoded 8 15.644 0.601 0.602 15.598 0.594 0.595
coded 8/80 17.877 0.769 0.842 17.318 0.750 0.816
coded 8/160 18.828 0.807 0.878 17.713 0.783 0.831
coded 8/240 19.345 0.831 0.921 17.861 0.799 $\mathbf{0.893}$
uncoded 10 16.053 0.650 0.652 16.000 0.644 0.648
coded 10/100 18.827 0.813 0.885 17.694 0.790 0.850
coded 10/200 19.937 0.846 0.897 18.009 0.814 0.871
coded 10/300 $\mathbf{20.529}$ $\mathbf{0.855}$ $\mathbf{0.907}$ $\mathbf{18.111}$ $\mathbf{0.817}$ 0.870
Appendix C MNIST results
In this section, we report the results obtained with the MNIST dataset.
C.1 Training
We present the evolution of the ELBO and its terms throughout the training process. The models were trained for 100 epochs using an Adam optimizer with a learning rate of $10^{-4}$, and a batch size of 128. Fig. 14 displays the results for configurations with 5 information bits, Fig. 15 for 8 information bits, and Fig. 16 for 10 information bits. The colors in all plots represent the various code rates.
The results are consistent with the ones obtained for FMNIST. Across all the configurations, coded models achieve superior bounds. The main differences in the ELBO come from the different performances in reconstruction. As we have observed across the different experiments, coded models are capable of better capturing the structure of the data, generating more detailed images and accurate reconstructions.
We observe that, as we decrease the code rate, we obtain better bounds in general. Adding redundancy does not increase the model’s flexibility, since the information bits determine the number of latent vectors. However, the introduction of ECCs in the model allows for latent spaces that better capture the structure of the images while employing the same number of latent vectors.
Figure 14: Evolution of the ELBO during training with 5 information bits on MNIST.
Figure 15: Evolution of the ELBO during training with 8 information bits on MNIST.
Figure 16: Evolution of the ELBO during training with 10 information bits on MNIST.
C.2 Reconstruction
We first evaluate the model’s performance in reconstructing data by examining its uncoded and coded versions across different configurations, varying the number of information bits and code rates. All the results obtained with MNIST are consistent with those presented in the main text for FMNIST.
In Table 2 we quantify the quality of the reconstructions measuring the PSNR in both training and test sets. In all the cases, coded models yield higher PSNR values, indicating a superior performance in reconstruction. This improvement is also evident through visual inspection of Fig. 17, where the coded models better capture the details in the images. As in FMNIST, we observe a general improvement of the PSNR as we increase the number of information bits and decrease the code rate.
As we discussed in the main text, the PSNR does not account for the semantic errors committed by the model. Therefore, we additionally evaluate the reconstruction accuracy, ensuring that the model successfully reconstructs images within the same class as the original ones. We also provide a confident reconstruction accuracy, for which we do not count errors when the MAP value of $q(\boldsymbol{m}|\boldsymbol{x})$ is below 0.4. In light of the results, we argue that introducing ECCs in the model allows for latent spaces that better capture the semantics of the images while employing the same number of latent vectors, outperforming the uncoded models in all the cases.
Figure 17: Example of reconstructed test images obtained with different model configurations. Observe that more details are visualized as we increase bits in the latent space and decrease the coding rate.
Figure 18: Example of erroneous reconstructions in MNIST using the 4 most-probable words (a posterior). The first column in each image shows the original input to the model, while the second column displays the reconstructions. The a posteriori probability of the word used for reconstruction is indicated in each row.
Table 2: Evaluation of reconstruction performance in MNIST.
Model
PSNR
Acc
Conf. Acc.
PSNR
Acc
Conf. Acc.
Entropy
uncoded 5 13.483 0.702 0.703 13.483 0.701 0.702 0.277
coded 5/50 14.983 0.887 0.923 14.888 0.887 0.920 2.073
coded 5/80 15.436 0.899 0.936 15.263 0.895 0.929 2.237
coded 5/100 15.590 0.905 0.931 15.352 0.898 0.924 2.382
uncoded 8 14.530 0.860 0.864 14.490 0.860 0.868 0.513
coded 8/80 16.878 0.937 0.964 16.042 0.912 0.947 3.105
coded 8/160 18.108 0.957 0.974 16.497 0.927 0.951 3.645
coded 8/240 19.984 0.967 0.978 16.688 0.936 0.957 3.881
uncoded 10 14.879 0.888 0.891 14.816 0.887 0.890 0.636
coded 10/100 17.584 0.945 0.972 16.795 0.928 $\mathbf{0.968}$ 4.080
coded 10/200 20.060 0.973 0.977 16.863 0.932 0.944 4.411
coded 10/300 $\mathbf{21.083}$ $\mathbf{0.979}$ $\mathbf{0.984}$ $\mathbf{17.114}$ $\mathbf{0.941}$ 0.945 4.810
We also report the average entropy of the variational posterior over the test set in Table 2. If we analyze the entropy together with the semantic accuracy, we can argue that coded VAE is aware that multiple vectors might be related to the same image class, and that the posterior shows larger uncertainties for images for which the model has not properly identified the class. We illustrate this argument in Fig. 18, where we show some images selected so that the MAP latent word of $q(\boldsymbol{m}|\boldsymbol{x})$ induces class reconstruction errors. We show the reconstruction of the 4 most probable latent vectors and their corresponding probabilities. Observe that the uncoded model is confident no matter the reconstruction outcome, while in the coded posterior, the uncertainty is much larger.
Table 4 shows the log-likelihood values obtained for the MNIST dataset with various model configurations. Coded models consistently outperform their uncoded counterparts for both the training and test sets, consistent with the findings observed using the FMNIST dataset.
C.3 Generation
In this section, we evaluate the model in the image generation task. In Fig. 19, we show examples of randomly generated images using different model configurations in MNIST. These results are consistent with the ones obtained in reconstruction, and with the ones obtained for FMNIST, as we observe that the coded models can generate more detailed and diverse images.
The improved inference provided by the repetition code can also be tested by generating images using the generative model and counting errors using the MAP solution of the variational posterior distribution. Table 4 reports the BER and WER for MNIST. Remarkably, for the same number of latent bits, coded models reduce both the BER and WER w.r.t. the uncoded case. Note also that the error rates grow with the number of latent bits, but this is expected due to the increased complexity of the inference process.
Figure 19: Example of randomly generated, uncurated images using different model configurations.
Table 3: Evaluation of the BER, WER in MNIST.
Model BER WER
uncoded 5 0.002 0.008
coded 5/50 0.007 0.034
coded 5/80 0.004 0.021
coded 5/100 0.009 0.045
uncoded 8 0.015 0.071
coded 8/80 0.020 0.147
coded 8/160 0.021 0.160
coded 8/240 0.023 0.167
uncoded 10 0.057 0.373
coded 10/100 0.030 0.258
coded 10/200 0.034 0.282
coded 10/300 0.041 0.331
Table 4: Evaluation of the log-likelihood (LL) in MNIST.
Model LL (train) LL (test)
uncoded 5 -149.049 -148.997
coded 5/50 -117.979 -119.094
coded 5/80 -$1$116.639
coded 5/100 -115.189 -117.200
uncoded 8 -127.079 -127.555
coded 8/80 -96.554 -104.692
coded 8/160 -96.014 -107.436
coded 8/240 -97.316 -111.312
uncoded 10 -120.594 -121.332
coded 10/100 -92.545 -99.373
coded 10/200 -86.072 -106.249
coded 10/300 -88.904 -110.799
Appendix D CIFAR10 results
In this section, we provide additional results using the CIFAR10 dataset with different model configurations.
D.1 Training
We present the evolution of the ELBO and its terms throughout the training process. The models were trained for 300 epochs using Adam optimizer with a learning rate of $10^{-4}$, and a batch size of 128. Fig. 20 displays the results for configurations with 70 information bits, Fig. 21 for 100 information bits, and Fig. 22 for 130 information bits. The colors in all plots represent the various code rates.
Across all the configurations, coded models achieve superior bounds. The main differences in the ELBO come from the different performances in reconstruction. As we have observed across the different experiments, coded models are capable of better capturing the structure of the data, generating more detailed images and accurate reconstructions.
Figure 20: Evolution of the ELBO during training with 70 information bits on CIFAR10.
Figure 21: Evolution of the ELBO during training with 100 information bits on CIFAR10.
Figure 22: Evolution of the ELBO during training with 130 information bits on CIFAR10.
In the coded case, we do not observe significant differences in the obtained bounds as we increase the number of information bits and reduce the code rate. However, the difference is notable if we compare the coded and uncoded models. Adding redundancy does not increase the model’s flexibility, since the information bits determine the number of latent vectors. However, the introduction of ECCs in the model allows for latent spaces that better capture the structure of the images while employing the same number of latent vectors.
It’s important to note that we are currently using feed-forward networks at the decoder’s input. However, this approach may not be suitable for the correlations present in our coded words. Utilizing an architecture capable of effectively leveraging these correlations among the coded bits could potentially enable us to better exploit the introduced redundancy.
D.2 Reconstruction and generation
We first evaluate the model’s performance in reconstructing data by examining its uncoded and coded versions across different configurations, varying the number of information bits and code rates.
In Table 6 we quantify the quality of the reconstructions measuring the PSNR in both training and test sets. Coded models yield higher PSNR values in train, and similar values in test, although the coded models with lower rates outperform the rest of the configurations. However, the improvement in reconstruction is evident through visual inspection of Fig. 23, where the coded models better capture the details in the images. We observe that the coded model yields images that better resemble the structure of the dataset, while the uncoded DVAE cannot decouple spatial information from the images and project it in the latent space.
We hypothesize that to adequately model complex images, transitioning to a hierarchical structure may be necessary. This would allow for the explicit modeling of both global and local information. However, despite employing this rather simple model, we observe that coded configurations outperform their uncoded counterparts in capturing colors and textures.
We also evaluate the model in the image generation task. In Fig. 24, we show examples of randomly generated images using different model configurations in CIFAR10. These results are consistent with the ones obtained in reconstruction, as we observe that the coded models can generate more detailed and diverse images. Additionally, we obtained the Fréchet Inception Distance (FID) score using the test set and 10k generated samples. For this, we used the implementation available at https://github.com/mseitzer/pytorch-fid. We can observe that the coded models significantly reduced the FID score in all the cases compared to their uncoded counterparts.For the coded models, we do not observe a clear influence of the code rate on the quality of the generations.
Figure 23: Example of reconstructed test images obtained with different model configurations. Observe that more details are visualized as we increase bits in the latent space and introduce redundancy.
Figure 24: Example of randomly generated, uncurated images using different model configurations.
Table 5: Evaluation of reconstruction performance in CIFAR10 with different model configurations.
Model PSNR (train) PSNR (test)
uncoded 70 17.985 17.596
coded 70/700 23.790 17.731
coded 70/1400 24.555 18.008
coded 70/2100 25.551 18.401
uncoded 100 18.509 18.334
coded 100/1000 24.754 18.229
coded 100/2000 24.866 18.927
coded 100/3000 25.646 18.920
uncoded 130 18.951 18.758
coded 130/1300 25.007 18.887
coded 130/2600 25.460 19.416
coded 130/3900 $\mathbf{25.515}$ $\mathbf{19.292}$
Table 6: Evaluation of the BER, WER, and FID in CIFAR10 with different model configurations.
Model BER WER FID
uncoded 70 0.162 1.000 177.524
coded 70/700 0.101 1.000 104.977
coded 70/1400 0.088 0.999 104.078
coded 70/2100 0.090 0.999 102.795
uncoded 100 0.182 1.000 172.063
coded 100/1000 0.123 1.000 107.887
coded 100/2000 0.114 1.000 101.182
coded 100/3000 0.138 1.000 107.287
uncoded 130 0.197 1.000 164.138
coded 130/1300 0.144 1.000 109.905
coded 130/2600 0.164 1.000 110.250
coded 130/3900 0.185 1.000 108.561
Appendix E Tiny ImageNet results
In this section, we provide additional results using the Tiny ImageNet dataset with different model configurations.
E.1 Training
We present the evolution of the ELBO and its terms throughout the training process. The models were trained for 300 epochs using an Adam optimizer with a learning rate of $10^{-4}$, and a batch size of 128. Fig. 25 displays the results for configurations with 70 information bits, Fig. 26 for 100 information bits, and Fig. 27 for 130 information bits. The colors in all plots represent the various code rates.
As in the rest of the datasets, coded models achieve superior bounds across all the configurations. The main differences in the ELBO come from the different performances in reconstruction. As we have observed across the different experiments, coded models are capable of better capturing the structure of the data, generating more detailed images and accurate reconstructions.
In the coded case, we do not observe significant differences in the obtained bounds as we increase the number of information bits and reduce the code rate. However, the difference is notable if we compare the coded and uncoded models. Adding redundancy does not increase the model’s flexibility, since the information bits determine the number of latent vectors. However, the introduction of ECCs in the model allows for latent spaces that better capture the structure of the images while employing the same number of latent vectors.
Figure 25: Evolution of the ELBO during training for the configurations with 70 information bits.
Figure 26: Evolution of the ELBO during training for the configurations with 100 information bits.
Figure 27: Evolution of the ELBO during training for the configurations with 130 information bits.
It’s important to note that we are currently using feed-forward networks at the decoder’s input. However, this approach may not be suitable for the correlations present in our coded words. Utilizing an architecture capable of effectively leveraging these correlations among the coded bits could potentially enable us to better exploit the introduced redundancy.
E.2 Reconstruction and generation
We first evaluate the model’s performance in reconstructing data by examining its uncoded and coded versions across different configurations, varying the number of information bits and code rates.
In Table 8 we quantify the quality of the reconstructions measuring the PSNR in both training and test sets. Coded models yield higher PSNR values in train, and similar values in test, although the coded models with lower rates outperform the rest of the configurations. However, the improvement in reconstruction is evident through visual inspection of Fig. 28, where the coded models better capture the details in the images. We observe that the coded model yields images that better resemble the structure of the dataset, while the uncoded DVAE cannot decouple spatial information from the images and project it in the latent space.
We hypothesize that to adequately model complex images, transitioning to a hierarchical structure may be necessary. This would allow for the explicit modeling of both global and local information. However, despite employing this rather simple model, we observe that coded configurations outperform their uncoded counterparts in capturing colors and textures.
Figure 28: Example of reconstructed test images obtained with different model configurations. Observe that more details are visualized as we increase the bits in the latent space and introduce redundancy.
Figure 29: Example of randomly generated, uncurated images using different model configurations.
Table 7: Evaluation of reconstruction performance in Tiny ImageNet with different model configurations.
Model PSNR (train) PSNR (test)
uncoded 70 15.598 15.402
coded 70/700 18.156 15.158
coded 70/1400 18.396 15.419
coded 70/2100 18.228 15.789
uncoded 100 16.012 15.774
coded 100/1000 18.298 15.677
coded 100/2000 18.647 15.892
coded 100/3000 18.729 16.167
uncoded 130 16.278 16.009
coded 130/1300 18.719 15.901
coded 130/2600 18.818 $\mathbf{16.329}$
coded 130/3900 $\mathbf{19.020}$ 16.288
Table 8: Evaluation of the BER, WER, and FID in Tiny ImageNet with different model configurations.
Model BER WER FID
uncoded 70 0.143 1.000 265.474
coded 70/700 0.096 0.998 171.993
coded 70/1400 0.104 1.000 170.496
coded 70/2100 0.096 0.998 176.245
uncoded 100 0.164 1.000 234.358
coded 100/1000 0.099 1.000 153.743
coded 100/2000 0.097 1.000 162.889
coded 100/3000 0.098 1.000 163.049
uncoded 130 0.200 1.000 219.003
coded 130/1300 0.129 1.000 165.064
coded 130/2600 0.114 1.000 164.759
coded 130/3900 0.128 1.000 170.603
We also evaluate the model in the image generation task. In Fig. 29, we show examples of randomly generated images using different model configurations in Tiny ImageNet. These results are consistent with the ones obtained in reconstruction, as we observe that the coded models can generate more detailed and diverse images. Additionally, we obtained the FID score using the test set and 10k generated samples. For this, we used the implementation available at https://github.com/mseitzer/pytorch-fid. We can observe that the coded models significantly reduced the FID score in all the cases compared to their uncoded counterparts.
For the coded models, we do not observe a clear influence of the code rate on the quality of the generations in CIFAR-10. However, in Tiny ImageNet, smaller code rates produce worse FID scores. We hypothesize that this may be due to the presence of artifacts in the generated images. Our experiments indicate that coded models with lower rates attempt to model fine details in images, which can lead to artifacts in generation.
Appendix F IWAE results
One could draw a parallel between the coded DVAE with repetition codes and the well-known IWAE [3], but the two approaches are fundamentally different. In the IWAE, independent samples are drawn from the variational posterior and propagated independently through the generative model to obtain a tighter variational bound on the marginal log-likelihood. In our method, we jointly propagate the output of the ECC encoder through the generative model, obtaining a single prediction and exploiting the introduced known correlations in the variational approximation of the posterior. In the case of repetition codes, the ECC encoder outputs are repeated bits, or repeated probabilities in the case of soft encoding. However, our approach extends beyond repetition codes, opening a new field for improved inference in discrete latent variable models.
In this work, we specifically utilize the redundancy introduced by the repetition code to correct potential errors made by the encoder through a soft decoding approach, leading to a more accurate approximation of $p(\boldsymbol{m}|\boldsymbol{x})$ and an improved proposal for sampling. The results obtained in the coded DVAE case cannot be achieved by training the uncoded DVAE with the IWAE objective. The following results compare the uncoded IWAE model with the coded DVAE trained on FMNIST. While the uncoded model shows slight performance gains with an increasing number of IWAE samples (which improves the evidence lower bound), it still underperforms compared to the coded model. Furthermore, when using 20 and 30 IWAE samples, the metrics slightly declined compared to using 10 samples, likely due to overfitting, as we applied a common early stopping point.
Figure 30: Evolution of the ELBO and the IWAE objectives for various configurations. Observe that the IWAE provides a tighter bound than the ELBO in the uncoded setting. However, coded models obtain even better bounds using the same number of samples/repetitions.
Table 9: Comparison of the metrics obtained with our method and the uncoded DVAE trained with the IWAE objective.
Model BER WER Entropy Acc. Conf. Acc. PSNR
uncoded 8 0.089 0.384 0.467 0.594 0.595 15.598
uncoded 8 IWAE 10 samples 0.063 0.372 1.309 0.617 0.640 14.282
uncoded 8 IWAE 20 samples 0.075 0.447 1.391 0.634 0.651 14.237
uncoded 8 IWAE 30 samples 0.074 0.438 1.564 0.619 0.641 13.757
coded 8/80 0.021 0.144 2.905 0.750 0.816 17.318
coded 8/160 0.027 0.189 3.637 0.783 0.831 17.713
coded 8/240 0.037 0.231 4.000 0.799 0.893 17.861
Appendix G Hierarchical coded DVAE results
Inspired by polar codes [2], we present a hierarchical coded DVAE with two layers of latent bits, as illustrated in Figure 31. In this model, the latent bits $\boldsymbol{m}_{1}$ are encoded using a repetition code in the first layer, producing $\boldsymbol{c}_{1}$ and $\boldsymbol{z}_{1}$. Concurrently, the bits in the second layer, $\boldsymbol{m}_{2}$, are linearly combined with $\boldsymbol{m}_{1}$ following $\boldsymbol{m}_{1,2}=\boldsymbol{m}_{1}\oplus\boldsymbol{m}_{2}$, considering a binary field or Galois field. The resulting vector is then encoded with another repetition code to produce $\boldsymbol{c}_{2}$, which is subsequently modulated into $\boldsymbol{z}_{2}$. Finally, both $\boldsymbol{z}_{1}$ and $\boldsymbol{z}_{2}$ are concatenated and passed through the decoder network to generate $\boldsymbol{x}$. The model provides stronger protection for $\boldsymbol{m}_{1}$, as it appears in both branches of the generative model. Inference follows a similar approach to the one employed in the coded DVAE, incorporating the linear combination of $\boldsymbol{m}_{1}$ and $\boldsymbol{m}_{2}$ used in the second branch.
Figure 31: Graphical model of the hierarchical Coded VAE with two layers.
We adopt the same variational family as used in the standard Coded VAE; however, in this case, we incorporate both hierarchical levels, leading to
$q_{\boldsymbol{\eta}}(\boldsymbol{m},\boldsymbol{z}|\boldsymbol{x})=q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1}|\boldsymbol{x})q_{\boldsymbol{\eta}}(\boldsymbol{m}_{2}|\boldsymbol{x})p(\boldsymbol{z}_{1}|\boldsymbol{c}_{1})p(\boldsymbol{z}_{2}|\boldsymbol{c}_{2}),$ (12)
where $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1}|\boldsymbol{x})$ is calculated following the same approach as in the coded DVAE with repetition codes, computing the all-are-zero and all-are-ones products of probabilities of the bits in $\boldsymbol{c}_{1}$ that are copies of the same message bit. The posterior $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{2}|\boldsymbol{x})$ considering both the encoder’s output and the inferred posterior distribution $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1}|\boldsymbol{x})$. Note that in this case, the decoder outputs the probabilities for both $\boldsymbol{c}_{1}$ and $\boldsymbol{c}_{2}$, with $\boldsymbol{c}_{2}$ being the encoded version of the linear combination $\boldsymbol{m}_{1,2}=\boldsymbol{m}_{1}\oplus\boldsymbol{m}_{2}$. Consequently, we first obtain $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1,2}|\boldsymbol{x})$ following the same approach as in the coded DVAE with repetition codes, and determine $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{2}|\boldsymbol{x})$ as
$q_{\boldsymbol{\eta}}(\boldsymbol{m}_{2}|\boldsymbol{x})=\prod_{u=1}^{M}\text{Ber}(p_{u}),$ (13)
$p_{u}=q_{\boldsymbol{\eta}}(m_{1,2,u}=1|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{1,u}=0|\boldsymbol{x})+q_{\boldsymbol{\eta}}(m_{1,2,u}=0|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{1,u}=1|\boldsymbol{x}).$ (14)
After obtaining $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1}|\boldsymbol{x})$ and $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{2}|\boldsymbol{x})$, we recalculate the posterior bit probabilities for the linear combination $q^{\prime}_{\boldsymbol{\eta}}(\boldsymbol{m}_{1,2}|\boldsymbol{x})$ as
$q^{\prime}_{\boldsymbol{\eta}}(\boldsymbol{m}_{1,2}|\boldsymbol{x})=\prod_{u=1}^{M}\text{Ber}(q_{u}),$ (15)
$q_{u}=q_{\boldsymbol{\eta}}(m_{1,u}=1|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{2,u}=0|\boldsymbol{x})+q_{\boldsymbol{\eta}}(m_{1,u}=0|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{2,u}=1|\boldsymbol{x}).$ (16)
Next, we apply the soft encoding approach to incorporate the repetition code structure at both levels of the hierarchy.The posterior probabilities $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1}|\boldsymbol{x})$ are repeated to obtain $q_{\boldsymbol{\eta}}(\boldsymbol{c}_{1}|\boldsymbol{x})$, and the posterior probabilities $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1,2}|\boldsymbol{x})$ are repeated to produce $q_{\boldsymbol{\eta}}(\boldsymbol{c}_{2}|\boldsymbol{x})$. Utilizing the reparameterization trick from Eq. 5, we sample $\boldsymbol{z}_{1}$ and $\boldsymbol{z}_{2}$, concatenate them to form $\boldsymbol{z}$, and pass this through the decoder to generate $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})$. The model is trained by maximizing the ELBO, given by
$\text{ELBO}=\mathbb{E}_{q_{\boldsymbol{\eta}}(\boldsymbol{m},\boldsymbol{z}|\boldsymbol{x})}\log p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})-\mathcal{D}_{KL}\big{(}q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1}|\boldsymbol{x})||p(\boldsymbol{m}_{1})\big{)}-\mathcal{D}_{KL}\big{(}q_{\boldsymbol{\eta}}(\boldsymbol{m}_{2}|\boldsymbol{x})||p(\boldsymbol{m}_{2})\big{)},$ (17)
where both $p(\boldsymbol{m}_{1})$ and $p(\boldsymbol{m}_{2})$ are assumed to be independent Bernoulli distributions with bit probabilities of $0.5$, consistent with the other scenarios.
We obtained results on the FMNIST dataset using a model with 5 information bits per branch and repetition rates of $R=1/10$ and $R=1/20$. In this case, we applied the same code rate to both branches, although varying code rates could be used to control the level of protection at each hierarchy level. Tables 10 and 11 present the metrics obtained for the different configurations. Specifically, Table 10 shows the overall metrics obtained with this structure, and Table 11 compares the error metrics across the two hierarchy levels. As expected, $\mathbf{m}_{2}$ shows poorer error metrics compared to $\mathbf{m}_{1}$, since the model provides more redundancy to $\mathbf{m}_{1}$ incorporating it in both branches. Although the overall metrics and generation quality are somewhat similar to those of the coded DVAE with 10 information bits (see tables in Figures 2 and 4), the introduced hierarchy results in a more interpretable latent space. In this setup, $\mathbf{m}_{1}$ captures global features (such as clothing types in the FMNIST dataset), while $\mathbf{m}_{2}$ controls individual features, as we can observe in Figure 32, where we show examples of the model’s generative outputs for fixed $\mathbf{m}_{1}$ and random samples of $\mathbf{m}_{2}$.
Figure 32: Examples of generated images using the hierarchical coded DVAE with (a) a 5/50 repetition code in each branch, and (b) a 5/100 repetition code in each branch. In all the examples provided, $\boldsymbol{m}_{1}$ was fixed while $\boldsymbol{m}_{2}$ was randomly sampled.
Table 10: Comparison of the obtained metrics for the coded DVAE with polar codes with different configurations, which we refer to as ’hierarchical coded DVAE’.
Model BER WER Acc Conf. Acc PSNR
hier. 5/50 0.099 0.400 0.753 0.800 17.130
hier. 5/100 0.050 0.330 0.784 0.870 17.513
Table 11: Comparison of the obtained error metrics in the different hierarchy levels.
Model BER $\boldsymbol{m}_{1}$ WER $\boldsymbol{m}_{1}$ BER $\boldsymbol{m}_{2}$ WER $\boldsymbol{m}_{2}$
hier. 5/50 0.079 0.259 0.119 0.362
hier. 5/100 0.026 0.110 0.075 0.287
Appendix H Ablation study
In this section, we conduct ablation studies on the hyperparameter $\beta$ of the model, responsible for regulating the decay of exponentials in the smoothing transformation, as well as on the number of trainable parameters in the models.
H.1 Ablation study on the hyperparameter $\beta$
Across all experiments, we have consistently configured the hyperparameter $\beta$, which controls the decay of exponentials in the smoothing transformation, to a value of 15. To illustrate its impact on the overall performance of the model, we conducted an ablation study on the value of this hyperparameter for both uncoded and coded cases.
The smoothing distribution employed for the reparameterization trick consists of two overlapping exponentials. The hyperparameter $\beta$ functions as a temperature term, regulating the decay of the distributions and, consequently, influencing the degree of overlapping. A lower $\beta$ value results in more overlapped tails, while a higher value leads to less overlapped distributions. A priori, we would like these distributions to be separated, allowing us to retrieve the true value of the bit and effectively use the latent structure of the model.
H.1.1 Coded model
We first evaluate the influence of the parameter $\beta$ in coded models. We take as a reference the coded model with 8 information bits and a rate $R=1/30$, and train it using $\beta=5,10,15,20$. We assess the performance of the model in reconstruction and generation tasks. We observe the model is fairly robust, achieving similar performance across configurations in most metrics.
In Fig. 33 we show examples of reconstructed images using the different configurations to assess reconstruction through visual examination, and Table 12 contains the associated reconstruction metrics. All the configurations achieve similar performances, although the models trained with $\beta=10$ and $\beta=15$ seem to be the best configurations for this scenario. Larger values may result in unstable training and inferior performance.
Figure 33: Example of reconstructed images obtained with different values of $\beta$ using the coded model with an 8/240 code.
Table 12: Evaluation of reconstruction performance in FMNIST with different values of $\beta$ using the coded model with an 8/240 code.
Model
PSNR
Acc
Conf. Acc.
PSNR
Acc
Conf. Acc.
Entropy
5 18.344 0.791 0.895 17.614 0.766 0.849 4.025
10 19.106 0.822 0.904 17.737 0.793 0.872 4.023
15 19.345 0.831 0.921 17.861 0.799 0.893 4.000
20 18.797 0.809 0.887 17.837 0.787 0.877 3.810
Figure 34: Example of randomly generated, uncurated images with different values of $\beta$ using the coded model with an 8/240 code.
Table 13: Evaluation of the BER and WER in FMNIST with different values of $\beta$ using the coded model with an 8/240 code.
Beta BER WER
5 0.150 0.726
10 0.080 0.480
15 0.037 0.231
20 0.065 0.399
Table 14: Evaluation of the log-likelihood (LL) in FMNIST with different values of $\beta$ using the coded model with an 8/240 code.
Beta LL (train) LL (test)
5 -228.448 -234.629
10 -229.379 -237.495
15 -231.679 -238.459
20 -229.627 -235.927
Next, we evaluate the model in the image generation task. Fig. 34 contains examples of randomly generated images using the different configurations. Table 14 reports the obtained BER and WER, and Table 14 the estimated log-likelihood of the different values of $\beta$. The model trained with $\beta=15$ stands out in terms of error metrics, although achieves similar log-likelihood values as the model trained with $\beta=10$. Again, these two configurations appear to be the most suitable in this scenario.
H.1.2 Uncoded model
We first evaluate the influence of the parameter $\beta$ in uncoded models. We take as a reference the coded model with 8 information bits and train it using $\beta=5,10,15,20$. We assess the performance of the model in reconstruction and generation tasks. We observe that the uncoded model is also robust, achieving similar performance across configurations.
In Fig. 35 we show examples of reconstructed images using the different configurations to assess reconstruction through visual examination, and Table 15 contains the associated reconstruction metrics. All the configurations achieve similar performances, although the models trained with $\beta=10$ and $\beta=15$ seem to be the best configurations for this scenario. Larger values may result in unstable training and inferior performance, as we can clearly observe in this case.
Figure 35: Example of reconstructed images obtained with different values of $\beta$ using an uncoded model 8 information bits.
Figure 36: Example of randomly generated, uncurated images with different values of $\beta$ using an uncoded model 8 information bits.
Table 15: Evaluation of reconstruction performance in FMNIST with different values of $\beta$ using an uncoded model 8 information bits.
Model
PSNR
Acc
Conf. Acc.
PSNR
Acc
Conf. Acc.
Entropy
5 14.239 0.503 0.501 14.237 0.503 0.491 0.231
10 15.624 0.606 0.603 15.571 0.598 0.598 0.357
15 15.644 0.601 0.602 15.598 0.594 0.595 0.467
20 13.717 0.464 0.466 13.743 0.460 0.462 0.383
Table 16: Evaluation of the BER and WER in FMNIST with different values of $\beta$ using an uncoded model 8 information bits.
Beta BER WER
5 0.203 0.852
10 0.086 0.384
15 0.089 0.384
20 0.278 0.939
Table 17: Evaluation of the log-likelihood (LL) in FMNIST with different values of $\beta$ using an uncoded model 8 information bits.
Model LL (train) LL (test)
5 -256.431 -257.983
10 -247.507 -249.460
15 -247.964 -249.880
20 -272.460 -273.554
Next, we evaluate the model in the image generation task. Fig. 36 contains examples of randomly generated images using the different configurations. Table 17 reports the obtained BER and WER, and Table 17 the estimated log-likelihood of the different values of $\beta$. The models trained with $\beta=15$ and $\beta=10$ clearly outperform the other two in this task, generating more diverse and detailed images, and obtaining better error metrics and log-likelihood values.
H.2 Ablation study on the number of trainable parameters
A consistent architecture was employed across all experiments, which is detailed in Section A. However, since the introduction of the code alters the dimensionality of the latent space, it is necessary to adjust the encoder’s output and the decoder’s input. This results in an augmentation of the trainable parameters in the coded cases compared to their uncoded counterparts.
Given that a higher number of parameters usually results in better performance, we conducted an ablation study on the model’s trainable parameters to confirm that the improved performance introduced by the coded models is not due to this factor. We adjusted the hidden dimensions of the encoder and decoder architectures to ensure both configurations (coded and uncoded) have roughly the same number of trainable parameters. We have conducted the ablation study using the uncoded model with 8 bits and the coded 8/240 model trained on FMNIST.
We adjusted the encoder’s last hidden dimension and the decoder’s first hidden dimension to equalize the parameter count between the uncoded and coded models. This adjustment was straightforward since the last layers of the encoder and the first layers of the decoder are feed-forward layers. We kept the latent dimension of the model unchanged, ensuring that the modification solely pertained to the neural network architecture.
Table 18: Parameter count.
Model # encoder parameters # decoder parameters
uncoded 8 6,592,008 19,341,185
uncoded 8 adjusted 6,717,538 19,581,035
coded 8/240 6,711,024 19,578,753
coded 8/240 adjusted 6,583,174 19,332,871
H.2.1 Evaluation
This section provides an empirical evaluation of the models trained with the adjusted parameter count, demonstrating that the enhanced performance observed in coded models does not result from an augmented number of trainable parameters. We found that the performance of the original and adjusted models is very similar, meaning that the conclusions drawn in the main text hold even in this scenario.
We first evaluate the reconstruction performance, measuring the PSNR and reconstruction accuracy in both train and test sets, which are included in Table 19. Then, we assess generation measuring the BER and WER, reported in Table 21. Finally, we compute the log-likelihood for train and test sets, shown in Table 21.
In terms of reconstruction, both the original and adjusted models exhibit very similar performance, observed in both reconstruction quality and accuracy. However, the adjusted models show slightly inferior results than the original ones. For the coded model, this might be attributed to reduced flexibility when decreasing the number of parameters. As for the uncoded model, the increased complexity while maintaining the same low-dimensional latent vector may not provide enough expressiveness to leverage the added flexibility in the architecture, potentially causing the observed decrease in performance. We observe the same behavior when we evaluate the BER and WER.
Analyzing the results, especially the log-likelihood values shown in Table 21, we can argue that increasing the flexibility in the architecture does not necessarily lead to improved performance in this scenario. The coded models exhibit similar performance with both the original and adjusted parameter counts, consistently outperforming the uncoded models. These results indicate that the performance enhancement is attributed to the introduction of ECCs in the latent space, rather than differences in the architecture required to handle the introduced redundancy.
Figure 37: Example of reconstructed images obtained with different configurations.
Figure 38: Example of randomly generated, uncurated images using different model configurations.
Table 19: Evaluation of reconstruction performance in FMNIST with the adjusted parameter count.
Model
PSNR
Acc
Conf. Acc.
PSNR
Acc
Conf. Acc.
Entropy
uncoded 8 15.644 0.601 0.602 15.598 0.594 0.595 0.659
uncoded 8 adj. 15.530 0.586 0.586 15.491 0.581 0.580 0.449
coded 8/240 19.345 0.831 0.921 17.861 0.799 0.893 4.609
coded 8/240 adj. 19.383 0.828 0.883 17.771 0.792 0.828 3.952
Table 20: Evaluation of the BER and WER in FMNIST with the adjusted parameter count.
Model BER WER
uncoded 8 0.089 0.384
uncoded 8 adj. 0.125 0.561
coded 8/240 0.037 0.231
coded 8/240 adj. 0.064 0.399
Table 21: Evaluation of the log-likelihood (LL) in FMNIST with the adjusted parameter count.
Model LL (train) LL (test)
uncoded 8 -247.964 -249.880
uncoded 8 adj. -250.543 -252.408
coded 8/240 -231.679 -238.459
coded 8/240 adj. -229.283 -238.302
Appendix I Evaluating log-likelihood using the soft-encoding model
The reparameterization trick introduced in (5) requires that bits are independent, a condition not met in $\boldsymbol{c}$ once the code’s structure is introduced. To address this issue, during training, we employ a soft encoding strategy. We assume the bits in $\boldsymbol{c}$ are independent and equally distributed according to $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$. Therefore, instead of directly repeating sampled bits in $\boldsymbol{m}$ to obtain $\boldsymbol{c}$ following $\boldsymbol{c}=\boldsymbol{m}^{T}\mathbf{G}$, we repeat the posterior probabilities for the copies of the same bit and sample $\boldsymbol{z}$ using the reparameterization trick in (5). Remark that, despite the soft encoding assumption during training, the generative results in Fig. 4, 6, 13, 19, 24, 24, and 29 are obtained through hard encoding. Namely, we sampled an information word $\boldsymbol{m}$ and obtained $\boldsymbol{c}$ by repeating its bits.
While the hard-encoding images are visually appealing, to evaluate the coded DVAE LL for a given image $\boldsymbol{x}$, we have to leverage the soft-encoding model since it is unrealistic for a sample from the soft-encoding model to produce equally repeated bits. In the soft-coded model, we sample bit probabilities from a prior distribution that we model through the product of $M$ independent Beta distributions, and we use a proposal distribution model similar to the Vamp-prior in [28]. Namely, a mixture model with components given by $q(\boldsymbol{m}|\boldsymbol{x})$ for different training points.
In the main text and Section H, we report the log-likelihood values for different model configurations trained on the FMNIST dataset. Table 4 presents the results obtained for the MNIST dataset. Due to their simplicity, these datasets do not require high-dimensional latent spaces, and competitive results can be achieved with just 8 or 10 information bits. However, for more complex datasets like CIFAR10 or Tiny ImageNet, high-dimensional latent spaces are necessary to capture spatial information, colors, and textures. For these high-dimensional datasets, we were unable to obtain valid log-likelihood estimates because the number of samples needed for importance sampling to converge was too large. However, given the difference in the level of detail in the reconstructed and generated images between coded and uncoded models (see Sections D and E), coded models are expected to better approximate the true marginal likelihood of the data.
Appendix J Variational inference at codeword level
Here, we present an alternative variational family that is valid for any ECC, including random codes. We assume that we have a deterministic mapping of the form $\boldsymbol{c}=\mathcal{C}(\boldsymbol{m})$. We assume a variational family of the form
$q_{\boldsymbol{\eta}}(\boldsymbol{c},\boldsymbol{z}|\boldsymbol{x})=q_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x})p(\boldsymbol{z}|\boldsymbol{c}),$ (18)
$q_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x})\propto p(\boldsymbol{c})q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}|\boldsymbol{x}),$ (19)
$q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}|\boldsymbol{x})=\prod_{j=1}^{M/R}\text{Ber}(g_{j,\boldsymbol{\eta}}(\boldsymbol{x})),$ (20)
where $g_{\boldsymbol{\eta}}(\boldsymbol{x})$ represents the output of the encoder with a parameter set denoted as $\boldsymbol{\eta}$. Note that $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}|\boldsymbol{x})$ corresponds to the uncoded posterior, which we subsequently constrain using the prior distribution $p(\boldsymbol{c})$ over the code words to obtain the coded posterior $q_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x})$. Then, the coded posterior distribution can be defined as a categorical distribution over the set of code words $\mathcal{C}(\boldsymbol{m})$, which is given by
$\begin{gathered}q_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x})=\text{Cat}\Big{(}\Big{[}\frac{1}{W}q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}_{1}|\boldsymbol{x}),\ldots,\frac{1}{W}q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}_{2^{M}}|\boldsymbol{x})\Big{]}\Big{)}\\ =\frac{1}{W}\prod_{\boldsymbol{c}_{i}\in\mathcal{C}(\boldsymbol{m})}\prod_{j=1}^{M/R}g_{j,\boldsymbol{\eta}}(\boldsymbol{x})^{c_{i,j}}(1-g_{j,\boldsymbol{\eta}}(\boldsymbol{x}))^{(1-c_{i,j})},\end{gathered}$ (21)
where $W=\sum_{\boldsymbol{c}_{i}\in\mathcal{C}(\boldsymbol{m})}q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}_{i}|\boldsymbol{x})$ is a constant for normalization.
Inference is done by maximizing the ELBO, which can be expressed as
[…]
In this case, due to the inability to compute the KL Divergence in closed form, we approximate it via Monte Carlo, sampling from the categorical distribution $q_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x})$. The reconstruction term also needs to be approximated via Monte Carlo. Since the use of channel coding introduces structural dependencies among the components of the vectors $\boldsymbol{c}$, we can no longer assume their independence. Consequently, the formulation of the smoothing transformation as independent mixtures introduced in the DVAE is no longer applicable. Hence, this approach involves sampling $\boldsymbol{c}^{\prime}$ from the categorical distribution $q_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x})$ and subsequently applying the transformation over the sampled word. Thus, we obtain a smooth transformation for each sample $\boldsymbol{c}^{\prime}$ using the inverse CDFs of $p(z_{j}|c_{j}=0)$ and $p(z_{j}|c_{j}=1)$, which are given by
$F^{-1}_{p(z_{j}|\boldsymbol{c}^{\prime}_{j}=0)}(\rho)=-\frac{1}{\beta}\log\big{(}1-\rho(1-e^{-\beta})\big{)},$ (23)
$F^{-1}_{p(z_{j}|\boldsymbol{c}^{\prime}_{j}=1)}(\rho)=\frac{1}{\beta}\log\big{(}\rho(1-e^{-\beta})+e^{-\beta}\big{)}+1.$ (24)
These are differentiable functions that convert samples $\rho$ from a uniform distribution $\mathcal{U}(0,1)$ into a sample from $q_{\eta}(\boldsymbol{z},\boldsymbol{c}=\boldsymbol{c}^{\prime}|\boldsymbol{x})$ following
$\begin{gathered}q_{\eta}(\boldsymbol{z},\boldsymbol{c}=\boldsymbol{c}^{\prime}|\boldsymbol{x})=\\ =\prod_{j=1}^{M/R}\Big{[}F^{-1}_{p(z_{j}|\boldsymbol{c}^{\prime}_{j}=1)}(\rho)^{c^{\prime}_{j}}+F^{-1}_{p(z_{j}|\boldsymbol{c}^{\prime}_{j}=0)}(\rho)^{(1-c^{\prime}_{j})}\Big{]}.\end{gathered}$ (25)
Thus, we can apply the reparameterization trick to obtain samples from the latent variable $\boldsymbol{z}$ and optimize the reconstruction term of the ELBO with respect to the parameters $\boldsymbol{\theta}$ of the decoder.
The KL Divergence term is approximated via Monte Carlo, drawing samples from $q_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x})$. As it is not possible to backpropagate through discrete variables, we approximate the gradients with respect to the parameters of the encoder using the REINFORCE leave-one-out estimator [23, 13], given by
$\begin{gathered}\widehat{g}_{LOO}=\\ =\frac{1}{S-1}\Bigg{[}\sum_{s=1}^{S}f_{\boldsymbol{\eta}}\big{(}\boldsymbol{z}^{(s)},\boldsymbol{c}^{(s)}\big{)}\bigtriangledown_{\boldsymbol{\eta}}\log q_{\boldsymbol{\eta}}\big{(}\boldsymbol{c}^{(s)}|\boldsymbol{x}\big{)}-\overline{f}_{\boldsymbol{\eta}}\sum_{s=1}^{S}\bigtriangledown_{\boldsymbol{\eta}}\log q_{\boldsymbol{\eta}}\big{(}\boldsymbol{c}^{(s)}|\boldsymbol{x}\big{)}\Bigg{]},\end{gathered}$ (26)
where
$f_{\boldsymbol{\eta}}(\boldsymbol{z}^{(s)},\boldsymbol{c}^{(s)})=\log\Bigg{(}\frac{q_{\boldsymbol{\eta}}(\boldsymbol{c}^{(s)}|\boldsymbol{x})}{p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z}^{(s)})p(\boldsymbol{c}^{(s)})}\Bigg{)},$ (27)
$\overline{f_{\boldsymbol{\eta}}}=\frac{1}{S}\sum_{s=1}^{S}f_{\boldsymbol{\eta}}(\boldsymbol{z}^{(s)},\boldsymbol{c}^{(s)}).$ (28)
Defining a distribution over the codebook can seem intuitive, but scalability becomes challenging as the size of the codebook increases. The reason is that the posterior distribution must be evaluated for all codewords during both inference and test time. However, it can still provide a bound, enabling the utilization of more complex codes with theoretical guarantees that can outperform the previously proposed repetition codes.
Algorithm 3 Training the model with inference at codeword level.
1:  Input: training data $\boldsymbol{x}_{i}$, codebook.
2:  repeat
3:     $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}|\boldsymbol{x}_{i})\leftarrow$ forward encoder $g_{\boldsymbol{\eta}}(\boldsymbol{x}_{i})$
4:     $q_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x}_{i})\leftarrow$ evaluate $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}|\boldsymbol{x}_{i})$ over the codebook and normalize
5:     $\tilde{\boldsymbol{c}}\leftarrow$ sample from $q_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x}_{i})$
6:     $\boldsymbol{z}\leftarrow$ modulate $\tilde{\boldsymbol{c}}$
7:     $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})\leftarrow$ forward decoder $f_{\boldsymbol{\theta}}(\boldsymbol{z})$
8:     Compute ELBO according to (22)
9:     Compute encoder’s gradients according to (26)
10:     $\boldsymbol{\theta},\boldsymbol{\eta}\leftarrow Update(ELBO)$
11:  until convergence
Appendix K Computational resources
All the experiments in this paper were conducted on a single GPU. Depending on availability, we used either a Titan X Pascal with 10GB of RAM, a Nvidia GeForce GTX with 10GB of RAM, or a Nvidia GeForce RTX 4090 with 24GB of RAM. Since training times varied significantly based on the hardware used, we were unable to provide comparable training times.
Appendix L Coded DVAE scheme
Figure 39: Graphic representation of the coded DVAE, illustrating both the generative and inference paths.
Appendix M Hierarchical coded DVAE scheme
Figure 40: Graphic representation of the hierarchical coded DVAE, illustrating both the generative and inference paths.

### v2

@@
-Protect Before Generate: Error Correcting Codes within Discrete Deep Generative Models
+Improved Variational Inference in Discrete VAEs using Error Correcting Codes
 Abstract
-Despite significant advancements in deep probabilistic models, learning low-dimensional discrete latent representations remains a challenging task.
-In this paper, we introduce a novel method that enhances variational inference in discrete latent variable models by leveraging Error-Correcting Codes to introduce redundancy in the latent representations.
-This redundancy is then exploited by the variational posterior to yield more accurate estimates, thereby narrowing the variational gap.
-Inspired by ECCs commonly used in digital communications and data storage, we demonstrate proof-of-concept using a Discrete Variational Autoencoder (DVAE) with binary latent variables and block repetition codes.
-We further extend this idea to a hierarchical structure based on polar codes, where certain latent bits are more robustly protected.
-Our method improves generation quality, data reconstruction, and uncertainty calibration compared to the uncoded DVAE, even when trained with tighter bounds such as the Importance Weighted Autoencoder (IWAE) objective.
-In particular, we demonstrate superior performance on MNIST, FMNIST, CIFAR10, and Tiny ImageNet datasets.
-The general approach of integrating ECCs into variational inference is compatible with existing techniques to boost variational inference, such as importance sampling or Hamiltonian Monte Carlo.
-We also outline the key properties ECCs must have to effectively enhance discrete variational inference.
-Keywords: deep generative models, variational inference, discrete representations, error correcting codes
+Despite advances in deep probabilistic models, learning discrete latent representations remains challenging.
+This work introduces a novel method to improve inference in discrete Variational Autoencoders by reframing the inference problem through a generative perspective.
+We conceptualize the model as a communication system, and propose to leverage Error Correcting Codes (ECCs) to introduce redundancy in latent representations, allowing the variational posterior to produce more accurate estimates and reduce the variational gap.
+We present a proof-of-concept using a Discrete Variational Autoencoder with binary latent variables and low-complexity repetition codes, extending it to a hierarchical structure for disentangling global and local data features.
+Our approach significantly improves generation quality, data reconstruction, and uncertainty calibration, outperforming the uncoded models even when trained with tighter bounds such as the Importance Weighted Autoencoder objective.
+We also outline the properties that ECCs should possess to be effectively utilized for improved discrete variational inference.
 1 Introduction
-Discrete latent space models seek to represent data using a finite set of features.
-Recent progress in generative models has increasingly favored these representations, as they are well-suited for datasets characterized by naturally discrete hidden states.
-However, effective learning of low-dimensional discrete latent representations is technically challenging.
-Vector Quantized-Variational Autoencoders [31, 21] stand out as solutions for this problem but rely on a non-probabilistic autoencoder, which does not provide uncertainty quantification in the latent space.
-To fit a fully probabilistic Variational Autoencoder (VAE) model [12], a common approach considers either Concrete [18] or Gumble-Softmax [11] approximations to sample from a discrete latent distribution in a reparameterizable manner [20, 17].
-However, this approach leads to instabilities since the gradient variance is sensitive to the temperature that controls these approximations.
-The DVAE in [22, 30, 29], augments the binary latent representations with a set of continuous random variables, pairing each bit with a continuous counterpart where reparameterization can be done in a more stable manner after marginalizing the latent bits.
-The key distinction between the DVAE [22] and the DVAE++ [30] lies in their smoothing transformations: while [22] introduces spike-and-exponential transformations, [30] uses overlapping exponential distributions.
-These overlapping transformations are generalized in [29], enabling tighter variational bounds.
-We demonstrate the effectiveness of our method over a simplified version of the DVAE++ [30].
-This work presents a novel method to improve variational inference and representation learning in generative models with discrete latent variables.
-In particular, for a latent variable model, we argue that one should use ECCs to introduce redundancy into the latent sample before the reconstruction decoder network processes it to generate the data.
-The variational approximation to the true posterior distribution can then exploit the added redundancy to provide more accurate estimates, reducing, in turn, the variational gap to the data likelihood.
-Our approach is based on well-known digital communications and data storage techniques where information is protected with ECCs before transmission/storage to reduce the overall error rate during recovery.
-For different datasets, our results demonstrate that, compared to the uncoded DVAE, the DVAE with ECCs (Coded-DVAE) achieves superior generation quality, better data reconstruction, and critically calibrated uncertainty in the latent space.
-In Fig.
-1, we highlight some representative results for both MNIST and FMNIST data sets.
-We note that the use of ECCs is a general design approach that is perfectly compatible with state-of-the-art techniques for improved variational inference, such as importance sampling [3, 27] or Hamiltonian Monte Carlo [32, 4].
-In summary, our main contributions are:
-  • We provide proof-of-concept results demonstrating that training deep generative models can be improved by ECC techniques, an idea that, to the best of our knowledge, is completely novel in the literature.
-  • We formulate a coded version of DVAE using block repetition codes. We show that encoding/decoding of the block repetition code can be efficiently done with linear complexity.
-  • We show that Coded-DVAE improves reconstruction, generation, and uncertainty calibration in the latent space when compared to the uncoded case using the same latent dimension, even when the uncoded DVAE is trained with tighter bounds such as the IWAE objective [3].
-  • We discuss the generalization of this method to other coding schemes and introduce a hierarchical structure, inspired by polar codes [2], that effectively separates high-level information from finer details.
-  • Through an extensive ablation study, we show that the enhancement in performance is not attributed to the increased dimensionality introduced by the redundancy from the ECC.
-Figure 1: Comparison between uncoded and coded DVAE models with 8 latent bits, where the uncoded models are identified by the number of latent bits, and the coded models by their code rate.
+Discrete latent space models represent data using a finite set of features, making them particularly well-suited for representing inherently discrete data modalities.
+Additionally, they provide benefits in terms of interpretability and computational efficiency compared to continuous representations [Van Den Oord et al., 2017, Jang et al., 2017, Vahdat et al., 2018a].
+Consequently, recent advancements in deep generative models have increasingly embraced discrete latent representations [Razavi et al., 2019, Ramesh et al., 2021, Rombach et al., 2022].
+However, model training and inference of low-dimensional discrete latent representations remains technically challenging due to non-differentiability, which often requires the use of approximations that can lead to unstable optimization [Jang et al., 2017, Kool et al., 2019].
+Currently, Vector Quantized-Variational Autoencoders [Van Den Oord et al., 2017, Razavi et al., 2019] stand out as state-of-the-art discrete latent variable models.
+They rely on a non-probabilistic encoder trained with a straight-through estimator, which lacks uncertainty quantification in the latent space.
+In contrast, fully probabilistic discrete Variational Autoencoders [Kingma and Welling, 2014] employ continuous relaxations such as Gumbel-Softmax [Jang et al., 2017] or Concrete [Maddison et al., 2017] for differentiable sampling [Ramesh et al., 2021, Lievin et al., 2020].
+However, these methods can be unstable, as gradient variance is highly sensitive to the temperature parameter.
+The Discrete Variational Autoencoder (DVAE) framework [Rolfe, 2016, Vahdat et al., 2018a, b] addresses this issue by augmenting binary latent representations with continuous variables, allowing for stable reparameterization after marginalizing the latent bits.
+In this setting, Boltzmann machines serve as discrete priors to enhance model flexibility, at the cost of degraded interpretability.
+This work introduces a novel approach to improve variational inference in discrete latent variable models, particularly discrete VAEs, by offering an alternative perspective on the inference problem.
+Rather than viewing VAEs as lossy compression methods employed to minimize data reconstruction error, we adopt a generative perspective, where inference seeks to recover the latent vector that generated a given observation.
+This allows us to conceptualize the generative model as a communication system, where we reinterpret latent vectors as messages transmitted through a nonlinear channel given by the VAE decoder, and then recovered by the VAE encoder.
+Based on this analogy, we propose to incorporate Error Correcting Codes to introduce redundancy into latent representations before they are processed by the decoder to generate data.
+Figure 1: ECC within a DVAE.
 Fig.
-(a) presents examples of uncurated generation for FMNIST.
+(a) illustrates the method and how adding redundancy transforms a dense $M$-dimensional latent space into a sparse $D$-dimensional one with $2^{M}$ valid vectors, enabling error correction in the variational posterior via minimum distance.
 Fig.
-(b) illustrates the evolution of the reconstruction and regularization terms of the loss.
-Fig.
-(c) depicts the graphical model of the uncoded DVAE, while Fig.
-(d) shows the graphical model of the coded DVAE.
+(b) compares the ELBO and IWAE objectives for coded and uncoded models, where coding schemes with $M=8$ and $D=80,160,240$ introduce $D/8$ redundancy bits per latent bit.
+Our approach builds on well-established techniques from digital communications and data storage, where information is protected using ECCs before transmission or storage to reduce the overall error rate during recovery [Moon, 2005].
+Notably, Shannon’s landmark work demonstrated that properly designed error correction schemes can achieve arbitrarily low error rates [Shannon, 1948].
+We use ECCs to deterministically introduce controlled redundancy to the latent vectors, increasing their Hamming distance and creating a higher-dimensional but sparser space where only a subset of vectors are valid.
+As illustrated in Fig.
+1(a), this increased separation between latent codes allows for some errors in inference while still enabling recovery of the correct latent code by identifying the closest valid vector.
+Hence, this structure facilitates error correction during inference.
+The integration of ECCs into discrete latent variable models is a novel approach that remains compatible with existing training methods for discrete generative models.
+Note that, while deep learning has been widely explored in digital communications [Guo et al., 2022, Wu et al., 2023, Shen et al., 2023, Ye et al., 2024, Chen et al., 2024], the use of ECCs as a design tool in machine learning remains largely unexplored.
+This work introduces a new research direction, leveraging communications theory to enhance deep generative models.
+In our experiments, we use a simplified version of DVAE++ [Vahdat et al., 2018b] (referred to as uncoded DVAE) and show that the added redundancy enables a more accurate variational approximation to the true posterior, using simple independent priors.
+We highlight that error rates and variational gaps are linked through bounds derived from mismatch hypothesis testing, showing that minimizing the variational gap tightens the upper bound on the error rate [Schlüter et al., 2013, Yang et al., 2024].
+Across different datasets, our results indicate that the DVAE with ECCs (Coded-DVAE) leads to reduced error rates in inference, resulting in smaller variational gaps.
+This improvement translates into superior generation quality, improved data reconstruction, and critically calibrated uncertainty in the latent space.
+Notably, Fig.
+1(b) shows that Coded-DVAE achieves a significantly tighter training bounds even when the baseline is trained using the Importance Weighted Autoencoder (IWAE) objective [Burda et al., 2016].
+Contributions.
+The primary contribution of this work is presenting a new perspective on the inference problem by framing the generative model as a communication system.
+We present proof-of-concept results that show how training discrete latent variable models can be improved by incorporating ECC techniques, an approach that, to the best of our knowledge, is novel in the existing literature.
+We then introduce a coded version of a discrete VAE, utilizing block repetition codes, and show that encoding and decoding can be done with linear complexity.
+Our results show that the Coded-DVAE model significantly reduces error rates during inference, leading to better data reconstruction, generation quality, and improved uncertainty calibration in the latent space.
+We further generalize this approach to other coding schemes and propose a hierarchical structure inspired by polar codes [Arikan, 2009], which effectively separates high-level information from finer details.
+[^1] Code: https://github.com/mariamartinezgarcia/codedVAE
 2 Our baseline: the uncoded DVAE
-This section introduces a simplified version of the uncoded DVAE [22, 30, 29], serving as the foundational model upon which the subsequent aspects of our work are constructed.
-Let $\boldsymbol{X}=\{\boldsymbol{x}_{0},\ldots,\boldsymbol{x}_{N}\}$ denote a collection of unlabelled data, where $\boldsymbol{x}_{i}$ represents a K-dimensional feature vector.
-While [22], [30] and [29] use Boltzmann machine priors, we consider a generative probabilistic model characterized by a simple low-dimensional binary latent variable $\boldsymbol{m}\in\{0,1\}^{M}$ comprising independent and identically distributed (i.i.d.) Bernoulli components $p(\boldsymbol{m})=\prod_{j=1}^{M}p(m_{j})=\prod_{j=1}^{M}\text{Ber}(\nu)$.
-Since backpropagation through discrete variables is generally not possible, a smoothing transformation of these binary variables is introduced.
-While the smoothing transformations proposed in [22] are limited to spike-and-X type of transformations, [30] show better results by using truncated exponential distributions:
-$p(\boldsymbol{z}|\boldsymbol{m})=\prod_{j=1}^{M}p(z_{j}|m_{j}),\quad p(z_{j}|m_{j})$ $=\frac{e^{-\beta(z_{j}-m_{j})}}{Z_{\beta}},$ (1)
-for $m_{j}\in\{0,1\}$, $z_{j}\in[0,1]$, and $Z_{\beta}=(1-e^{-\beta})/\beta$.
-The parameter $\beta$ serves as an inverse temperature term, similar to the one in the Gumbel-Softmax relaxation [11].
-Given the simplicity of the defined binary prior, the complexity of the model is primarily determined by the likelihood function $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})=p(f_{\boldsymbol{\theta}}(\boldsymbol{z}))$, where the likelihood is a Neural Network (NN) (referred to as the decoder) with parameter set $\boldsymbol{\theta}$.
-Variational family and inference
-Following [22], we assume an amortized variational family of the following form:
-$q_{\boldsymbol{\eta}}(\boldsymbol{m},\boldsymbol{z}|\boldsymbol{x})$ $=q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})p(\boldsymbol{z}|\boldsymbol{m}),\quad q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})=\prod_{j=1}^{M}\text{Ber}(g_{j,\boldsymbol{\eta}}(\boldsymbol{x})),$ (2)
-where $g_{\boldsymbol{\eta}}(\boldsymbol{x})$ represents a parameterized function; here, a NN (referred to as the encoder) with parameter set $\boldsymbol{\eta}$.
+This section introduces a simplified version of the DVAE++ [Vahdat et al., 2018b] (uncoded DVAE), which serves as the foundational model upon which the subsequent aspects of our work are constructed.
+A key advantage of this framework is its stable training process, facilitated by reparameterization through smoothing transformations, in contrast to more unstable methods like Gumbel-Softmax [Jang et al., 2017].
+To ensure that any performance improvements in the Coded-VAE stem solely from the coding scheme rather than a complex prior, we adopt a simple independent prior.
+Figure 2: Graphical model of the uncoded DVAE.
+Let $\boldsymbol{X}=\{\boldsymbol{x}_{0},\ldots,\boldsymbol{x}_{N}\}$ denote a collection of unlabeled data, where $\boldsymbol{x}_{i}\in\mathbb{R}^{K}$ represents a $K$-dimensional feature vector.
+While Rolfe [2016], Vahdat et al.
+[2018b] and Vahdat et al.
+[2018a] use Boltzmann machine priors, we consider a generative probabilistic model characterized by a simple low-dimensional binary latent variable $\boldsymbol{m}\in\{0,1\}^{M}$ comprising independent and identically distributed (i.i.d.) Bernoulli components $p(\boldsymbol{m})=\prod_{j=1}^{M}p(m_{j})=\prod_{j=1}^{M}\text{Ber}(\nu)$, where $\nu=p(m_{j}=1)$.
+Since backpropagation through discrete samples is generally not possible, a smoothing transformation $p(\boldsymbol{z}|\boldsymbol{m})=\prod_{j=1}^{M}p(z_{j}|m_{j})$ of these binary variables is introduced.
+Following Vahdat et al.
+[2018b], we introduce truncated exponential distributions given by
+$p(z|m=1)=\frac{e^{\beta(z-1)}}{Z_{\beta}},\quad p(z|m=0)=\frac{e^{-\beta z}}{Z_{\beta}},$ (1)
+for $m\in\{0,1\}$, $z\in[0,1]$, and $Z_{\beta}=(1-e^{-\beta})/\beta$.
+The parameter $\beta$ serves as an inverse temperature term, similar to the one in the Gumbel-Softmax relaxation [Jang et al., 2017].
+Given the simplicity of the defined binary prior, the complexity of the model is primarily determined by the likelihood function $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})=p(f_{\boldsymbol{\theta}}(\boldsymbol{z}))$, the decoder function, where the likelihood is a Neural Network (NN) (referred to as the decoder) with parameter set $\boldsymbol{\theta}$.
+2.1 Variational family and inference
+Following Rolfe [2016], we assume an amortized variational family of the form $q_{\boldsymbol{\eta}}(\boldsymbol{m},\boldsymbol{z}|\boldsymbol{x})=q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})p(\boldsymbol{z}|\boldsymbol{m})$, where
+$q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ $=\prod_{j=1}^{M}\text{Ber}(m_{j};q_{j})=\prod_{j=1}^{M}\text{Ber}(m_{j};g_{j,\boldsymbol{\eta}}(\boldsymbol{x}))$ (2)
+where $\text{Ber}(m_{j};q_{j})$ is a Bernoulli with parameter $q_{j}=q_{\boldsymbol{\eta}}(m_{j}=1|\boldsymbol{x})$ and $g_{\boldsymbol{\eta}}(\boldsymbol{x})$ is a NN (referred to as the data encoder) with parameter set $\boldsymbol{\eta}$.
 Inference is achieved by maximizing the Evidence Lower Bound (ELBO), which can be expressed as
 […]
-where the first term corresponds to the reconstruction of the observed data and the second term is the Kullback-Leibler (KL) Divergence between the variational family and the binary prior distribution, which acts as a regularization term.
-This can be computed in closed form as $\mathcal{D}_{KL}\big{(}q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})||p(\boldsymbol{m})\big{)}=\sum_{j=1}^{M}\Big{[}q_{j}\log\frac{q_{j}}{\nu}+(1-q_{j})\log\frac{1-q_{j}}{1-\nu}\Big{]}$, where $q_{j}=q_{\boldsymbol{\eta}}(m_{j}=1|\boldsymbol{x})$.
-The reconstruction term needs to be approximated via Monte Carlo.
+where the first term corresponds to the reconstruction of the observed data and the second is the Kullback-Leibler (KL) Divergence between the variational family and the prior distribution, which acts as a regularization term.
+This can be computed in closed form as $\mathcal{D}_{KL}\big{(}q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})||p(\boldsymbol{m})\big{)}=\sum_{j=1}^{M}\Big{[}q_{j}\log\frac{q_{j}}{\nu}+(1-q_{j})\log\frac{1-q_{j}}{1-\nu}\Big{]}$, with $\nu=p(m_{j}=1)$.
 Since $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})$ does not depend on the binary latent variable $\boldsymbol{m}$, we can marginalize the posterior distribution as
-$q_{\boldsymbol{\eta}}(\boldsymbol{z}|\boldsymbol{x})$ $=\prod_{j=1}^{M}q_{\boldsymbol{\eta}}(z_{j}|\boldsymbol{x}),\quad q_{\eta}(z_{j}|\boldsymbol{x})=\sum_{k=0}^{1}q_{\eta}(m_{j}=k|\boldsymbol{x})p(z_{j}|m_{j}=k).$ (4)
-As shown in [30], the corresponding inverse Cumulative Density Function (CDF) is given by
+$\begin{gathered}q_{\boldsymbol{\eta}}(\boldsymbol{z}|\boldsymbol{x})=\prod_{j=1}^{M}q_{\boldsymbol{\eta}}(z_{j}|\boldsymbol{x}),\\ \quad q_{\eta}(z_{j}|\boldsymbol{x})=\sum_{k=0}^{1}q_{\eta}(m_{j}=k|\boldsymbol{x})p(z_{j}|m_{j}=k).\end{gathered}$ (4)
+As shown in Vahdat et al. [2018b], the corresponding inverse Cumulative Density Function (CDF) is given by
 $F^{-1}_{q_{\boldsymbol{\eta}}(z_{j}|\boldsymbol{x})}(\rho)=-\frac{1}{\beta}\log\left(\frac{-b+\sqrt{b^{2}-4c}}{2}\right),$ (5)
 where $b=\big{(}\rho+e^{-\beta}(q_{j}-\rho)\big{)}/(1-q_{j})-1$ and $c=-[q_{j}e^{-\beta}]/(1-q_{j})$.
-The (5) is a differentiable function that converts a sample $\rho$ from an independent uniform distribution $\mathcal{U}(0,1)$ into a sample from $q_{\eta}(\boldsymbol{z}|\boldsymbol{x})$.
-Thus, we can apply the reparameterization trick to sample from the latent variable $\boldsymbol{z}$ and optimize the ELBO with respect to the model’s parameters.
-Algorithm 1 outlines the training procedure for the uncoded DVAE.
-Algorithm 1 Training the model with uncoded inference.
-1:  Input: training data $\boldsymbol{x}_{i}$.
-2:  repeat
-3:     $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{m}|\boldsymbol{x}_{i})\leftarrow$ forward encoder $g_{\boldsymbol{\eta}}(\boldsymbol{x}_{i})$
-4:     $\boldsymbol{z}\leftarrow$ sample from (5)
-5:     $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})\leftarrow$ forward decoder $f_{\boldsymbol{\theta}}(\boldsymbol{z})$
-6:     Compute ELBO according to (3)
-7:     $\boldsymbol{\theta},\boldsymbol{\eta}\leftarrow Update(ELBO)$
-8:  until convergence
-3 Improving inference by adding redundancy to latent vectors
-In the DVAE framework [22, 30, 29], the authors use Boltzmann machines as priors instead of the independent prior $p(\boldsymbol{m})$ presented in Section 2.
-While these complex priors increase the model’s flexibility and can produce competitive results, our objective is to enhance inference through model design by maintaining the simpler independent prior.
-This approach would improve interpretability and encourage the model to learn independent components in the latent space, which is essential for capturing potentially disentangled representations.
-VAEs [12] are often viewed as lossy compression models, where the goal is to minimize reconstruction error while imposing regularization through a prior distribution.
-However, our approach is better understood from a generative standpoint.
-We first sample a latent vector $\boldsymbol{m}$, generate an observation $\boldsymbol{x}$, and focus on minimizing the error rate when recovering $\boldsymbol{m}$ from $\boldsymbol{x}$.
-Achieving this requires the variational approximation to be sufficiently accurate.
-In fields where reliable data transmission or storage is important, introducing ECCs is a well-established approach to reduce the error rate when estimating a discrete source $\boldsymbol{m}$ transmitted through a noisy channel with output $\boldsymbol{x}$.
-Estimating $\boldsymbol{m}$ from $\boldsymbol{x}$ implies approximating the true and unknown posterior distribution $p(\boldsymbol{m}|\boldsymbol{x})$ with a proposed $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$.
-The gap between $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ and $p(\boldsymbol{m}|\boldsymbol{x})$ is precisely the variational gap.
-We propose employing ECCs to safeguard $\boldsymbol{m}$ with controlled and known redundancy that can be leveraged by the variational posterior $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ by design.
-This way, it is possible to reduce the mistakes committed when comparing $\boldsymbol{m}$ with samples drawn from $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$, obtaining a tighter approximation to the true posterior $p(\boldsymbol{m}|\boldsymbol{x})$, therefore reducing the gap to optimal inference.
-ECCs play a crucial role in information theory and digital communications by enabling reliable data transmission over unreliable channels [19].
-They introduce redundancy into the transmitted data, allowing the receiver to detect errors and, in many cases, correct them without retransmission.
-In his seminal work, Shannon [25] demonstrated that arbitrarily reliable communication is possible through error correction.
-Our approach builds on the idea that the generative model in Fig.
-1(c) can be conceptualized as a communication system, where the bits sampled from $p(\boldsymbol{m})$ undergo continuous modulation into $\boldsymbol{z}$ and are then transmitted through a nonlinear communication channel (in this setting, the decoder NN) characterized by the input/output response $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})=p(f_{\boldsymbol{\theta}}(\boldsymbol{z}))$.
-In this scenario, the complexity of the channel is essential since it is necessary to account for the intricate nature of the data at its output (e.g., complex images).
-Following this idea, the process of inference via $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ can be thought of as deciphering the latent variable $\boldsymbol{m}$ given the observed data $\boldsymbol{x}$, where the encoder NN plays the role of the channel equalizer, trying to reverse the channel’s effects without knowing the bit correlations from the ECC.
-4 Coded DVAE
+Equation (5) defines a differentiable function that converts a sample $\rho$ from an independent uniform distribution $\mathcal{U}(0,1)$ into a sample from $q_{\boldsymbol{\eta}}(\boldsymbol{z}|\boldsymbol{x})$.
+Thus, we can apply the reparameterization trick to sample from the latent variable $\boldsymbol{z}$ and optimize the ELBO.
+3 On the variational gap and the inference error
+While VAEs are typically viewed as lossy compression methods, this work introduces an alternative perspective on the inference problem from a generative viewpoint: we first sample a latent vector $\boldsymbol{m}$, generate an observation $\boldsymbol{x}$, and aim to minimize the error rate when recovering $\boldsymbol{m}$ from $\boldsymbol{x}$.
+This requires approximating the true, unknown posterior distribution $p_{\boldsymbol{\theta}}(\boldsymbol{m}|\boldsymbol{x})$ with a sufficiently accurate proposed distribution $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$.
+The gap between $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ and $p_{\boldsymbol{\theta}}(\boldsymbol{m}|\boldsymbol{x})$ is precisely the variational gap, which can be related to the error rate in inference, i.e.
+comparing the true $\boldsymbol{m}$ with samples drawn from $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$.
+In statistical classification, particularly in the context of multiple hypothesis testing, classification error is the primary performance metric.
+Bayes’ decision rule (also known as Bayes’ test) minimizes classification error, assuming the true probability distributions governing the classification problem are known.
+In the case of model mismatch, e.g., when a posterior distribution is learned using variational inference, statistical bounds can be used to relate the classification error and the model estimation error [Schlüter et al., 2013, Yang et al., 2024].
+In particular, for a generative model $\boldsymbol{x}\sim p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{m})$ with discrete latent vector $\boldsymbol{m}$, let $1-A_{\boldsymbol{\theta}}$ be the error rate when estimating $\boldsymbol{m}$ from $\boldsymbol{x}$ using the true posterior $p_{\boldsymbol{\theta}}(\boldsymbol{m}|\boldsymbol{x})$ and $1-A_{\boldsymbol{\eta}}$ be the error rate using the variational approximate posterior $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$.
+Then, $\Delta\doteq A_{\boldsymbol{\theta}}-A_{\boldsymbol{\eta}}$ can be bounded as
+$0\leq\Delta^{2}$ $\leq 1-e^{-2\mathcal{D}_{KL}(q_{\boldsymbol{\eta}}(\boldsymbol{m},\boldsymbol{x})||p_{\boldsymbol{\theta}}(\boldsymbol{m},\boldsymbol{x}))},$ (6)
+where $\mathcal{D}_{KL}(q_{\boldsymbol{\eta}}(\boldsymbol{m},\boldsymbol{x})||p_{\boldsymbol{\theta}}(\boldsymbol{m},\boldsymbol{x}))=\mathbb{E}_{p(\boldsymbol{x})}[\mathcal{D}_{KL}(q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})||p_{\boldsymbol{\theta}}(\boldsymbol{m}|\boldsymbol{x}))]$ is the variational gap.
+For fixed $\boldsymbol{\theta}$ (and thus for fixed $A_{\boldsymbol{\theta}}$), any improvement in $\boldsymbol{\eta}$ that reduces the expected variational gap also tightens an upper bound on $\Delta\doteq A_{\boldsymbol{\theta}}-A_{\boldsymbol{\eta}}$.
+In fields where reliable data transmission or storage is crucial, introducing ECCs is a well-established approach to reduce the error rate when estimating a discrete source $\boldsymbol{m}$ transmitted through a noisy channel with output $\boldsymbol{x}$.
+Building on this, we propose using ECCs to protect $\boldsymbol{m}$ with controlled redundancy that the variational posterior $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ can leverage.
+This way, it is possible to reduce error rates in inference and refine the variational parameters $\boldsymbol{\eta}$ on the enhanced model, thus narrowing the variational gap.
+Notably, according to (6), a smaller variational gap results in a tighter upper bound on the error rates.
+In the following sections, we show that simple coding schemes enhance inference, improving generation, reconstruction, log-likelihoods, and uncertainty calibration in the latent space.
+4 Coded-DVAE
+Figure 3: Graphical model of the Coded-DVAE.
+Model BER BER MAP WER WER MAP LL test
+uncoded 5 0.051 0.051 0.195 0.190 -267.703
+[…]
+[…]
+[…]
+uncoded 8 0.089 0.088 0.384 0.376 -249.880
+coded 8/80 0.021 0.003 0.144 0.021 -232.992
+coded 8/160 0.027 0.001 0.189 0.009 -235.819
+coded 8/240 0.037 0.004 0.231 0.023 -238.459
+uncoded 10 0.142 0.144 0.622 0.644 -244.997
+coded 10/100 0.040 0.005 0.321 0.036 -230.772
+coded 10/200 0.044 0.006 0.341 0.039 -234.849
+coded 10/300 0.045 0.002 0.349 0.014 -238.647
+Figure 4: Evaluation of generation in FMNIST.
+Uncurated randomly generated samples (left) and evaluation metrics (right), including BER and WER sampled from the variational posterior, BER and WER using the MAP value, and test LL.
 This section extends the previously described DVAE, introducing an ECC over $\boldsymbol{m}$.
-We refer to this model as coded DVAE.
-In ECCs, we augment the dimensionality of the binary latent space from $M$ to $D$ in a controlled and deterministic manner, where $R=M/D$ is the coding rate.
-An ECC is typically designed so that the $2^{M}$ possible codewords are separated as much as possible in the space of binary vectors of $D$ bits.
-This facilitates algorithms in detecting and/or correcting errors by searching for the nearest code word.
-A random choice of the codewords brings what is known as a random block code [25].
-While they are known to be very robust and amenable to theoretical analysis, their lack of structure makes them computationally intractable since we have to rely on codeword enumeration during the encoding/decoding process.
-In Appendix J, we include the formulation of a random code’s encoding/decoding process within the DVAE model.
-Instead, we adopt a much simpler linear coding scheme, namely repetition codes.
-In a repetition code, each bit of the original message $\boldsymbol{m}$ is repeated multiple times to create the encoded message $\boldsymbol{c}$.
+We refer to this model as Coded-DVAE.
+The corresponding graphical model is shown in Fig.
+3, where the diamond node represents a deterministic transformation.
+In ECCs, we augment the dimensionality of the binary latent space from $M$ to $D$, introducing redundancy in a controlled and deterministic manner, where $R=M/D$ is the coding rate.
+An ECC is designed to maximize the separation between the $2^{M}$ possible codewords within the $D$-bit binary space, enabling efficient error correction by searching for the nearest valid codeword (see Fig.
+1(a)).
+A random selection of codewords results in random block codes [Shannon, 1948], which are very robust and amenable to theoretical analysis.
+However, their lack of structure makes them impractical, as encoding and decoding require exhaustive codeword enumeration.
+Appendix N provides a detailed formulation of the encoding and decoding process for a random code within the Coded-DVAE.
+We adopt a much simpler linear coding scheme, namely repetition codes, where each bit of the original message $\boldsymbol{m}$ (i.e., each information bit) is duplicated multiple times to form the encoded message $\boldsymbol{c}$.
 Intuitively, the more times an information bit is repeated, the better it is protected.
 Our experiments consider uniform $(M,D)$ repetition codes where all bits are repeated $L$ times, resulting in codewords of dimension $D=ML$ and a coding rate of $R=1/L$.
-Note that repetition codes represent a special case of linear ECCs since each codeword can be deterministically computed by multiplying a binary vector $\boldsymbol{m}$ by an $M\times D$ generator matrix $\mathbf{G}$, such that $\boldsymbol{c}=\boldsymbol{m}^{T}\mathbf{G}$, where $u$-th row, with $u=1,\ldots,M$, has entries equal to one at columns $L(u-1)+1,L(u-1)+2,\dots,Lu$, and zero elsewhere.
+Note that repetition codes represent a special case of linear ECCs since each codeword can be computed by multiplying a binary vector $\boldsymbol{m}$ by an $M\times D$ generator matrix $\mathbf{G}$, such that $\boldsymbol{c}=\boldsymbol{m}^{T}\mathbf{G}$, where $k$-th row, with $k=1,\ldots,M$, has entries equal to one at columns $L(k-1)+1,L(k-1)+2,\dots,Lk$, and zero elsewhere.
 For example, for $M=3$ and $L=2$, the generator matrix of the $(3,6)$ repetition code is
 $\mathbf{G}=\begin{bmatrix}1&1&0&0&0&0\\ 0&0&1&1&0&0\\ 0&0&0&0&1&1\end{bmatrix}.$ (7)
-The generative process of the coded DVAE follows similarly to the uncoded case, and it is represented in Fig.
-1(d).
+The generative process of the Coded-DVAE is similar to the uncoded version and is illustrated in Fig.
+3.
 We assume the same prior distribution $p(\boldsymbol{m})$, but in this case the samples $\boldsymbol{m}$ are deterministically encoded using $\mathbf{G}$.
-Now, the smoothing $\boldsymbol{z}$ transformations are defined over $\boldsymbol{c}$
-$p(\boldsymbol{z}|\boldsymbol{c})=\prod_{j=1}^{D}p(z_{j}|c_{j}),\quad p(z_{j}|c_{j})$ $=\frac{e^{-\beta(z_{j}-c_{j})}}{Z_{\beta}},$ (7)
-for $z_{j}\in[0,1]$, $c_{j}\in\{0,1\}$ and $Z_{\beta}=(1-e^{-\beta})/\beta$.
+Consequently, the smoothing transformations are now defined over $\boldsymbol{c}$
+$p(\boldsymbol{z}|\boldsymbol{c})=\prod_{j=1}^{D}p(z_{j}|c_{j}),$ (8)
+with $p(z_{j}|c_{j})$ following Eq.
+(1).
 The likelihood $p(\boldsymbol{x}|\boldsymbol{z})$ is again of the form $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})=p(f_{\boldsymbol{\theta}}(\boldsymbol{z}))$.
-Note that, compared to the uncoded case, we have a larger input dimensionality to the decoder NN $f_{\boldsymbol{\theta}}(\boldsymbol{z})$.
-When comparing uncoded vs.
-coded DVAEs, the structure of the decoder NN $f_{\boldsymbol{\theta}}(\boldsymbol{z})$ (detailed in Appendix A) is equal in both cases except for the first Multilayer Perceptron (MLP) layer that attacks the input $\boldsymbol{z}$.
-Therefore, if a rate $R=1/L$ repetition code is used, the number of additional parameters of the $f_{\boldsymbol{\theta}}(\boldsymbol{z})$ NN is given by $(L-1)\times h$, where $h$ is the dimension of the first hidden space of $f_{\boldsymbol{\theta}}(\boldsymbol{z})$.
-Variational family and inference
-The repetition code introduces correlations between the bits in $\boldsymbol{c}$ that we will exploit to obtain an improved variational bound. We again assume a variational family factorizing as
-$q_{\boldsymbol{\eta}}(\boldsymbol{m},\boldsymbol{z}|\boldsymbol{x})=q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})p(\boldsymbol{z}|\boldsymbol{c})$ (8)
-where $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})=\prod_{u=1}^{M}q_{\boldsymbol{\eta}}(m_{u}|\boldsymbol{x})$ is computed in two steps.
-First, we construct an encoder NN $\mathbf{g}_{\boldsymbol{\eta}}(\boldsymbol{x})$ similar to that of (2), that retrieves the probabilities of the bits in $\boldsymbol{c}$ from $\boldsymbol{x}$ without exploiting the correlations introduced by the repetition code:
-$q^{u}_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x})$ $=\prod_{j=1}^{D}\text{Ber}(g_{j,\boldsymbol{\eta}}(\boldsymbol{x})),$ (9)
-where the $u$ superscript serves as a reminder that this posterior does not exploit the redundancy introduced by the ECC.
-Now, we utilize the known redundancy introduced by the ECC to constrain the solution of $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}|\boldsymbol{x})$, given that each bit from $\boldsymbol{m}$ has been repeated $L$ times to create $\boldsymbol{c}$.
-To do so, we follow a soft decoding approach, where the marginal posteriors of the information bits are derived from the marginal posteriors of the encoded bits, exploiting the repetition code’s known structure.
-In the case of repetition codes, we compute the all-are-zero and the all-are-ones products of probabilities of the bits in $\boldsymbol{c}$ that are copies of the same message bit and renormalize as
-$q(m_{u}=1|\boldsymbol{x})$ $=\frac{1}{Z}\prod_{j=L(u-1)+1}^{Lu}g_{j,\boldsymbol{\eta}}(\boldsymbol{x})\doteq\frac{g^{+}_{u,\boldsymbol{\eta}}(\boldsymbol{x})}{Z},$ (10)
-$q(m_{u}=0|\boldsymbol{x})$ $=\frac{1}{Z}\prod_{j=L(u-1)+1}^{Lu}(1-g_{j,\boldsymbol{\eta}}(\boldsymbol{x}))\doteq\frac{g^{-}_{u,\boldsymbol{\eta}}(\boldsymbol{x})}{Z},$ (11)
-for $u=1,\ldots,M$ and $Z=\left(g^{+}_{u,\boldsymbol{\eta}}(\boldsymbol{x})+g^{-}_{u,\boldsymbol{\eta}}(\boldsymbol{x})\right)^{-1}$.
-This approach can be seen as a soft majority voting strategy, enabling the recovery of the original information vector even if some bits in the inferred encoded word are corrupted.
-All operations in (10) preserve the gradients concerning the parameters in the encoder $\mathbf{g}_{\boldsymbol{\eta}}(\boldsymbol{x})$.
-We implement them in the log domain for stability.
-When compared to the uncoded case, as in the likelihood term, we consider the same NN structure for the encoder $\mathbf{g}_{\boldsymbol{\eta}}(\boldsymbol{x})$ where both cases only differ in the last MLP layer.
-The additional overhead in the coded cases requires $(L-1)\times h^{\prime}$ parameters in the last layer, where $h^{\prime}$ is the dimension at the output of the last layer.
-In Appendix H.2, we conduct an ablation study on the number of trainable parameters to demonstrate that the improvement in performance does not stem from this increase in the number of parameters, but rather from the incorporation of the ECC in the latent space.
-Algorithm 2 Training the coded DVAE with repetition codes.
-1:  Input: training data $\boldsymbol{x}_{i}$, matrix $\mathbf{G}$.
-2:  repeat
-3:     $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}|\boldsymbol{x}_{i})\leftarrow$ forward encoder $g_{\boldsymbol{\eta}}(\boldsymbol{x}_{i})$
-4:     $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x}_{i})\leftarrow$ soft decoding by aggregating $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}|\boldsymbol{x}_{i})$ according to (10)
-5:     $q_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x}_{i})\leftarrow$ repeat posterior bit probabilities $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x}_{i})$ according to $\mathbf{G}$
-6:     $\boldsymbol{z}\leftarrow$ sample from (5)
-7:     $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})\leftarrow$ forward decoder $f_{\boldsymbol{\theta}}(\boldsymbol{z})$
-8:     Compute ELBO according to (3)
-9:     $\boldsymbol{\theta},\boldsymbol{\eta}\leftarrow Update(ELBO)$
-10:  until convergence
-Soft encoding for efficient reparameterization
-Given the variational family in (8), the ELBO matches the expression in (3).
-However, the reparameterization trick in (5) requires independent bits, which is not the case in $\boldsymbol{c}$.
-To efficiently circumvent this issue during training, we employ a soft encoding approach.
-With soft encoding, a marginal probability is computed for each bit in the codeword $\boldsymbol{c}$, taking into account the structure of the ECC and the marginal probabilities of the information bits.
-For a repetition code, this involves simply replicating the posterior probabilities $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})=\prod_{u=1}^{M}q_{\boldsymbol{\eta}}(m_{u}|\boldsymbol{x})$ for each copy of the same information bit.
-Hence, we treat the bits in $\boldsymbol{c}$ as independent but distributed according to $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$.
-The Algorithm 2 shows the training pseudo-code.
-When marginalizing $\boldsymbol{c}$ using the soft encoding marginals, we disregard potential correlations between the coded bits.
-For instance, with repetition codes, sampling from the marginals could produce inconsistent bits, leading to an invalid codeword.
-However, since we do not sample the coded bits during training but instead propagate their marginal probabilities, we consider this approximation to have minimal negative impact.
-In fact, it can be seen as a form of probabilistic dropout, which enhances robustness during training.
-It is important to note that when sampling from the generative model in test time, we use hard bits encoded into valid codewords, yielding visually appealing samples, indicating that our training approach is reliable.
-Related work
-While the use of deep neural networks and generative models in digital communications problems have been profusely reported in recent years (see [36], [5], [9], [33], and [26] for representative examples), the use of ECC techniques as a design tool in machine learning is scarce.
-The most prominent example is [7], where the authors address multiclass learning problems via ECCs.
-In [1], the authors proposed using linear codes for applications that may require removing the trace of a sample from the system, e.g., a user requests their data to be deleted or corrupted data is discovered.
-They address a regression problem by introducing a coded learning protocol that employs linear encoders to divide the training data before the learning phase.
-More recently, in [35] the authors introduced ECCs to improve code-to-code translation using transformers.
-4.1 Reconstruction
-We first evaluate the model’s performance of reconstructing data by examining its uncoded and coded versions across different configurations, varying the number of information bits and code rates.
-The introduction of the repetition code led to improved reconstruction and smaller KL values, indicating that the posterior latent features are disentangled and less correlated.
-In Appendices B, C, D and E, we show the behavior of the ELBO loss function for all the models and datasets.
-Image reconstruction quality.
-In the table included in Fig.
-2, we first quantify the quality of the reconstructions in FMNIST by measuring the Peak Signal-To-Noise Ratio (PSNR) in the test set.
-The results for the rest of the datasets are provided in sections B, C, D and E of the Appendix.
-In all the cases, the coded models yield higher PSNR values than their uncoded counterparts, indicating a superior performance in reconstruction.
-This improvement is also evident by visual inspection of Fig.
-2, where the coded models exhibit a greater ability to capture details in the images for the same latent dimension.
-We observe a general improvement in PSNR as we increase the number of information bits, i.e., as we augment the latent dimensionality of the model.
-This increase in the number of available latent vectors provides greater flexibility, enabling the models to capture the underlying structure of the data more effectively.
-We also observe a general improvement in PSNR as we decrease the code rate, i.e., as we add more redundancy.
-Note that adding redundancy does not increase the model’s flexibility, since the information bits determine the number of latent vectors.
-However, coded models yield more accurate and detailed reconstructions.
+Compared to the uncoded case, the input dimensionality to the decoder $f_{\boldsymbol{\theta}}(\boldsymbol{z})$ is $L$ times larger due to the introduced redundancy.
+Therefore, the overall architecture of the decoder (detailed in Appendix C) remains unchanged, except for the first Multilayer Perceptron (MLP) layer that processes the input $\boldsymbol{z}$.
+When using a repetition code with rate $R=1/L$, the additional parameters of $f_{\boldsymbol{\theta}}(\boldsymbol{z})$ amount to $(L-1)\times h$, where $h$ is the dimension of the first hidden layer.
+4.1 Variational family and inference
+Since $\boldsymbol{c}$ is a deterministic transformation of $\boldsymbol{m}$, its randomness is entirely determined by $\boldsymbol{m}$.
+Therefore, following the recognition model described in Fig.
+3, we assume a variational family factorizing as
+$q_{\boldsymbol{\eta}}(\boldsymbol{m},\boldsymbol{z}|\boldsymbol{x})=q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})p(\boldsymbol{z}|\boldsymbol{c}),$ (9)
+where $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})=\prod_{j=1}^{M}q_{\boldsymbol{\eta}}(m_{j}|\boldsymbol{x})$ is computed in two steps.
+First, we construct an encoder $g_{\boldsymbol{\eta}}(\boldsymbol{x})$ similar to the one in (2), which estimates $\boldsymbol{c}$ from $\boldsymbol{x}$
+$q^{u}_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x})$ $=\prod_{j=1}^{D}q^{u}_{\boldsymbol{\eta}}(c_{j}|\boldsymbol{x})=\prod_{j=1}^{D}\text{Ber}(c_{j};g_{j,\boldsymbol{\eta}}(\boldsymbol{x})),$ (10)
+where the superscript $u$ indicates that this posterior does not utilize the ECC (uncoded).
+Now, we leverage the known redundancy introduced by the ECC to constrain the solution of $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}|\boldsymbol{x})$, given that each bit from $\boldsymbol{m}$ has been repeated $L$ times to create $\boldsymbol{c}$.
+To do so, we follow a soft decoding approach, where the marginal posteriors of the information bits are derived from the marginal posteriors of the encoded bits, exploiting the code’s known structure.
+In the case of repetition codes, we compute the all-are-zero and the all-are-ones products of probabilities of the bits in $\boldsymbol{c}$ that are copies of the same message bit and re-normalize as
+$\begin{gathered}q(m_{k}=1|\boldsymbol{x})=\frac{1}{Z}\prod_{j=L(k-1)+1}^{Lk}q^{u}_{\boldsymbol{\eta}}(c_{j}=1|\boldsymbol{x}),\\ q(m_{k}=0|\boldsymbol{x})=\frac{1}{Z}\prod_{j=L(k-1)+1}^{Lk}q^{u}_{\boldsymbol{\eta}}(c_{j}=0|\boldsymbol{x}),\end{gathered}$ (11)
+for $k=1,\ldots,M$ and $Z$ is the normalization constant. For the $M=3$, $L=2$ toy example in (7)
+$q_{\boldsymbol{\eta}}(m_{1}=1|\boldsymbol{x})\propto q_{\boldsymbol{\eta}}^{u}(c_{1}=1|\boldsymbol{x})q_{\boldsymbol{\eta}}^{u}(c_{2}=1|\boldsymbol{x}),$
+$q_{\boldsymbol{\eta}}(m_{2}=1|\boldsymbol{x})\propto q_{\boldsymbol{\eta}}^{u}(c_{3}=1|\boldsymbol{x})q_{\boldsymbol{\eta}}^{u}(c_{4}=1|\boldsymbol{x}),$
+$q_{\boldsymbol{\eta}}(m_{3}=1|\boldsymbol{x})\propto q_{\boldsymbol{\eta}}^{u}(c_{5}=1|\boldsymbol{x})q_{\boldsymbol{\eta}}^{u}(c_{6}=1|\boldsymbol{x}).$
+This approach can be seen as a soft majority voting strategy.
+All operations in (11) are differentiable and implemented in the log domain.
+We consider the same architecture for the encoder $g_{\boldsymbol{\eta}}(\boldsymbol{x})$ in (10) as in the uncoded case, with the only difference being the final MLP layer.
+The additional overhead in the coded case requires $(L-1)\times h^{\prime}$ parameters in the last layer, where $h^{\prime}$ is the output dimension.
+In Appendix K.2, we conduct an ablation study on the number of trainable parameters to show that performance gains do not stem from the increased parameter count.
 Model PSNR Acc Conf. Acc Entropy
@@
 coded 10/300 $\mathbf{18.111}$ $\mathbf{0.817}$ 0.870 5.076
-Figure 2: Reconstruction performance over the test set in FMNIST.
+Figure 5: Reconstruction performance in FMNIST.
 The figure at the left shows an example of reconstructed test images obtained with different model configurations.
@@
 The table at the right includes reconstruction metrics.
-Acc is the semantic accuracy and Conf.
+Acc is the semantic accuracy, and Conf.
 Acc the confident semantic accuracy.
 Entropy is the average entropy of $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ in the test set.
+4.2 Efficient reparameterization
+Given the variational family in (9), the ELBO remains unchanged from (3).
+However, the reparameterization trick in (5) assumes independent bits, which does not hold for $\boldsymbol{c}$.
+To address this issue efficiently during training, we adopt a soft encoding approach.
+We compute a marginal probability for each bit in $\boldsymbol{c}$, leveraging the ECC structure and the marginal posterior probabilities of the information bits in $\boldsymbol{m}$.
+For a repetition code, this involves simply replicating the posterior probabilities $q_{\boldsymbol{\eta}}(m_{k}|\boldsymbol{x}),k=1,\ldots,M$, for each copy of the same information bit [^2] For the example in (7), this results in $p_{\boldsymbol{\theta}}(\boldsymbol{c}|\boldsymbol{x})\approx q_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x})=\text{Ber}(c_{1};q_{1})\text{Ber}(c_{2};q_{1})\text{Ber}(c_{3};q_{2})\text{Ber}(c_{4};q_{2})\text{Ber}(c_{5};q_{3})\text{Ber}(c_{6};q_{3})$..Hence, we treat the bits in $\boldsymbol{c}$ as independent but distributed according to $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$.
+The training algorithm can be found in Appendix B.
+When marginalizing $\boldsymbol{c}$ using the soft encoding marginals, we disregard potential correlations between the coded bits.
+Consequently, sampling from the marginals may produce an invalid codeword.
+However, since we do not sample the coded bits during training but instead propagate their marginal probabilities, we consider this method to have minimal negative impact.
+In fact, it can be seen as a form of probabilistic dropout, which enhances robustness during training.
+Since encoded words may contain inconsistent bits, the decoder learns to utilize the correlated inputs from repeated bits rather than disregarding them.
+Remark that when sampling from the generative model in test time, we use hard bits encoded into valid codewords, yielding visually appealing samples, indicating effective training.
+5 Results
+This section empirically evaluates the Coded-DVAE with repetition codes and its uncoded counterpart on reconstruction and generation tasks.
+In particular, we present results for MNIST [Deng, 2012], FMNIST [Xiao et al., 2017], CIFAR10 [Krizhevsky et al., 2009], and Tiny ImageNet [Le and Yang, 2015].
+Additionally, we compare the coded model to the uncoded DVAE trained with the IWAE objective (see Fig.
+1 and Appendix H).
+All experiments use a fixed architecture (see Appendix C), modifying the encoder’s output and decoder’s input to accommodate the repetition code.
+We identify the models based on the number of information bits for uncoded models and the code rate for coded models.
+5.1 Generation
+We first evaluate the models using the image generation task.
+Fig.
+4 presents examples of randomly generated images using various model configurations on FMNIST.
+Additional results for other datasets are provided in the Supplementary Material.
+All models produce more detailed and diverse images as the number of information bits increases.
+However, when the number of latent vectors becomes too high for the dataset’s complexity, some codebook entries remain unspecialized during training.
+This results in generation artifacts, where images contain overlapping features from different object classes.
+A visual inspection of Fig.
+4 suggests that such artifacts occur more frequently in the uncoded model.
+Error metrics.
+The effectiveness of the ECC-based inference can be assessed by generating images and measuring errors in recovering $\boldsymbol{m}$ by either drawing one sample from the variational posterior $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ or using the Maximum a Posteriori (MAP) estimate.
+As shown in the table included in Fig.
+4, the coded models significantly reduce both Bit Error Rate (BER) and Word Error Rate (WER) compared to the uncoded case for the same number of latent bits.
+Note also that the error rates grow with the number of latent bits, which is expected due to the increased complexity of the inference process.
+Log-likelihood (LL).
+Aditionally, we estimated the data LL using importance sampling with 300 samples per observation.
+As shown in the table included in Fig.
+4, coded models consistently outperform their uncoded counterparts on both train and test sets, aligning with the rest of the results.
+LL values generally improve with more information bits, reflecting increased model flexibility.
+However, reducing the code rate does not enhance LL, suggesting possible decoder overfitting, where reconstruction improves, but LL declines.
+This may be due to the use of feed-forward networks at the decoder’s input, which might not fully capture the correlations in coded words (see Appendix L for details).
+5.2 Reconstruction
+We also assess the model’s performance in image reconstruction. In the table in Fig. 5, we report the Peak Signal-To-Noise Ratio (PSNR) of reconstructions on the FMNIST test set, calculated as
+$\text{PSNR}=20\log\Big{(}\frac{\text{max}(x_{i,j})}{\text{RMSE}(\boldsymbol{x},\boldsymbol{x}^{\prime})}\Big{)},$ (12)
+where $\text{max}(x_{i,j})$ is the highest pixel value, RMSE represents the Root Mean Square Error, and $\boldsymbol{x}^{\prime}$ the reconstructed image.
+Results for the other datasets can be found in the Supplementary Material.
+In all cases, coded models outperform their uncoded counterparts in reconstruction quality, as indicated by higher PSNR values, a trend also visible in Fig.
+5.
+As the number of information bits increases, PSNR improves, reflecting greater model flexibility to capture data structure.
+We also observe improved PSNR with reduced code rates, as more redundancy is added.
+However, this increased redundancy does not enhance the model’s flexibility, which is determined by the number of information bits.
+Figure 6: Example of erroneous reconstructions in FMNIST.
+The first column in each image shows the original images, while the second column displays the reconstructions.
+The $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ probability is indicated in each row.
 Semantic Accuracy.
-As the PSNR operates at the pixel level, it does not account for the semantic errors committed by the model.
-For example, if the model incorrectly reconstructs a nine instead of a four in the MNIST dataset, the PSNR may still yield a large value due to the similarity between the two images.
-Nonetheless, this would represent a severe failure in correct reconstruction of the intended class.
-Therefore, we additionally evaluate the reconstruction accuracy, ensuring that the model successfully reconstructs images within the same class as the original ones.
-For this purpose, we trained an image classifier for each dataset and compared the reconstructed images’ predicted labels against the originals’ ground truth labels.
-Additionally, we provide a confident reconstruction accuracy.
-While the reconstruction accuracy is computed across the entire dataset partitions, for the confident accuracy, we only consider those images projected into a latent vector with a probability exceeding 0.4.111Namely, we do not count errors when the Maximum a Posteriori (MAP) value of $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ is below 0.4..
-Results for FMNIST are detailed in the table included in Fig.2, and corresponding results for MNIST are available in Table 2 within Appendix C.
-In light of the results, we can conclude that introducing an ECC in the model allows for latent spaces that better capture the semantics of the images while employing the same number of latent vectors, significantly outperforming the uncoded models in terms of accuracy in all the cases.
-Figure 3: Example of erroneous reconstructions in FMNIST using the 4 most-probable a posteriori latent vectors.
-The first column in each image shows the original input to the model, while the second column displays the reconstructions.
-The $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ probability is indicated in each row.
+Since the PSNR operates at the pixel level, it does not account for the semantic errors committed by the model.
+To address this, we evaluate reconstruction accuracy in Fig.
+5, checking whether the model reconstructs images within the correct semantic class (type of clothing in FMNIST).
+For this, we trained an image classifier for each dataset and compared the predicted labels of the reconstructed images with the ground truth labels.
+We also report a confident reconstruction accuracy, which is calculated by considering only the images projected into a latent vector with posterior probability above 0.4  [^3] We do not count errors when the MAP of $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})<0.4$..
+Results indicate that incorporating an ECC enables the model to produce latent spaces that better capture the semantics of the data.
 Posterior Uncertainty Calibration.
-Finally, also in the table included in Fig.2, we report the average entropy of the variational posterior $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ over the test set.
-The low entropy observed in the uncoded models suggests a low uncertainty when the model projects data points into the latent space, which could be advantageous if the model consistently assigned high probability to the correct latent vectors.
-However, the semantic accuracy results demonstrate this is not true in the uncoded model.
-In other words, the uncoded variational family projects images into the wrong class with high confidence.
-This indicates the uncertainty of the uncoded case is severely miscalibrated.
-Coded models, on the other hand, improve semantic accuracy and present a larger entropy.
-This suggests that i) the coded DVAE is aware that multiple latent vectors might be related to the image class and ii) that the model posterior shows large uncertainties (high entropy) for certain images for which the model has not properly identified the class.
-We illustrate this in Fig.
-3, where we show some images that were selected so that the MAP latent word from $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ induces class reconstruction errors.
-We display the reconstruction of the 4 most probable latent vectors and their corresponding probabilities.
-Observe that the uncoded model is confident no matter the reconstruction outcome while, in the coded posterior, the uncertainty is much larger.
-These results are also observed for MNIST, indicating that the posterior distribution in coded models exhibits a better uncertainty calibration.
-Note also that the increase in the number of latent bits (from 8 to 10) does not result in an excessive increase in the entropy despite the exponential growth of the number of vectors.
-4.2 Generation
-Model BER WER LL train LL test
-uncoded 5 0.051 0.195 -266.157 -267.703
-coded 5/50 0.011 0.046 -239.379 -241.882
-coded 5/80 0.008 0.039 -227.550 -232.992
-coded 5/100 0.010 0.049 -238.206 -241.404
-uncoded 8 0.089 0.384 -247.964 -249.880
-coded 8/80 0.021 0.144 -227.550 -232.992
-coded 8/160 0.027 0.189 -228.585 -235.819
-coded 8/240 0.037 0.231 -231.679 -238.459
-uncoded 10 0.142 0.622 -242.842 -244.997
-coded 10/100 0.040 0.321 -222.011 -230.772
-coded 10/200 0.044 0.341 -223.748 -234.849
-coded 10/300 0.045 0.349 -226.504 -238.647
-Figure 4: Evaluation of generation in FMNIST.
-The figure at the left shows an example of randomly generated, uncurated FMNIST images.
-The table at the right shows the quantitative results on the evaluation of the BER, WER, and log-likelihood (LL).
-In this section, we evaluate the model for the image generation task.
+Finally, we report the average entropy of the variational posterior $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$ over the test set in the table included in Fig.
+5.
+The low entropy in the uncoded models indicates low uncertainty when the model maps data points into the latent space, which could be beneficial if the model consistently assigns high probability to the correct latent vectors.
+However, the semantic accuracy results show that this is not the case for the uncoded model, as it often projects images into vectors corresponding to the wrong semantic class with high confidence.
+Figure 7: Reconstruction results for CIFAR10 (top) and Tiny ImageNet (bottom).
+Figure 8: Generation results for CIFAR10 (left) and Tiny ImageNet (right).
+Coded models, in contrast, improve semantic accuracy and exhibit higher entropy.
+This indicates that (i) the Coded-DVAE recognizes multiple latent vectors that might be related to a given semantic class, and (ii) the model’s posterior demonstrates significant uncertainty (high entropy) for images where the class is not well identified.
+This is illustrated in Fig.
+6, where we show images with class reconstruction errors caused by the MAP latent word from $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$.
+The reconstructions of the three most probable latent vectors, along with their respective probabilities, are displayed.
+Notably, the uncoded model shows high confidence regardless of the reconstruction result, while the coded posterior shows greater uncertainty.
+Note that this does not imply an uninformative distribution, as information bits remain recoverable.
+Additionally, increasing the number of latent bits does not lead to an excessive increase in entropy, despite the exponential growth in the number of vectors.
+5.3 Results on CIFAR10 and ImageNet
+Since MNIST-like datasets are relatively simple, it is challenging to fully assess the benefits of introducing ECCs in our model.
+To address this, we present additional results on more complex datasets, CIFAR10 and Tiny ImageNet, which feature colored images with more intricate shapes, patterns, and greater diversity.
+We trained both uncoded and coded models with different configurations to better understand the impact of the ECC.
+For context, the DVAE++ model [Vahdat et al., 2018b] required 128 binary latent variables to achieve state-of-the-art performance on these datasets, using a more complex model with Boltzmann Machine priors.
 In Fig.
-4, we show examples of randomly generated images using different model configurations in FMNIST.
-Results for the rest of the datasets are available in Appendices B, C, D, and E.
-These results are consistent with the ones obtained in reconstruction since we can observe that the coded models can generate more detailed and diverse images.
-Both uncoded and coded models generate more intricate and varied images with increased information bits.
-However, if the number of latent vectors becomes too large for the dataset’s complexity, not all words in the codebook are specialized during model training.
-This leads to generation artifacts, images where different classes of objects are overlapped.
-A visual inspection of Fig.
-4 suggests these artifacts are more frequent in the uncoded case.
-Note that, since we are dealing with discrete latent variables, we could simply detect and prune uninformative vectors.
-Accuracy metrics in generation.
-The improved inference given by the repetition code can also be tested by generating images using the generative model and counting errors using the MAP solution of the variational distribution $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})$.
-The table included in Fig.
-4 reports the BER and WER for FMNIST.
-As expected, at the same number of latent bits, the coded models significantly reduce both the BER and WER w.r.t.
-the uncoded case.
-Note also that the error rates grow with the number of latent bits, which is expected due to the increased complexity of the inference process.
-We may commit more errors by taking the MAP, but errors typically fall in consistent reconstructions (latent words that also reconstruct the same type of image), as the results presented in Section 4.1 indicated.
-Log-likelihood.
-We additionally estimated the log-likelihood (LL).
-Results for FMNIST with different model configurations, estimated through importance sampling with 300 samples per observation, are presented in the table included in Fig.
-4, please refer to Appendix I for further details.
-We observe that coded models consistently outperform their uncoded counterparts for both train and test sets, aligning with the results previously presented.
-Thus, we can argue that the introduction of repetition codes in the definition of the model allows for an improved inference and tighter posterior approximation.
-We draw similar conclusions for other datasets.
-We observe a general improvement in LL values as we increase the number of information bits, i.e., as we augment the latent dimensionality of the model and its flexibility.
-However, reducing the code rate does not lead to an improvement in log-likelihood.
-We argue that this might indicate overfitting of the decoder, as the LL deteriorates while reconstruction metrics improve.
-We must note that we use feed-forward networks at the decoder’s input.
-However, this may not be appropriate for the correlations we present in our coded words.
-We might overcome this overfitting tendency by using an architecture that properly leverages these coded bits correlations.
-Figure 5: Reconstruction results for CIFAR10 (left) and Tiny ImageNet (right).
-Figure 6: Generation results for CIFAR10 (left) and Tiny ImageNet (right).
-4.3 Additional results with CIFAR10 and Tiny ImageNet
-Since MNIST-like datasets are rather simple, it is difficult to assess the true gain in performance resulting from the introduction of ECCs proposed in our model.
-This section presents additional results using CIFAR10 and Tiny ImageNet, which contain colored images with more intricate shapes, patterns, and greater diversity than the previous datasets.
-We trained uncoded and coded models using different configurations to gain intuition regarding the effect of introducing the ECC.
-For a reference, in the case of the DVAE++ [30], the authors needed 128 binary latent variables to achieve state-of-the-art performance in generation and reconstruction for this dataset.
-They employed a more intricate model than the one introduced in this study, featuring Boltzmann Machine priors.
-In Fig.
-5 we show examples of reconstruction using different configurations of the model and in Fig.
-6 we show examples of randomly generated images.
-Additional results are provided in Appendices D and E.
-The results are consistent with those presented in previous sections; but in this case, the difference in performance is even more pronounced.
-We observe that the uncoded DVAE cannot decouple spatial information from the images and project it in the latent space.
-Nevertheless, the coded DVAE shows particular promise for learning low-dimensional discrete latent representations in complex datasets.
-Note that we used a rather simple architecture as we want to focus on the gain obtained only by introducing ECCs in the latent space.
+7, we show reconstruction examples, and in Fig.
+8, we present randomly generated images.
+Additional results are available in Appendices F and G.
+The results align with those from previous sections, but the performance difference is even more pronounced in this case.
+We find that the uncoded DVAE struggles to decouple spatial information in the images, while the Coded-DVAE shows particular promise for learning low-dimensional discrete latent representations, even for complex datasets.
+It is important to note that we used a simple architecture to isolate the gains achieved purely through the introduction of ECCs in the latent space, keeping a simple independent prior.
 6 Beyond repetition codes
@@
 We believe this opens a new path for designing latent probabilistic models with discrete latent variables.
-Although a detailed analysis of the joint design of ECC and encoder-decoder networks is beyond the scope of this work, we will outline key properties that any ECCs must satisfy to be integrated within this framework.
-  • Scalable hard encoding $\big{(}\boldsymbol{m}\rightarrow\boldsymbol{c}\big{)}$.
-Our model requires hard encoding for generation once the model is trained.
-This process should have linear complexity in $M$.
-  • Scalable soft encoding $\big{(}p(\boldsymbol{m})\rightarrow p(\boldsymbol{c})\big{)}$.
-Soft encoding is required during training for reparameterization.
-This process should also have linear complexity in $M$.
-  • Scalable soft decoding $\big{(}p(\boldsymbol{c})\rightarrow p(\boldsymbol{m})\big{)}$.
-Our model employs soft-in soft-out (SISO) decoding during inference.
-This process should again be linearly complex in $M$.
-  • Differentiability. Both encoding and decoding processes must be differentiable w.r.t the inputs to enable gradient computation and backpropagation.
-Figure 7: Graphical model of the hierarchical coded DVAE.
-Since Shannon’s seminal work [25], researchers have developed effective ECC schemes that meet these properties, including state-of-the-art ECCs such as Low Density Parity Check (LDPC) codes [8], or polar codes [2].
-Efforts have also focused on developing efficient SISO decoders, such as the sum-product algorithm [15].
-Inspired by polar codes [2], we present a hierarchical coded DVAE with two layers of latent bits.
+Although a detailed analysis of the joint design of ECC and encoder-decoder networks is beyond the scope of this work, in Appendix I we outline key properties that any ECCs must satisfy to be integrated within this framework.
+Drawing inspiration from polar codes [Arikan, 2009], we introduce a hierarchical Coded-DVAE with two layers.
+The graphical model is shown in Fig.
+9.
 In this model, the latent bits $\boldsymbol{m}_{1}$ are encoded using a repetition code in the first layer, producing $\boldsymbol{c}_{1}$ and $\boldsymbol{z}_{1}$.
 Simultaneously, the vector $\boldsymbol{m}_{2}$ is linearly combined with $\boldsymbol{m}_{1}$ using modulo 2 operations $\big{(}\boldsymbol{m}_{1}\oplus\boldsymbol{m}_{2}\big{)}$ and then encoded using a repetition code, yielding $\boldsymbol{c}_{2}$ and $\boldsymbol{z}_{2}$.
-Both soft vectors are concatenated and fed to the decoder NN to generate $\boldsymbol{x}$.
-The model provides stronger protection for $\boldsymbol{m}_{1}$, as it appears in both branches of the generative model.
-Inference follows a similar approach to the coded DVAE, incorporating the linear combination of $\boldsymbol{m}_{1}$ and $\boldsymbol{m}_{2}$ used in the second branch.
-This hierarchical structure allows the model to effectively separate high-level information from finer details, as we show in the results presented in Appendix G.
+Both soft vectors are concatenated and fed to the decoder to generate $\boldsymbol{x}$.
+Since $\boldsymbol{m}_{1}$ appears in both generative branches, it receives stronger protection.
+Inference follows a similar approach to the Coded-DVAE, incorporating the linear combination of $\boldsymbol{m}_{1}$ and $\boldsymbol{m}_{2}$ used in the second branch.
+This hierarchical structure allows the model to separate high-level information from finer details, as we show in the results presented in Appendix J.
+The bits in $\boldsymbol{m}_{1}$ offer stronger protection, allowing the model to encode the semantic information (type of clothing in FMNIST).
+As a result, during inference, it is more likely to project the image into a latent vector $\boldsymbol{m}_{1}$ that represents the correct semantic class.
+In contrast, $\boldsymbol{m}_{2}$ provides weaker protection and encodes image details that are less crucial for reconstruction.
+Fig.
+9 demonstrates that by keeping $\boldsymbol{m}_{1}$ fixed (semantic information) and sampling $\boldsymbol{m}_{2}$ randomly (fine details) we can generate diverse images within the same semantic class.
+Figure 9: Hierarchical Coded-DVAE. Graphical model (right) and generated images with a 5/100 code per branch (left). Here, $\boldsymbol{m}_{1}$ is fixed, and $\boldsymbol{m}_{2}$ is randomly sampled.
 7 Conclusion
@@
 By integrating redundancy into the latent space, the variational family can effectively refine the inference network’s output according to the structure of the ECC.
-Our findings underscore the efficacy of simple and efficient ECCs, like repetition codes, showcasing remarkable improvements over a lightweight version of the DVAE introduced in [30].
+Our findings underscore the efficacy of simple and efficient ECCs, like repetition codes, showcasing remarkable improvements over a lightweight version of the DVAE introduced in Vahdat et al.
+[2018b].
 Furthermore, our work reveals numerous avenues for future research.
-Firstly, investigating decoder architectures capable of efficiently utilizing the correlations and structure introduced by the ECCs, in contrast to the feed-forward networks employed in this study.
-We also contemplate exploring more complex and robust coding schemes, conducting theoretical analyses aligned with Shannon’s channel capacity and mutual information concepts to determine the fundamental parameters of the ECC needed to achieve reliable variational inference, exploring different modulations, and integrating these concepts into state-of-the-art models based on discrete representations.
-7 Acknowledgements
-The work was partially supported by the Ministerio de Ciencia, Innovación y Universidades, Agencia Estatal de Investigación, under grant PID2021-123182OB-I00, MCIN/AEI/10.13039/501100011033, co-financed by European Regional Development Fund (ERDF), “A way of making Europe”.
+Firstly, investigating architectures capable of efficiently utilizing the correlations and structure introduced by the ECCs, in contrast to the feed-forward networks employed in this study.
+We also contemplate exploring more complex and robust coding schemes, conducting theoretical analyses aligned with Shannon’s channel capacity and mutual information to determine the fundamental parameters of the ECC needed to achieve reliable inference, exploring different modulations, and integrating these concepts into state-of-the-art models based on discrete representations.
+Acknowledgements.
+María Martínez-García was supported by the Generación de Conocimiento grant PID2022-142506NA-I00.
+Grace Villacrés was supported by the Comunidad de Madrid within the 2023-2026 agreement with Universidad Rey Juan Carlos for the granting of direct subsidies for the promotion, encouragement of research and technology transfer, line of Action A Emerging Doctors, under Project OrdeNGN (Ref.
+F1177).
+David Mitchell was supported by the National Science Foundation under Grant No.
+CCF-2145917 and he acknowledges travel support from the European Union’s Horizon 2020 research and innovation program under Grant Agreement No.
+951847.
 Pablo M.
-Olmos also acknowledges the support by BBVA Foundation (Beca Leonardo a Investigadores y Creadores Culturales 2024) and by the Comunidad de Madrid under grants IND2022/TIC-23550 and ELLIS Unit Madrid.
-David Mitchell acknowledges support from the National Science Foundation under Grant No.
-CCF-2145917 and travel support from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 951847.
-Grace Villacrés acknowledges support from Comunidad de Madrid within the 2023-2026 agreement with Universidad Rey Juan Carlos for the granting of direct subsidies for the promotion and encouragement of research and technology transfer, Line of Action A: Emerging Doctors, under Project OrdeNGN.
-Appendix A Architecture
+Olmos was supported by the Comunidad de Madrid IND2022/TIC-23550, IDEA-CM project (TEC-2024/COM-89), the ELLIS Unit Madrid (European Laboratory for Learning and Intelligent Systems), the 2024 Leonardo Grant for Scientific Research and Cultural Creation from the BBVA Foundation, and by projects MICIU/AEI/10.13039/501100011033/FEDER and UE (PID2021-123182OB-I00; EPiCENTER).
+We would like to thank Isabel Valera for her valuable feedback on earlier drafts of the manuscript, as well as for the discussions that helped us improve the presentation of our method to the machine learning community.
+Appendix A Uncoded training algorithm
+The following pseudo-code describes the training process for the uncoded DVAE.
+It’s important to note that the main difference from the training of the Coded-DVAE lies in the fact that the encoder directly outputs $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{m}|\boldsymbol{x}_{i})$, which is used to sample $\boldsymbol{z}$.
+Therefore, we skip the soft decoding and coding steps.
+Algorithm 1 Training the model with uncoded inference.
+1:  Input: training data $\boldsymbol{x}_{i}$.
+2:  repeat
+3:     $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{m}|\boldsymbol{x}_{i})\leftarrow$ forward encoder $g_{\boldsymbol{\eta}}(\boldsymbol{x}_{i})$
+4:     $\boldsymbol{z}\leftarrow$ sample from (5)
+5:     $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})\leftarrow$ forward decoder $f_{\boldsymbol{\theta}}(\boldsymbol{z})$
+6:     Compute ELBO according to (3)
+7:     $\boldsymbol{\theta},\boldsymbol{\eta}\leftarrow Update(ELBO)$
+8:  until convergence
+Appendix B Coded training algorithm
+The following pseudo-code describes the training process for the Coded-DVAE.
+Here, we utilize soft decoding to leverage the added redundancy and retrieve the marginal posteriors of the information bits $\boldsymbol{m}$, correcting potential errors in $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}|\boldsymbol{x}_{i})$.
+We then apply the soft encoding technique to incorporate the structure of the code and sample $\boldsymbol{z}$ using the reparameterization trick as described in (5).
+Algorithm 2 Training the Coded-DVAE with repetition codes.
+1:  Input: training data $\boldsymbol{x}_{i}$, matrix $\mathbf{G}$.
+2:  repeat
+3:     $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}|\boldsymbol{x}_{i})\leftarrow$ forward encoder $g_{\boldsymbol{\eta}}(\boldsymbol{x}_{i})$
+4:     $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x}_{i})\leftarrow$ soft decoding by aggregating $q_{\boldsymbol{\eta}}^{u}(\boldsymbol{c}|\boldsymbol{x}_{i})$ according to (11)
+5:     $q_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x}_{i})\leftarrow$ repeat posterior bit probabilities $q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x}_{i})$ according to $\mathbf{G}$
+6:     $\boldsymbol{z}\leftarrow$ sample from (5)
+7:     $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})\leftarrow$ forward decoder $f_{\boldsymbol{\theta}}(\boldsymbol{z})$
+8:     Compute ELBO according to (3)
+9:     $\boldsymbol{\theta},\boldsymbol{\eta}\leftarrow Update(ELBO)$
+10:  until convergence
+Appendix C Architecture
 In this section, we detail the architecture used to obtain the experimental results with FMNIST and MNIST (28x28 gray-scale images).
-Note that across experiments we only modify the output layer of the encoder and the input layer of the decoder to adapt to the different configurations of the model.
+Note that, across the experiments, we only modify the output layer of the encoder and the input layer of the decoder to adapt to the different configurations of the model.
 This modification leads to a minimal alteration in the total number of parameters.
-In Section H.2, we conduct an ablation study on the number of trainable parameters to show that the enhancement in performance is not attributed to the increased dimensionality introduced by redundancy.
+In Section K.2, we conduct an ablation study on the number of trainable parameters to show that the enhancement in performance is not attributed to the increased dimensionality introduced by redundancy.
 For the additional CIFAR10 experiments, we change the input of the encoder and the output of the decoder to process the 32x32 color images.
@@
 The rest of the architecture remains unchanged.
-A.1 Encoder
+These architectures are comprehensively described in the following subsections.
+C.1 Encoder
 The encoder NN consists of 3 convolutional layers followed by two fully connected layers.
-We employed Leaky ReLU as the intermediate activation function and a Sigmoid as the output activation, as the encoder outputs bit probabilities.
+We employed Leaky ReLU as the intermediate activation function and a Sigmoid as the output activation, since the encoder outputs bit probabilities.
 The full architecture is detailed in Fig.
-8.
+10.
 Figure 10: Block diagram of the encoder architecture for FMNIST and MNIST.
-A.2 Decoder
-The decoder architecture is inspired by the one proposed in [24].
-It is composed of two fully connected layers, followed by transposed convolutional layers with residual connections and Squeeze-and-Excitation (SE) layers [10].
+C.2 Decoder
+The decoder architecture is inspired by the one proposed in Schuster and Krogh [2023].
+It is composed of two fully connected layers, followed by transposed convolutional layers with residual connections and Squeeze-and-Excitation (SE) layers [Hu et al., 2018].
 We employed Leaky ReLU as the intermediate activation function and a Sigmoid as output activation, given that we consider datasets with gray-scale images.
 The complete architecture is detailed in Fig.
-9.
+11.
 Figure 11: Block diagram of the decoder architecture for FMNIST and MNIST.
-Appendix B FMNIST results
+Appendix D FMNIST results
 In this section, we present supplementary results obtained with the FMNIST dataset.
-B.1 Training
+D.1 Training
 We present the evolution of the ELBO and its terms throughout the training process.
@@
 As we have observed in the different experiments, coded models are capable of generating more detailed images and accurate reconstructions.
-Introducing the repetition code also leads to smaller KL values, indicating that the posterior latent features are disentangled and less correlated.
+Introducing the repetition code also leads to smaller KL values, indicating that the posterior latent features are potentially disentangled and less correlated.
 We observe that, as we decrease the code rate, we obtain better bounds in general.
@@
 Figure 14: Evolution of the ELBO during training with 10 information bits on FMNIST.
-B.2 Reconstruction and generation
+D.2 Reconstruction and generation
 In this section, we augment the results presented in the main text, including outcomes obtained with the training dataset in Table 1.
 We include again the results obtained in the test to facilitate comparison.
+Additionally, we assess reconstruction quality using the Structural Similarity Index Measure (SSIM) to provide a more comprehensive evaluation.
 The results remain consistent across the two data partitions, and the analysis conducted for the test set also applies to training data.
-In all the cases, the coded models yield higher PSNR values than their uncoded counterparts, indicating a superior performance in reconstruction.
-We observe a general improvement in PSNR as we increase the number of information bits (i.e., as we augment the latent dimensionality of the model) and decrease the code rate (i.e., as we introduce more redundancy).
-As we discussed in the main text, the PSNR does not account for the semantic errors committed by the model.
+In all the cases, the coded models yield higher PSNR and SSIM values than their uncoded counterparts, indicating a superior performance in reconstruction.
+We observe a general improvement in PSNR and SSIM as we increase the number of information bits (i.e., as we augment the latent dimensionality of the model) and decrease the code rate (i.e., as we introduce more redundancy).
+As we discussed in the main text, the reconstruction metrics such as the PSNR and SSIM do not account for the semantic errors committed by the model.
 Therefore, we additionally report the semantic accuracy and the confident semantic accuracy.
@@
 We observe that coded models better capture the semantics of the images while employing the same number of latent vectors, significantly outperforming the uncoded models in terms of accuracy in all the cases.
-In Fig 13 we include additional examples of randomly generated images using different model configurations.
+Figure 15: Example of randomly generated, uncurated images using different model configurations.
+In Fig.
+15, we include additional examples of randomly generated images using different model configurations.
 We observe that coded models can generate more detailed and diverse images than their uncoded counterparts.
-Figure 13: Example of randomly generated, uncurated images using different model configurations.
 Table 1: Evaluation of reconstruction performance in FMNIST.
@@
 PSNR
+SSIM
 Acc
@@
 PSNR
+SSIM
 Acc
 Conf. Acc.
-uncoded 5 14.490 0.541 0.541 14.477 0.536 0.536
-coded 5/50 16.375 0.656 0.702 16.241 0.647 0.700
-coded 5/80 16.824 0.694 0.751 16.624 0.688 0.748
-coded 5/100 17.001 0.708 0.760 16.702 0.700 0.757
-uncoded 8 15.644 0.601 0.602 15.598 0.594 0.595
-coded 8/80 17.877 0.769 0.842 17.318 0.750 0.816
-coded 8/160 18.828 0.807 0.878 17.713 0.783 0.831
-coded 8/240 19.345 0.831 0.921 17.861 0.799 $\mathbf{0.893}$
-uncoded 10 16.053 0.650 0.652 16.000 0.644 0.648
-coded 10/100 18.827 0.813 0.885 17.694 0.790 0.850
-coded 10/200 19.937 0.846 0.897 18.009 0.814 0.871
-coded 10/300 $\mathbf{20.529}$ $\mathbf{0.855}$ $\mathbf{0.907}$ $\mathbf{18.111}$ $\mathbf{0.817}$ 0.870
-Appendix C MNIST results
+uncoded 5 14.490 0.438 0.541 0.541 14.477 0.437 0.536 0.536
+coded 5/50 16.375 0.559 0.656 0.702 16.241 0.551 0.647 0.700
+coded 5/80 16.824 0.585 0.694 0.751 16.624 0.574 0.688 0.748
+coded 5/100 17.001 0.596 0.708 0.760 16.702 0.580 0.700 0.757
+uncoded 8 15.644 0.513 0.601 0.602 15.598 0.509 0.594 0.595
+coded 8/80 17.877 0.647 0.769 0.842 17.318 0.619 0.750 0.816
+coded 8/160 18.828 0.690 0.807 0.878 17.713 0.641 0.783 0.831
+coded 8/240 19.345 0.713 0.831 0.921 17.861 0.653 0.799 $\mathbf{0.893}$
+uncoded 10 16.053 0.542 0.650 0.652 16.000 0.536 0.644 0.648
+coded 10/100 18.827 0.690 0.813 0.885 17.694 0.639 0.790 0.850
+coded 10/200 19.937 0.735 0.846 0.897 18.009 0.659 0.814 0.871
+coded 10/300 $\mathbf{20.529}$ $\mathbf{0.754}$ $\mathbf{0.855}$ $\mathbf{0.907}$ $\mathbf{18.111}$ $\mathbf{0.662}$ $\mathbf{0.817}$ 0.870
+Appendix E MNIST results
 In this section, we report the results obtained with the MNIST dataset.
-C.1 Training
+E.1 Training
 We present the evolution of the ELBO and its terms throughout the training process.
@@
 The colors in all plots represent the various code rates.
+Figure 16: Evolution of the ELBO during training with 5 information bits on MNIST.
+Figure 17: Evolution of the ELBO during training with 8 information bits on MNIST.
+Figure 18: Evolution of the ELBO during training with 10 information bits on MNIST.
 The results are consistent with the ones obtained for FMNIST.
@@
 However, the introduction of ECCs in the model allows for latent spaces that better capture the structure of the images while employing the same number of latent vectors.
-Figure 14: Evolution of the ELBO during training with 5 information bits on MNIST.
-Figure 15: Evolution of the ELBO during training with 8 information bits on MNIST.
-Figure 16: Evolution of the ELBO during training with 10 information bits on MNIST.
-C.2 Reconstruction
+E.2 Reconstruction
 We first evaluate the model’s performance in reconstructing data by examining its uncoded and coded versions across different configurations, varying the number of information bits and code rates.
 All the results obtained with MNIST are consistent with those presented in the main text for FMNIST.
-In Table 2 we quantify the quality of the reconstructions measuring the PSNR in both training and test sets.
-In all the cases, coded models yield higher PSNR values, indicating a superior performance in reconstruction.
+In Table 2 we quantify the quality of the reconstructions measuring the PSNR and the SSIM in both training and test sets.
+In all the cases, coded models yield higher PSNR and SSIM values, indicating a superior performance in reconstruction.
 This improvement is also evident through visual inspection of Fig.
-17, where the coded models better capture the details in the images.
-As in FMNIST, we observe a general improvement of the PSNR as we increase the number of information bits and decrease the code rate.
-As we discussed in the main text, the PSNR does not account for the semantic errors committed by the model.
-Therefore, we additionally evaluate the reconstruction accuracy, ensuring that the model successfully reconstructs images within the same class as the original ones.
-We also provide a confident reconstruction accuracy, for which we do not count errors when the MAP value of $q(\boldsymbol{m}|\boldsymbol{x})$ is below 0.4.
-In light of the results, we argue that introducing ECCs in the model allows for latent spaces that better capture the semantics of the images while employing the same number of latent vectors, outperforming the uncoded models in all the cases.
+19, where the coded models better capture the details in the images.
+Similar to the results observed on FMNIST, we find that both PSNR and SSIM generally improve as the number of information bits increases and the code rate decreases.
 Figure 19: Example of reconstructed test images obtained with different model configurations.
@@
 PSNR
+SSIM
 Acc
@@
 PSNR
+SSIM
 Acc
@@
 Entropy
-uncoded 5 13.483 0.702 0.703 13.483 0.701 0.702 0.277
-coded 5/50 14.983 0.887 0.923 14.888 0.887 0.920 2.073
-coded 5/80 15.436 0.899 0.936 15.263 0.895 0.929 2.237
-coded 5/100 15.590 0.905 0.931 15.352 0.898 0.924 2.382
-uncoded 8 14.530 0.860 0.864 14.490 0.860 0.868 0.513
-coded 8/80 16.878 0.937 0.964 16.042 0.912 0.947 3.105
-coded 8/160 18.108 0.957 0.974 16.497 0.927 0.951 3.645
-coded 8/240 19.984 0.967 0.978 16.688 0.936 0.957 3.881
-uncoded 10 14.879 0.888 0.891 14.816 0.887 0.890 0.636
-coded 10/100 17.584 0.945 0.972 16.795 0.928 $\mathbf{0.968}$ 4.080
-coded 10/200 20.060 0.973 0.977 16.863 0.932 0.944 4.411
-coded 10/300 $\mathbf{21.083}$ $\mathbf{0.979}$ $\mathbf{0.984}$ $\mathbf{17.114}$ $\mathbf{0.941}$ 0.945 4.810
+uncoded 5 13.491 0.434 0.702 0.703 13.483 0.430 0.701 0.702 0.277
+coded 5/50 14.983 0.601 0.887 0.923 14.888 0.598 0.887 0.920 2.073
+coded 5/80 15.436 0.641 0.899 0.936 15.263 0.634 0.895 0.929 2.237
+coded 5/100 15.590 0.652 0.905 0.931 15.352 0.641 0.898 0.924 2.382
+uncoded 8 14.530 0.558 0.860 0.864 14.490 0.558 0.860 0.868 0.513
+coded 8/80 16.878 0.739 0.937 0.964 16.042 0.699 0.912 0.947 3.105
+coded 8/160 18.108 0.802 0.957 0.974 16.497 0.736 0.927 0.951 3.645
+coded 8/240 19.984 0.838 0.967 0.978 16.688 0.752 0.936 0.957 3.881
+uncoded 10 14.879 0.591 0.888 0.891 14.816 0.589 0.887 0.890 0.636
+coded 10/100 17.584 0.777 0.945 0.972 16.795 0.744 0.928 $\mathbf{0.968}$ 4.080
+coded 10/200 20.060 0.875 0.973 0.977 16.863 0.765 0.932 0.944 4.411
+coded 10/300 $\mathbf{21.083}$ $\mathbf{0.902}$ $\mathbf{0.979}$ $\mathbf{0.984}$ $\mathbf{17.114}$ $\mathbf{0.781}$ $\mathbf{0.941}$ 0.945 4.810
+As we discussed in the main text, reconstruction metrics such as PSNR and SSIM do not account for the semantic errors committed by the model.
+Therefore, we additionally evaluate the reconstruction accuracy, ensuring that the model successfully reconstructs images within the same class as the original ones.
+We also provide a confident reconstruction accuracy, for which we do not count errors when the MAP value of $q(\boldsymbol{m}|\boldsymbol{x})$ is below 0.4.
+In light of the results, we argue that introducing ECCs in the model allows for latent spaces that better capture the semantics of the images while employing the same number of latent vectors, outperforming the uncoded models in all the cases.
 We also report the average entropy of the variational posterior over the test set in Table 2.
-If we analyze the entropy together with the semantic accuracy, we can argue that coded VAE is aware that multiple vectors might be related to the same image class, and that the posterior shows larger uncertainties for images for which the model has not properly identified the class.
+If we consider the entropy together with the semantic accuracy, we can argue that coded VAE is aware that multiple vectors might be related to the same image class, and that the posterior shows larger uncertainties for images for which the model has not properly identified the class.
 We illustrate this argument in Fig.
-18, where we show some images selected so that the MAP latent word of $q(\boldsymbol{m}|\boldsymbol{x})$ induces class reconstruction errors.
+20, where we show some images selected so that the MAP latent word of $q(\boldsymbol{m}|\boldsymbol{x})$ induces class reconstruction errors.
 We show the reconstruction of the 4 most probable latent vectors and their corresponding probabilities.
@@
 Coded models consistently outperform their uncoded counterparts for both the training and test sets, consistent with the findings observed using the FMNIST dataset.
-C.3 Generation
+E.3 Generation
 In this section, we evaluate the model in the image generation task.
 In Fig.
-19, we show examples of randomly generated images using different model configurations in MNIST.
-These results are consistent with the ones obtained in reconstruction, and with the ones obtained for FMNIST, as we observe that the coded models can generate more detailed and diverse images.
-The improved inference provided by the repetition code can also be tested by generating images using the generative model and counting errors using the MAP solution of the variational posterior distribution.
+21, we show examples of randomly generated images using different model configurations in MNIST.
+These results are consistent with the ones obtained in reconstruction, and with the ones obtained for FMNIST, where we observe that the coded models can generate more detailed and diverse images.
+The improved inference provided by the repetition code can also be tested by generating images using the generative model and counting errors using the MAP solution of the variational posterior distribution, as well as sampling from the approximate posterior distribution.
 Table 4 reports the BER and WER for MNIST.
@@
 Table 3: Evaluation of the BER, WER in MNIST.
-Model BER WER
-uncoded 5 0.002 0.008
-coded 5/50 0.007 0.034
-coded 5/80 0.004 0.021
-coded 5/100 0.009 0.045
-uncoded 8 0.015 0.071
-coded 8/80 0.020 0.147
-coded 8/160 0.021 0.160
-coded 8/240 0.023 0.167
-uncoded 10 0.057 0.373
-coded 10/100 0.030 0.258
-coded 10/200 0.034 0.282
-coded 10/300 0.041 0.331
-Table 4: Evaluation of the log-likelihood (LL) in MNIST.
+Model BER BER MAP WER WER MAP
+uncoded 5 0.002 0.001 0.008 0.004
+coded 5/50 0.007 0.000 0.034 0.000
+coded 5/80 0.004 0.000 0.021 0.000
+coded 5/100 0.009 0.000 0.045 0.000
+uncoded 8 0.015 0.013 0.071 0.057
+[…]
+[…]
+[…]
+uncoded 10 0.057 0.052 0.373 0.353
+[…]
+[…]
+[…]
+Table 4: Evaluation of the LL in MNIST.
 Model LL (train) LL (test)
@@
 coded 10/300 -88.904 -110.799
-Appendix D CIFAR10 results
+Appendix F CIFAR10 results
 In this section, we provide additional results using the CIFAR10 dataset with different model configurations.
-D.1 Training
+F.1 Training
 We present the evolution of the ELBO and its terms throughout the training process.
@@
 The colors in all plots represent the various code rates.
+Figure 22: Evolution of the ELBO during training with 70 information bits on CIFAR10.
+Figure 23: Evolution of the ELBO during training with 100 information bits on CIFAR10.
+Figure 24: Evolution of the ELBO during training with 130 information bits on CIFAR10.
 Across all the configurations, coded models achieve superior bounds.
@@
 As we have observed across the different experiments, coded models are capable of better capturing the structure of the data, generating more detailed images and accurate reconstructions.
-Figure 20: Evolution of the ELBO during training with 70 information bits on CIFAR10.
-Figure 21: Evolution of the ELBO during training with 100 information bits on CIFAR10.
-Figure 22: Evolution of the ELBO during training with 130 information bits on CIFAR10.
 In the coded case, we do not observe significant differences in the obtained bounds as we increase the number of information bits and reduce the code rate.
@@
 Utilizing an architecture capable of effectively leveraging these correlations among the coded bits could potentially enable us to better exploit the introduced redundancy.
-D.2 Reconstruction and generation
+F.2 Reconstruction and generation
 We first evaluate the model’s performance in reconstructing data by examining its uncoded and coded versions across different configurations, varying the number of information bits and code rates.
-In Table 6 we quantify the quality of the reconstructions measuring the PSNR in both training and test sets.
-Coded models yield higher PSNR values in train, and similar values in test, although the coded models with lower rates outperform the rest of the configurations.
+In Table 6 we quantify the quality of the reconstructions measuring the PSNR and the SSIM in both training and test sets.
+Coded models yield higher PSNR and SSIM values in train, and similar values in test, although the coded models with lower rates outperform the rest of the configurations.
 However, the improvement in reconstruction is evident through visual inspection of Fig.
-23, where the coded models better capture the details in the images.
-We observe that the coded model yields images that better resemble the structure of the dataset, while the uncoded DVAE cannot decouple spatial information from the images and project it in the latent space.
-We hypothesize that to adequately model complex images, transitioning to a hierarchical structure may be necessary.
-This would allow for the explicit modeling of both global and local information.
+25, where the coded models better capture the details in the images.
+We observe that the coded model produces images that better resemble the structure of the dataset, while the uncoded DVAE cannot decouple spatial information from the images and project it in the latent space.
+We hypothesize that, to adequately model complex images, transitioning to a hierarchical structure may be necessary.
+This would allow for explicit modeling of both global and local information.
 However, despite employing this rather simple model, we observe that coded configurations outperform their uncoded counterparts in capturing colors and textures.
@@
 In Fig.
-24, we show examples of randomly generated images using different model configurations in CIFAR10.
+26, we show examples of randomly generated images using different model configurations in CIFAR10.
 These results are consistent with the ones obtained in reconstruction, as we observe that the coded models can generate more detailed and diverse images.
@@
 For this, we used the implementation available at https://github.com/mseitzer/pytorch-fid.
-We can observe that the coded models significantly reduced the FID score in all the cases compared to their uncoded counterparts.For the coded models, we do not observe a clear influence of the code rate on the quality of the generations.
+We can observe that the coded models significantly reduced the FID score in all the cases compared to their uncoded counterparts.
+For the coded models, we do not observe a clear influence of the code rate on the quality of the generations.
 Figure 25: Example of reconstructed test images obtained with different model configurations. Observe that more details are visualized as we increase bits in the latent space and introduce redundancy.
 Figure 26: Example of randomly generated, uncurated images using different model configurations.
-Table 5: Evaluation of reconstruction performance in CIFAR10 with different model configurations.
-Model PSNR (train) PSNR (test)
-uncoded 70 17.985 17.596
-coded 70/700 23.790 17.731
-coded 70/1400 24.555 18.008
-coded 70/2100 25.551 18.401
-uncoded 100 18.509 18.334
-coded 100/1000 24.754 18.229
-coded 100/2000 24.866 18.927
-coded 100/3000 25.646 18.920
-uncoded 130 18.951 18.758
-coded 130/1300 25.007 18.887
-coded 130/2600 25.460 19.416
-coded 130/3900 $\mathbf{25.515}$ $\mathbf{19.292}$
+Table 5: Reconstruction metrics in CIFAR10 with different model configurations.
+Model
+PSNR
+SSIM
+PSNR
+SSIM
+uncoded 70 17.985 0.340 17.596 0.324
+coded 70/700 23.790 0.709 17.731 0.323
+coded 70/1400 24.555 0.748 18.008 0.357
+coded 70/2100 25.551 0.748 18.401 0.385
+uncoded 100 18.509 0.381 18.334 0.370
+coded 100/1000 24.754 0.754 18.229 0.375
+coded 100/2000 24.866 0.761 18.927 0.432
+coded 100/3000 25.646 0.793 18.920 0.426
+uncoded 130 18.951 0.419 18.758 .408
+coded 130/1300 25.007 0.767 18.887 0.426
+coded 130/2600 25.460 0.784 $\mathbf{19.416}$ $\mathbf{0.464}$
+coded 130/3900 $\mathbf{25.515}$ $\mathbf{0.785}$ 19.292 0.456
 Table 6: Evaluation of the BER, WER, and FID in CIFAR10 with different model configurations.
-Model BER WER FID
-uncoded 70 0.162 1.000 177.524
-coded 70/700 0.101 1.000 104.977
-coded 70/1400 0.088 0.999 104.078
-coded 70/2100 0.090 0.999 102.795
-uncoded 100 0.182 1.000 172.063
-coded 100/1000 0.123 1.000 107.887
-coded 100/2000 0.114 1.000 101.182
-coded 100/3000 0.138 1.000 107.287
-uncoded 130 0.197 1.000 164.138
-coded 130/1300 0.144 1.000 109.905
-coded 130/2600 0.164 1.000 110.250
-coded 130/3900 0.185 1.000 108.561
-Appendix E Tiny ImageNet results
+Model BER
+BER
+MAP
+WER
+WER
+MAP
+FID
+uncoded 70 0.162 0.162 1.000 0.999 177.524
+coded 70/700 0.101 0.081 1.000 0.988 104.977
+coded 70/1400 0.088 0.044 0.999 0.917 104.078
+coded 70/2100 0.090 0.029 0.999 0.811 102.795
+uncoded 100 0.182 0.179 1.000 1.000 172.063
+coded 100/1000 0.123 0.104 1.000 0.999 107.887
+coded 100/2000 0.114 0.060 1.000 0.990 101.182
+coded 100/3000 0.138 0.077 1.000 0.997 107.287
+uncoded 130 0.197 0.194 1.000 1.000 164.138
+coded 130/1300 0.144 0.115 1.000 0.999 109.905
+coded 130/2600 0.164 0.102 1.000 1.000 110.250
+coded 130/3900 0.185 0.132 1.000 1.000 108.561
+Appendix G Tiny ImageNet results
 In this section, we provide additional results using the Tiny ImageNet dataset with different model configurations.
-E.1 Training
+G.1 Training
 We present the evolution of the ELBO and its terms throughout the training process.
@@
 As we have observed across the different experiments, coded models are capable of better capturing the structure of the data, generating more detailed images and accurate reconstructions.
+Figure 27: Evolution of the ELBO during training for the configurations with 70 information bits.
+Figure 28: Evolution of the ELBO during training for the configurations with 100 information bits.
+Figure 29: Evolution of the ELBO during training for the configurations with 130 information bits.
 In the coded case, we do not observe significant differences in the obtained bounds as we increase the number of information bits and reduce the code rate.
@@
 However, the introduction of ECCs in the model allows for latent spaces that better capture the structure of the images while employing the same number of latent vectors.
-Figure 25: Evolution of the ELBO during training for the configurations with 70 information bits.
-Figure 26: Evolution of the ELBO during training for the configurations with 100 information bits.
-Figure 27: Evolution of the ELBO during training for the configurations with 130 information bits.
 It’s important to note that we are currently using feed-forward networks at the decoder’s input.
@@
 Utilizing an architecture capable of effectively leveraging these correlations among the coded bits could potentially enable us to better exploit the introduced redundancy.
-E.2 Reconstruction and generation
+G.2 Reconstruction and generation
+Figure 30: Example of reconstructed test images obtained with different model configurations.
+Observe that more details are visualized as we increase the bits in the latent space and introduce redundancy.
 We first evaluate the model’s performance in reconstructing data by examining its uncoded and coded versions across different configurations, varying the number of information bits and code rates.
-In Table 8 we quantify the quality of the reconstructions measuring the PSNR in both training and test sets.
-Coded models yield higher PSNR values in train, and similar values in test, although the coded models with lower rates outperform the rest of the configurations.
+In Table 8, we quantify the quality of the reconstructions measuring the PSNR and the SSIM in both training and test sets.
+Coded models yield higher PSNR and SSIM values in train, and similar values in test, although coded models with lower rates outperform the rest of the configurations.
 However, the improvement in reconstruction is evident through visual inspection of Fig.
-28, where the coded models better capture the details in the images.
-We observe that the coded model yields images that better resemble the structure of the dataset, while the uncoded DVAE cannot decouple spatial information from the images and project it in the latent space.
-We hypothesize that to adequately model complex images, transitioning to a hierarchical structure may be necessary.
+30, where the coded models better capture the details in the images.
+We observe that the coded model produces images that better resemble the structure of the dataset, while the uncoded DVAE cannot decouple spatial information from the images and project it in the latent space.
+We hypothesize that, to adequately model complex images, transitioning to a hierarchical structure may be necessary.
 This would allow for the explicit modeling of both global and local information.
 However, despite employing this rather simple model, we observe that coded configurations outperform their uncoded counterparts in capturing colors and textures.
-Figure 28: Example of reconstructed test images obtained with different model configurations.
-Observe that more details are visualized as we increase the bits in the latent space and introduce redundancy.
-Figure 29: Example of randomly generated, uncurated images using different model configurations.
-Table 7: Evaluation of reconstruction performance in Tiny ImageNet with different model configurations.
-Model PSNR (train) PSNR (test)
-uncoded 70 15.598 15.402
-coded 70/700 18.156 15.158
-coded 70/1400 18.396 15.419
-coded 70/2100 18.228 15.789
-uncoded 100 16.012 15.774
-coded 100/1000 18.298 15.677
-coded 100/2000 18.647 15.892
-coded 100/3000 18.729 16.167
-uncoded 130 16.278 16.009
-coded 130/1300 18.719 15.901
-coded 130/2600 18.818 $\mathbf{16.329}$
-coded 130/3900 $\mathbf{19.020}$ 16.288
-Table 8: Evaluation of the BER, WER, and FID in Tiny ImageNet with different model configurations.
-Model BER WER FID
-uncoded 70 0.143 1.000 265.474
-coded 70/700 0.096 0.998 171.993
-coded 70/1400 0.104 1.000 170.496
-coded 70/2100 0.096 0.998 176.245
-uncoded 100 0.164 1.000 234.358
-coded 100/1000 0.099 1.000 153.743
-coded 100/2000 0.097 1.000 162.889
-coded 100/3000 0.098 1.000 163.049
-uncoded 130 0.200 1.000 219.003
-coded 130/1300 0.129 1.000 165.064
-coded 130/2600 0.114 1.000 164.759
-coded 130/3900 0.128 1.000 170.603
 We also evaluate the model in the image generation task.
 In Fig.
-29, we show examples of randomly generated images using different model configurations in Tiny ImageNet.
-These results are consistent with the ones obtained in reconstruction, as we observe that the coded models can generate more detailed and diverse images.
+31, we show examples of randomly generated images using different model configurations in Tiny ImageNet.
+These results are consistent with the ones obtained in reconstruction, where we observe that the coded models can generate more detailed and diverse images.
 Additionally, we obtained the FID score using the test set and 10k generated samples.
@@
 Our experiments indicate that coded models with lower rates attempt to model fine details in images, which can lead to artifacts in generation.
-Appendix F IWAE results
-One could draw a parallel between the coded DVAE with repetition codes and the well-known IWAE [3], but the two approaches are fundamentally different.
+Figure 31: Example of randomly generated, uncurated images using different model configurations.
+Table 7: Reconstruction metrics in Tiny ImageNet with different model configurations.
+Model
+PSNR
+SSIM
+PSNR
+SSIM
+uncoded 70 15.598 0.207 15.402 0.196
+coded 70/700 18.156 0.367 15.158 0.191
+coded 70/1400 18.396 0.383 15.419 0.203
+coded 70/2100 18.228 0.369 15.789 0.223
+uncoded 100 16.012 0.229 15.774 0.215
+coded 100/1000 18.298 0.378 15.677 0.218
+coded 100/2000 18.647 0.404 15.892 0.232
+coded 100/3000 18.729 0.408 16.167 0.248
+uncoded 130 16.278 0.243 16.009 0.228
+coded 130/1300 18.719 0.409 15.900 0.233
+coded 130/2600 18.818 $\mathbf{0.433}$ $\mathbf{16.329}$ $\mathbf{0.259}$
+coded 130/3900 $\mathbf{19.020}$ 0.415 16.288 $\mathbf{0.259}$
+Table 8: Evaluation of the BER, WER, and FID in Tiny ImageNet with different model configurations.
+Model BER
+BER
+MAP
+WER
+WER
+MAP
+FID
+uncoded 70 0.143 0.140 1.000 0.999 265.474
+coded 70/700 0.096 0.072 0.998 0.978 171.993
+coded 70/1400 0.104 0.066 1.000 0.972 170.496
+coded 70/2100 0.096 0.034 0.998 0.832 176.245
+uncoded 100 0.164 0.162 1.000 1.000 234.358
+coded 100/1000 0.099 0.074 1.000 0.996 153.743
+coded 100/2000 0.097 0.053 1.000 0.981 162.889
+coded 100/3000 0.098 0.035 1.000 0.925 163.049
+uncoded 130 0.200 0.198 1.000 1.000 219.003
+coded 130/1300 0.129 0.107 1.000 0.999 165.064
+coded 130/2600 0.114 0.060 1.000 0.996 164.759
+coded 130/3900 0.128 0.070 1.000 0.999 170.603
+Appendix H IWAE results
+One could draw a parallel between the Coded-DVAE with repetition codes and the well-known IWAE [Burda et al., 2016], but the two approaches are fundamentally different.
 In the IWAE, independent samples are drawn from the variational posterior and propagated independently through the generative model to obtain a tighter variational bound on the marginal log-likelihood.
@@
 However, our approach extends beyond repetition codes, opening a new field for improved inference in discrete latent variable models.
-In this work, we specifically utilize the redundancy introduced by the repetition code to correct potential errors made by the encoder through a soft decoding approach, leading to a more accurate approximation of $p(\boldsymbol{m}|\boldsymbol{x})$ and an improved proposal for sampling.
-The results obtained in the coded DVAE case cannot be achieved by training the uncoded DVAE with the IWAE objective.
-The following results compare the uncoded IWAE model with the coded DVAE trained on FMNIST.
-While the uncoded model shows slight performance gains with an increasing number of IWAE samples (which improves the evidence lower bound), it still underperforms compared to the coded model.
-Furthermore, when using 20 and 30 IWAE samples, the metrics slightly declined compared to using 10 samples, likely due to overfitting, as we applied a common early stopping point.
+In our approach, we specifically utilize the redundancy introduced by the repetition code to correct potential errors made by the encoder through a soft decoding approach.
+This results in a more accurate approximation of the posterior $p(\boldsymbol{m}|\boldsymbol{x})$ and an improved proposal for sampling, resulting in improved performance over uncoded models, even those trained with the tighter IWAE bound.
+To compute the IWAE objective in the uncoded case, we draw samples from the posterior using the reparameterization trick described in Eq.
+(5), and compute the importance weights as $\boldsymbol{w}=\frac{p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})p(\boldsymbol{z})}{q_{\boldsymbol{\eta}}(\boldsymbol{z}|\boldsymbol{x})}$.
+Here, the prior and posterior over $\boldsymbol{z}$ are obtained by marginalizing out $\boldsymbol{m}$ in $p(\boldsymbol{z},\boldsymbol{m})=p(\boldsymbol{m})p(\boldsymbol{z}|\boldsymbol{m})$ and $q_{\boldsymbol{\eta}}(\boldsymbol{m},\boldsymbol{z}|\boldsymbol{x})=q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})p(\boldsymbol{z}|\boldsymbol{m})$, respectively.
+Additionally, we implemented two extensions of the IWAE bound to confirm that the observed improvements are not due to a known shortcoming of the reparametrized gradient estimator of the IWAE bound.
+Following Daudel et al.
+[2023], we implemented the VR-IWAE generalization, which introduces a hyperparameter $\alpha\in[0,1)$ into the objective function.
+This formulation interpolates between the IWAE bound (recovered when $\alpha=0$) and the ELBO (recovered when $\alpha=1$).
+For our experiments, we selected a midpoint value of $\alpha=0.5$.
+Additionally, we implemented the doubly-reparametrized (DR) gradient estimator associated with the VR-IWAE objective, as described in Daudel et al.
+[2023].
+We trained the models on FMNIST using 10, 20, and 30 posterior samples, matching the number of repetitions used in our coded models.
+We included an additional experiment training an uncoded model with the IWAE bound considering 100 samples.
+From the results, outlined in Table 9, we observe that training uncoded models with tighter bounds (even if we use bounds that are meant to become tighter as the number of samples increases) does not lead to substantial improvements in performance.
+These models consistently underperform relative to their coded counterparts, which are trained using the ELBO as the objective function.
+For a fair comparison, we use the same neural network architecture for the uncoded and the coded posterior (differing only in the final dense layer to match the respective output dimensionalities), the same network architecture for the decoder (differing only in the first dense layer to match the respective input dimensionalities), and the same simple independent prior.
+Fig.
+32 shows the evolution of the different objectives, with coded models obtaining better bounds than the uncoded IWAE models, even when considering 100 samples.
+In Table 9, we present test metrics, with coded models again overperforming their uncoded counterparts trained with the IWAE objective.
+We observe that some metrics slightly decline augmenting the number of samples, likely due to overfitting, since we applied a common early stopping point.
 Figure 32: Evolution of the ELBO and the IWAE objectives for various configurations.
@@
 However, coded models obtain even better bounds using the same number of samples/repetitions.
-Table 9: Comparison of the metrics obtained with our method and the uncoded DVAE trained with the IWAE objective.
-Model BER WER Entropy Acc. Conf. Acc. PSNR
-uncoded 8 0.089 0.384 0.467 0.594 0.595 15.598
-uncoded 8 IWAE 10 samples 0.063 0.372 1.309 0.617 0.640 14.282
-uncoded 8 IWAE 20 samples 0.075 0.447 1.391 0.634 0.651 14.237
-uncoded 8 IWAE 30 samples 0.074 0.438 1.564 0.619 0.641 13.757
-coded 8/80 0.021 0.144 2.905 0.750 0.816 17.318
-coded 8/160 0.027 0.189 3.637 0.783 0.831 17.713
-coded 8/240 0.037 0.231 4.000 0.799 0.893 17.861
-Appendix G Hierarchical coded DVAE results
-Inspired by polar codes [2], we present a hierarchical coded DVAE with two layers of latent bits, as illustrated in Figure 31.
+Table 9: Comparison of test metrics between our method and the uncoded DVAE trained with the IWAE objective.
+Model BER WER Entropy Acc. Conf. Acc. PSNR SSIM
+uncoded 8 0.089 0.384 0.467 0.594 0.595 15.598 0.509
+uncoded 8 IWAE 10 samples 0.063 0.372 1.309 0.617 0.640 14.282 0.470
+uncoded 8 IWAE 20 samples 0.075 0.447 1.391 0.634 0.651 14.237 0.460
+uncoded 8 IWAE 30 samples 0.074 0.438 1.564 0.619 0.641 13.757 0.457
+uncoded 8 IWAE 100 samples 0.107 0.583 2.274 0.596 0.663 12.996 0.433
+uncoded 8 IWAE ($\alpha=0.5$, 10 samples) 0.059 0.353 0.929 0.630 0.638 14.582 0.488
+uncoded 8 IWAE ($\alpha=0.5$, 20 samples) 0.067 0.388 1.066 0.638 0.644 14.432 0.479
+uncoded 8 IWAE ($\alpha=0.5$, 30 samples) 0.053 0.349 1.165 0.630 0.644 14.178 0.472
+uncoded 8 IWAE (DR, $\alpha=0.5$, 10 samples) 0.052 0.312 1.004 0.623 0.629 14.542 0.483
+uncoded 8 IWAE (DR, $\alpha=0.5$, 20 samples) 0.056 0.353 1.041 0.600 0.614 14.231 0.463
+uncoded 8 IWAE (DR, $\alpha=0.5$, 30 samples) 0.057 0.352 1.068 0.602 0.607 14.016 0.449
+coded 8/80 0.021 0.144 2.905 0.750 0.816 17.318 0.619
+coded 8/160 0.027 0.189 3.637 0.783 0.831 17.713 0.641
+coded 8/240 0.037 0.231 4.000 0.799 0.893 17.861 0.653
+Appendix I Beyond Repetition Codes
+We have presented compelling proof-of-concept results that incorporating ECCs, like repetition codes, into DVAEs can improve performance.
+We believe this opens a new path for designing latent probabilistic models with discrete latent variables.
+Although a detailed analysis of the joint design of ECC and encoder-decoder networks is beyond the scope of this work, we will outline key properties that any ECCs must satisfy to be integrated within this framework.
+  • Scalable hard encoding $\big{(}\boldsymbol{m}\rightarrow\boldsymbol{c}\big{)}$.
+Our model requires hard encoding for generation once the model is trained.
+This process should have linear complexity in $M$.
+  • Scalable soft encoding $\big{(}p(\boldsymbol{m})\rightarrow p(\boldsymbol{c})\big{)}$.
+Soft encoding is required during training for reparameterization.
+This process should also have linear complexity in $M$.
+  • Scalable soft decoding $\big{(}p(\boldsymbol{c})\rightarrow p(\boldsymbol{m})\big{)}$.
+Our model employs soft-in soft-out (SISO) decoding during inference.
+This process should again be linearly complex in $M$.
+  • Differentiability. Both encoding and decoding processes must be differentiable w.r.t. the inputs to enable gradient computation and backpropagation.
+Since Shannon’s landmark work [Shannon, 1948], researchers have been developing feasible ECC schemes that meet (or approach) the aforementioned properties for ever-increasing values of $M$ and $R$.
+This is driven by the high throughput demands in digital communications and limited power constraints in the case of wireless communications.
+In this regard, state-of-the-art ECC schemes such as Low Density Parity Check (LDPC) codes [Gallager, 1962] can be designed to meet the above criteria, leveraging their linear algebraic structure.
+Similarly, encoding and decoding algorithms for polar codes [Arikan, 2009] have complexity of $\mathcal{O}(D\log D)$, approaching this target.
+Additionally, in digital communications, soft decision decoding is known to provide savings of up to 2-3 dB of signal energy over hard decision decoding (binary input).
+For this reason, significant effort has been also made in the communications field to develop efficient and powerful SISO decoders, such as the sum-product algorithm (SPA) [Kschischang et al., 2001] for LDPC codes.
+Appendix J Hierarchical Coded-DVAE results
+Figure 33: Graphical model of the hierarchical Coded VAE with two layers.
+Inspired by polar codes [Arikan, 2009], we present a hierarchical Coded-DVAE with two layers of latent bits, as illustrated in Fig.
+33.
 In this model, the latent bits $\boldsymbol{m}_{1}$ are encoded using a repetition code in the first layer, producing $\boldsymbol{c}_{1}$ and $\boldsymbol{z}_{1}$.
@@
 The model provides stronger protection for $\boldsymbol{m}_{1}$, as it appears in both branches of the generative model.
-Inference follows a similar approach to the one employed in the coded DVAE, incorporating the linear combination of $\boldsymbol{m}_{1}$ and $\boldsymbol{m}_{2}$ used in the second branch.
-Figure 31: Graphical model of the hierarchical Coded VAE with two layers.
-We adopt the same variational family as used in the standard Coded VAE; however, in this case, we incorporate both hierarchical levels, leading to
+Inference follows a similar approach to the one employed in the Coded-DVAE, incorporating the linear combination of $\boldsymbol{m}_{1}$ and $\boldsymbol{m}_{2}$ used in the second branch.
+We adopt the same variational family as used in the standard Coded-VAE; however, in this case, we incorporate both hierarchical levels, leading to
 $q_{\boldsymbol{\eta}}(\boldsymbol{m},\boldsymbol{z}|\boldsymbol{x})=q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1}|\boldsymbol{x})q_{\boldsymbol{\eta}}(\boldsymbol{m}_{2}|\boldsymbol{x})p(\boldsymbol{z}_{1}|\boldsymbol{c}_{1})p(\boldsymbol{z}_{2}|\boldsymbol{c}_{2}),$ (13)
-where $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1}|\boldsymbol{x})$ is calculated following the same approach as in the coded DVAE with repetition codes, computing the all-are-zero and all-are-ones products of probabilities of the bits in $\boldsymbol{c}_{1}$ that are copies of the same message bit.
+where $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1}|\boldsymbol{x})$ is calculated following the same approach as in the Coded-DVAE with repetition codes, computing the all-are-zero and all-are-ones products of probabilities of the bits in $\boldsymbol{c}_{1}$ that are copies of the same message bit.
 The posterior $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{2}|\boldsymbol{x})$ considering both the encoder’s output and the inferred posterior distribution $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1}|\boldsymbol{x})$.
 Note that in this case, the decoder outputs the probabilities for both $\boldsymbol{c}_{1}$ and $\boldsymbol{c}_{2}$, with $\boldsymbol{c}_{2}$ being the encoded version of the linear combination $\boldsymbol{m}_{1,2}=\boldsymbol{m}_{1}\oplus\boldsymbol{m}_{2}$.
-Consequently, we first obtain $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1,2}|\boldsymbol{x})$ following the same approach as in the coded DVAE with repetition codes, and determine $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{2}|\boldsymbol{x})$ as
-$q_{\boldsymbol{\eta}}(\boldsymbol{m}_{2}|\boldsymbol{x})=\prod_{u=1}^{M}\text{Ber}(p_{u}),$ (13)
-$p_{u}=q_{\boldsymbol{\eta}}(m_{1,2,u}=1|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{1,u}=0|\boldsymbol{x})+q_{\boldsymbol{\eta}}(m_{1,2,u}=0|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{1,u}=1|\boldsymbol{x}).$ (14)
+Consequently, we first obtain $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1,2}|\boldsymbol{x})$ following the same approach as in the Coded-DVAE with repetition codes, and determine $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{2}|\boldsymbol{x})$ as
+$q_{\boldsymbol{\eta}}(\boldsymbol{m}_{2}|\boldsymbol{x})=\prod_{j=1}^{M}\text{Ber}(p_{j}),$ (14)
+$p_{j}=q_{\boldsymbol{\eta}}(m_{1,2,j}=1|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{1,j}=0|\boldsymbol{x})+q_{\boldsymbol{\eta}}(m_{1,2,j}=0|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{1,j}=1|\boldsymbol{x}).$ (15)
 After obtaining $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1}|\boldsymbol{x})$ and $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{2}|\boldsymbol{x})$, we recalculate the posterior bit probabilities for the linear combination $q^{\prime}_{\boldsymbol{\eta}}(\boldsymbol{m}_{1,2}|\boldsymbol{x})$ as
-$q^{\prime}_{\boldsymbol{\eta}}(\boldsymbol{m}_{1,2}|\boldsymbol{x})=\prod_{u=1}^{M}\text{Ber}(q_{u}),$ (15)
-$q_{u}=q_{\boldsymbol{\eta}}(m_{1,u}=1|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{2,u}=0|\boldsymbol{x})+q_{\boldsymbol{\eta}}(m_{1,u}=0|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{2,u}=1|\boldsymbol{x}).$ (16)
-Next, we apply the soft encoding approach to incorporate the repetition code structure at both levels of the hierarchy.The posterior probabilities $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1}|\boldsymbol{x})$ are repeated to obtain $q_{\boldsymbol{\eta}}(\boldsymbol{c}_{1}|\boldsymbol{x})$, and the posterior probabilities $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1,2}|\boldsymbol{x})$ are repeated to produce $q_{\boldsymbol{\eta}}(\boldsymbol{c}_{2}|\boldsymbol{x})$.
+$q^{\prime}_{\boldsymbol{\eta}}(\boldsymbol{m}_{1,2}|\boldsymbol{x})=\prod_{j=1}^{M}\text{Ber}(q_{j}),$ (16)
+$q_{j}=q_{\boldsymbol{\eta}}(m_{1,j}=1|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{2,j}=0|\boldsymbol{x})+q_{\boldsymbol{\eta}}(m_{1,j}=0|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{2,j}=1|\boldsymbol{x}).$ (17)
+Next, we apply the soft encoding approach to incorporate the repetition code structure at both levels of the hierarchy.
+The posterior probabilities $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1}|\boldsymbol{x})$ are repeated to obtain $q_{\boldsymbol{\eta}}(\boldsymbol{c}_{1}|\boldsymbol{x})$, and the posterior probabilities $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1,2}|\boldsymbol{x})$ are repeated to produce $q_{\boldsymbol{\eta}}(\boldsymbol{c}_{2}|\boldsymbol{x})$.
 Utilizing the reparameterization trick from Eq.
-5, we sample $\boldsymbol{z}_{1}$ and $\boldsymbol{z}_{2}$, concatenate them to form $\boldsymbol{z}$, and pass this through the decoder to generate $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})$.
+(5), we sample $\boldsymbol{z}_{1}$ and $\boldsymbol{z}_{2}$, concatenate them to form $\boldsymbol{z}$, and pass this through the decoder to generate $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})$.
 The model is trained by maximizing the ELBO, given by
@@
 As expected, $\mathbf{m}_{2}$ shows poorer error metrics compared to $\mathbf{m}_{1}$, since the model provides more redundancy to $\mathbf{m}_{1}$ incorporating it in both branches.
-Although the overall metrics and generation quality are somewhat similar to those of the coded DVAE with 10 information bits (see tables in Figures 2 and 4), the introduced hierarchy results in a more interpretable latent space.
-In this setup, $\mathbf{m}_{1}$ captures global features (such as clothing types in the FMNIST dataset), while $\mathbf{m}_{2}$ controls individual features, as we can observe in Figure 32, where we show examples of the model’s generative outputs for fixed $\mathbf{m}_{1}$ and random samples of $\mathbf{m}_{2}$.
-Figure 32: Examples of generated images using the hierarchical coded DVAE with (a) a 5/50 repetition code in each branch, and (b) a 5/100 repetition code in each branch.
+Although the overall metrics and generation quality are similar to those of the Coded-DVAE with 10 information bits (see tables in Figs.
+5 and 4), the introduced hierarchy results in a more interpretable latent space.
+In this setup, $\mathbf{m}_{1}$ captures global features (such as clothing types in the FMNIST dataset), while $\mathbf{m}_{2}$ controls individual features, as we can observe in Fig.
+34, where we show examples of the model’s generative outputs for fixed $\mathbf{m}_{1}$ and random samples of $\mathbf{m}_{2}$.
+Figure 34: Examples of generated images using the hierarchical Coded-DVAE with (a) a 5/50 repetition code in each branch, and (b) a 5/100 repetition code in each branch.
 In all the examples provided, $\boldsymbol{m}_{1}$ was fixed while $\boldsymbol{m}_{2}$ was randomly sampled.
-Table 10: Comparison of the obtained metrics for the coded DVAE with polar codes with different configurations, which we refer to as ’hierarchical coded DVAE’.
+Table 10: Comparison of the obtained metrics for the Coded-DVAE with polar codes with different configurations, which we refer to as ‘hierarchical Coded-DVAE’.
 Model BER WER Acc Conf. Acc PSNR
@@
 hier. 5/100 0.026 0.110 0.075 0.287
-Appendix H Ablation study
+Appendix K Ablation study
 In this section, we conduct ablation studies on the hyperparameter $\beta$ of the model, responsible for regulating the decay of exponentials in the smoothing transformation, as well as on the number of trainable parameters in the models.
-H.1 Ablation study on the hyperparameter $\beta$
+K.1 Ablation study on the hyperparameter $\beta$
 Across all experiments, we have consistently configured the hyperparameter $\beta$, which controls the decay of exponentials in the smoothing transformation, to a value of 15.
@@
 A priori, we would like these distributions to be separated, allowing us to retrieve the true value of the bit and effectively use the latent structure of the model.
-H.1.1 Coded model
+K.1.1 Coded model
 We first evaluate the influence of the parameter $\beta$ in coded models.
 We take as a reference the coded model with 8 information bits and a rate $R=1/30$, and train it using $\beta=5,10,15,20$.
-We assess the performance of the model in reconstruction and generation tasks.
-We observe the model is fairly robust, achieving similar performance across configurations in most metrics.
+We assess the performance of the model in reconstruction and
+Figure 35: Example of reconstructed images obtained with different values of $\beta$ using the coded model with an 8/240 code.
+generation tasks. We observe the model is fairly robust, achieving similar performance across configurations in most metrics.
 In Fig.
@@
 Larger values may result in unstable training and inferior performance.
-Figure 33: Example of reconstructed images obtained with different values of $\beta$ using the coded model with an 8/240 code.
 Table 12: Evaluation of reconstruction performance in FMNIST with different values of $\beta$ using the coded model with an 8/240 code.
@@
 Again, these two configurations appear to be the most suitable in this scenario.
-H.1.2 Uncoded model
+K.1.2 Uncoded model
 We first evaluate the influence of the parameter $\beta$ in uncoded models.
@@
 Figure 37: Example of reconstructed images obtained with different values of $\beta$ using an uncoded model 8 information bits.
-Figure 36: Example of randomly generated, uncurated images with different values of $\beta$ using an uncoded model 8 information bits.
 Table 15: Evaluation of reconstruction performance in FMNIST with different values of $\beta$ using an uncoded model 8 information bits.
@@
 20 13.717 0.464 0.466 13.743 0.460 0.462 0.383
+Figure 38: Example of randomly generated, uncurated images with different values of $\beta$ using an uncoded model 8 information bits.
 Table 16: Evaluation of the BER and WER in FMNIST with different values of $\beta$ using an uncoded model 8 information bits.
@@
 The models trained with $\beta=15$ and $\beta=10$ clearly outperform the other two in this task, generating more diverse and detailed images, and obtaining better error metrics and log-likelihood values.
-H.2 Ablation study on the number of trainable parameters
-A consistent architecture was employed across all experiments, which is detailed in Section A.
+K.2 Ablation study on the number of trainable parameters
+A consistent architecture was employed across all experiments, which is detailed in Section C.
 However, since the introduction of the code alters the dimensionality of the latent space, it is necessary to adjust the encoder’s output and the decoder’s input.
@@
 coded 8/240 adjusted 6,583,174 19,332,871
-H.2.1 Evaluation
+K.2.1 Evaluation
 This section provides an empirical evaluation of the models trained with the adjusted parameter count, demonstrating that the enhanced performance observed in coded models does not result from an augmented number of trainable parameters.
@@
 Figure 39: Example of reconstructed images obtained with different configurations.
-Figure 38: Example of randomly generated, uncurated images using different model configurations.
 Table 19: Evaluation of reconstruction performance in FMNIST with the adjusted parameter count.
@@
 coded 8/240 adj. 19.383 0.828 0.883 17.771 0.792 0.828 3.952
+Figure 40: Example of randomly generated, uncurated images using different model configurations.
 Table 20: Evaluation of the BER and WER in FMNIST with the adjusted parameter count.
@@
 coded 8/240 adj. -229.283 -238.302
-Appendix I Evaluating log-likelihood using the soft-encoding model
+Appendix L Evaluating log-likelihood using the soft-encoding model
 The reparameterization trick introduced in (5) requires that bits are independent, a condition not met in $\boldsymbol{c}$ once the code’s structure is introduced.
@@
 Remark that, despite the soft encoding assumption during training, the generative results in Fig.
-4, 6, 13, 19, 24, 24, and 29 are obtained through hard encoding.
+4, 8, 15, 21, 26, 26, and 31 are obtained through hard encoding.
 Namely, we sampled an information word $\boldsymbol{m}$ and obtained $\boldsymbol{c}$ by repeating its bits.
-While the hard-encoding images are visually appealing, to evaluate the coded DVAE LL for a given image $\boldsymbol{x}$, we have to leverage the soft-encoding model since it is unrealistic for a sample from the soft-encoding model to produce equally repeated bits.
-In the soft-coded model, we sample bit probabilities from a prior distribution that we model through the product of $M$ independent Beta distributions, and we use a proposal distribution model similar to the Vamp-prior in [28].
+While the hard-encoding images are visually appealing, to evaluate the Coded-DVAE LL for a given image $\boldsymbol{x}$, we have to leverage the soft-encoding model since it is unrealistic for a sample from the soft-encoding model to have equally repeated bits.
+In the soft-coded model, we sample bit probabilities from a prior distribution that we model through the product of $M$ independent Beta distributions, and we use a proposal distribution model similar to the Vamp-prior in Tomczak and Welling [2018].
 Namely, a mixture model with components given by $q(\boldsymbol{m}|\boldsymbol{x})$ for different training points.
-In the main text and Section H, we report the log-likelihood values for different model configurations trained on the FMNIST dataset.
+In the main text and Section K, we report the LL values for different model configurations trained on the FMNIST dataset.
 Table 4 presents the results obtained for the MNIST dataset.
@@
 For these high-dimensional datasets, we were unable to obtain valid log-likelihood estimates because the number of samples needed for importance sampling to converge was too large.
-However, given the difference in the level of detail in the reconstructed and generated images between coded and uncoded models (see Sections D and E), coded models are expected to better approximate the true marginal likelihood of the data.
-Appendix J Variational inference at codeword level
+However, given the difference in the level of detail in the reconstructed and generated images between coded and uncoded models (see Sections F and G), coded models are expected to better approximate the true marginal likelihood of the data.
+Appendix M Connection to previous works on VAEs as source coding methods
+Effective learning of low-dimensional discrete latent representations from complex data is a technically challenging task.
+In this work, we propose a novel method to improve inference in discrete VAEs within a fully probabilistic framework, introducing a new perspective on the inference problem.
+While previous studies have analyzed VAEs using rate-distortion (RD) theory [Chen et al., 2022, Townsend et al., 2019, Van Den Oord et al., 2017], our approach stems from a different yet complementary perspective that remains compatible with these works.
+In the literature, VAEs are often interpreted as lossy compression models (e.g., VAEs as source coding methods), but our contribution is best understood from a generative perspective.
+We conceptualize inference via $q(\boldsymbol{m}|\boldsymbol{x})$ as a decoding process, where the goal is to recover the discrete latent variable from the observed data.
+We sample a latent vector $\boldsymbol{m}$, generate a data point $\boldsymbol{x}$ (e.g., an image), and aim to minimize the error rate in recovering $\boldsymbol{m}$ from $\boldsymbol{x}$.
+Achieving this requires the variational approximation $q(\boldsymbol{m}|\boldsymbol{x})$ to closely match the true posterior $p(\boldsymbol{m}|\boldsymbol{x})$.We demonstrate that incorporating redundancy into the generative process reduces errors in estimating $\boldsymbol{m}$, leading to a more accurate latent posterior approximation.
+This improvement directly enhances our model’s overall performance, as confirmed by our experimental results.
+Additionally, error rates and variational gaps are related through bounds derived from mismatch hypothesis testing.
+These bounds demonstrate that minimizing the variational gap by maximizing the ELBO effectively tightens an upper bound on the error rate [Schlüter et al., 2013, Yang et al., 2024].
+In RD theory, the primary focus is on compression within the latent space, typically analyzed from an encoder/decoder and reconstruction (distortion) perspective.
+RD theory establishes theoretical limits on achievable compression rates and describes how practical models may diverge from these limits.
+Using RD practical compression methods, Townsend et al.
+[2019] demonstrates how asymmetric numeral systems (ANS) can be integrated with a VAE to improve its compression rate by jointly encoding sequences of data points, bringing performance closer to RD theoretical limits.
+Similarly, Chen et al.
+[2022] shows that a complex prior distribution in a VAE using an autoregressive invertible flow narrows the gap between the approximate and the true posterior distribution, thereby enhancing the overall performance of the VAE.
+We note that even using an independent prior, the hierarchical code structure outlined in Section J naturally decouples information across different latent spaces at various conceptual levels.
+Specifically, the most relevant information (class label) is captured by the most protected latent space, while the other space captures fine-grained features.
+This effect is not straightforward to enforce through direct design of a more complex prior.
+Our method complements all these efforts by introducing redundancy in the generative pathway to enhance variational inference, leading to more accurate latent posterior approximations.
+Notably, even when using a complex prior distribution, ECCs can still be leveraged to enhance inference and overall model performance.
+In this work, we demonstrate that our approach yields more robust models even with a simple, fixed independent prior, as evidenced by improved LL, generation quality, and reconstruction metrics.
+Moreover, in Section J, we show that integrating a hierarchical ECC with the same independent prior leads to even greater performance gains.
+It is also important to discuss the VQ-VAE [Van Den Oord et al., 2017], one of the most relevant works related to our approach, which is well-known for effectively learning compressed discrete representations of complex data.
+We believe that highlighting key differences between VQ-VAEs and our method can provide useful insights.
+A key distinction lies in the structure and dimensionality of the latent space.
+In image modeling, VQ-VAEs typically employ a latent matrix where indices correspond to codewords in a codebook.
+This design allows different codewords to capture specific patches of the original image, improving reconstruction and generation (the latter through an autoregressive prior on the latent representations).
+However, representing each data point as a grid of embeddings complicates interpretability.
+In contrast, our method encodes the entire image into the latent space, rather than partitioning it into patches.
+Another major difference is the nature of the encoder.
+VQ-VAEs rely on a non-probabilistic encoder that maps inputs to the nearest latent code using a distance-based metric.
+This deterministic mapping limits the model’s ability to capture and quantify uncertainty in the latent space.
+Our approach, on the other hand, operates within a fully probabilistic framework, allowing for uncertainty quantification in the learned latent representations.
+Additionally, our method is fully differentiable, enabling seamless gradient computation and backpropagation for end-to-end training.
+Our approach also allows us to leverage the reparameterization trick introduced in DVAE++ [Vahdat et al., 2018b], which we found to be more stable than continuous relaxations like the Gumbel-Softmax used in VQ-VAEs with stochastic quantization [Williams et al., 2020].
+Appendix N Variational inference at codeword level
 Here, we present an alternative variational family that is valid for any ECC, including random codes.
@@
 The KL Divergence term is approximated via Monte Carlo, drawing samples from $q_{\boldsymbol{\eta}}(\boldsymbol{c}|\boldsymbol{x})$.
-As it is not possible to backpropagate through discrete variables, we approximate the gradients with respect to the parameters of the encoder using the REINFORCE leave-one-out estimator [23, 13], given by
+Since it is not possible to backpropagate through discrete variables, we approximate the gradients with respect to the parameters of the encoder using the REINFORCE leave-one-out estimator [Salimans and Knowles, 2014, Kool et al., 2019], given by
 $\begin{gathered}\widehat{g}_{LOO}=\\ =\frac{1}{S-1}\Bigg{[}\sum_{s=1}^{S}f_{\boldsymbol{\eta}}\big{(}\boldsymbol{z}^{(s)},\boldsymbol{c}^{(s)}\big{)}\bigtriangledown_{\boldsymbol{\eta}}\log q_{\boldsymbol{\eta}}\big{(}\boldsymbol{c}^{(s)}|\boldsymbol{x}\big{)}-\overline{f}_{\boldsymbol{\eta}}\sum_{s=1}^{S}\bigtriangledown_{\boldsymbol{\eta}}\log q_{\boldsymbol{\eta}}\big{(}\boldsymbol{c}^{(s)}|\boldsymbol{x}\big{)}\Bigg{]},\end{gathered}$ (27)
@@
 Defining a distribution over the codebook can seem intuitive, but scalability becomes challenging as the size of the codebook increases.
-The reason is that the posterior distribution must be evaluated for all codewords during both inference and test time.
+The reason for this is that the posterior distribution must be evaluated for all codewords during both inference and test time.
 However, it can still provide a bound, enabling the utilization of more complex codes with theoretical guarantees that can outperform the previously proposed repetition codes.
@@
 11:  until convergence
-Appendix K Computational resources
+Appendix O Computational resources
 All the experiments in this paper were conducted on a single GPU.
@@
 Since training times varied significantly based on the hardware used, we were unable to provide comparable training times.
-Appendix L Coded DVAE scheme
-Figure 39: Graphic representation of the coded DVAE, illustrating both the generative and inference paths.
-Appendix M Hierarchical coded DVAE scheme
-Figure 40: Graphic representation of the hierarchical coded DVAE, illustrating both the generative and inference paths.
+Appendix P Coded-DVAE scheme
+Figure 41: Graphic representation of the Coded-DVAE, illustrating both the generative and inference paths.
+Appendix Q Hierarchical Coded-DVAE scheme
+Figure 42: Graphic representation of the hierarchical Coded-DVAE, illustrating both the generative and inference paths.